### v1

# MoA: Mixture of Sparse Attention for Automatic Large Language Model Compression

Abstract
Sparse attention can effectively mitigate the significant memory and throughput demands of Large Language Models (LLMs) in long contexts. Existing methods typically employ a uniform sparse attention mask, applying the same sparse pattern across different attention heads and input lengths. However, this uniform approach fails to capture the diverse attention patterns inherent in LLMs, ignoring their distinct accuracy-latency trade-offs. To address this challenge, we propose the Mixture of Attention (MoA), which automatically tailors distinct sparse attention configurations to different heads and layers. MoA constructs and navigates a search space of various attention patterns and their scaling rules relative to input sequence lengths. It profiles the model, evaluates potential configurations, and pinpoints the optimal sparse attention compression plan. MoA adapts to varying input sizes, revealing that some attention heads expand their focus to accommodate longer sequences, while other heads consistently concentrate on fixed-length local contexts. Experiments show that MoA increases the effective context length by 3.9×3.9$\times$3.9 × with the same average attention span, boosting retrieval accuracy by 1.57.1×1.5-7.1$\times$1.5 - 7.1 × over the uniform-attention baseline across Vicuna-7B, Vicuna-13B, and Llama3-8B models. Moreover, MoA narrows the capability gaps between sparse and dense models, reducing the maximum relative performance drop from $9\%-36\%$ to within $5\%$ across two long-context understanding benchmarks. MoA achieves a 1.21.4×1.2-1.4$\times$1.2 - 1.4 × GPU memory reduction and boosts decode throughput by $\times5$ 6.7 × for 7B and 13B dense models on a single GPU, with minimal impact on performance.
footnotetext: Corresponding authors: Yu Wang (yu-wang@tsinghua.edu.cn), Xuefei Ning (foxdoraame@gmail.com)
1 Introduction
Figure 1: Retrieval accuracy of the Vicuna-7B model using different attention methods across varying input lengths and retrieval positions on the LongEval dataset. This "Needle In a Haystack" experiment [19] takes massive key-value pairs as inputs and tests the accuracy to retrieve values based on given keys from diverse positions. (a) Original model with a full attention span; (b) StreamingLLM with half the attention span, showing reduced effectiveness beyond the span; (c) MoA with half the attention span, maintaining effectiveness beyond the span.
Large Language Models (LLMs) exhibit remarkable versatility across numerous applications [6, 57, 63]. Central to LLM is the attention mechanism [62], which computes interactions among tokens within a certain span, thereby enabling context understanding. Scaling input length is crucial for enhancing LLM capabilities [7, 60], including fact retrieval, summarization, few-shot learning, question answering and so on [4, 70]. However, the ever-growing attention computation and Key-Value Cache (KV-Cache) pose significant efficiency challenges [54, 69, 26, 33].
Previous work proposes sparse attention methods to address the efficiency challenges of long contexts in generative LLMs. These methods typically employ a uniform, fixed-span sliding window mask across all heads and input lengths, limiting attention to local contexts only [69, 26]. This approach allows the LLM to take long inputs with a fixed attention span, keeping bounded attention computation and KV caching overhead. Following previous works [7, 60], we quantify the effective context length as the maximum input length where content retrieval accuracy exceeds a 90% threshold. In principle, fixed-span local attention can gradually aggregate global information through multiple model layers, yielding a longer effective context length than each attention span [18]. Nonetheless, we reveal that uniform masks, like StreamingLLM [69], hardly extend effective context length beyond the span, as shown in Figure 6. Figure 1(b) further illustrates such limitation: with a 50% attention span mask, StreamingLLM fails to accurately retrieve content from the earlier half of the input and performs even worse at longer input lengths. Figure 2 reveals one possible explanation for the problem: while some attention heads focus on local contexts, others encompass the broad span of the entire input sequence. Consequently, the uniform approach fails to achieve a long effective context length as it limits the attention span of the global-context heads, while excessively allocates compute and memory budget for local-context heads. Additionally, as the input length increases, some attention heads need a faster increase in attention span than others to avoid serious performance degradation, as shown in Table 1. Unfortunately, the uniform approaches do not include heterogeneous rules to scale the attention spans differently for various heads. Besides, existing model compression methods [44, 40, 68, 36, 31, 37] use general language modeling corpora to decide the compression plan, which cannot accurately profile the influence of compression on long-context tasks.
In this work, we propose Mixture of Attention (MoA), a training-free sparse attention method. As illustrated in Figure 3, MoA constructs the search space of heterogeneous elastic rules of attention spans. For automatic LLM compression, MoA first utilizes gradient-based profiling to inspect the influences of each attention position on the prediction loss. Based on the profiling results, MoA tailors heterogeneous sparse attention configurations for each model layer and attention head. During profiling, MoA employs a calibration dataset with long-range dependencies and uses the original dense model’s response instead of the human-written response as the reference to calculate the loss. This ensures an accurate profiling of the attention influences to facilitate better compression results. Our contributions are summarized as follows.
  • Heterogeneous Elastic Rules. We propose heterogeneous elastic rules for masks of each attention head. We formulate MoA compression search space to include a diverse range of elastic rules that tailor the local attention span relative to the input length for each attention head. The heterogeneous elastic rules improve the fact retrieval accuracy of MoA from 25% to 98% compared with masks with uniform span and scaling function for each head.
  • Calibration Dataset Construction We emphasize the importance of data engineering in LLM compression. Our findings demonstrate that using datasets with long-range dependencies and referencing the original LLM’s responses are crucial for accurately profiling the influences of compression.
  • Automatic Optimization. We propose an automatic pipeline to find the optimal compression plan encompassing heterogeneous elastic rules for various attention heads. This pipeline can efficiently find the optimal plan within several hours, for example, two hours for compressing Vicuna-13B.
Experiments show that MoA achieves $\times5$ to 6.7×6.7$\times$6.7 × throughput improvements on 7B and 13B dense LLMs at a 50% density (the average of KV-Cache length / input length), with only 1% average relative degradation in retrieval accuracy. Additionally, MoA achieves over 90% retrieval accuracy with just 25% average density, far surpassing sparse attention baselines that need a density of 75% to 100% for similar performance. On long-context understanding benchmarks, MoA performs comparably to dense models, with a maximum relative performance drop of less than 5%, which is about one-sixth of that observed with the uniform sparse attention baseline. Our code is available at https://github.com/thu-nics/MoA.
2 Preliminary and Related work
2.1 Attention mechanism
The Multi-Head Self Attention (MHA) mechanism [62] is crucial to the functionality of LLMs. It starts with an input sequence transformed into query (Q), key (K), and value (V) matrices through linear projections. These matrices, combined with the cached K and V (KV-Cache) from previous sequences, compute the attention matrix (A). This calculation is modified by a causal mask (M) to ensure autoregressive properties, resulting in the output (O), as depicted in Equation 1:
$\mathbf{S}=\mathbf{Q}\mathbf{K}^{T},\quad\mathbf{A}=\text{softmax}(\mathbf{S}+\mathbf{M}),\quad\mathbf{O}=\mathbf{A}\mathbf{V}$ (1)
Autoregressive inference in LLMs involves two stages: prefill and decode. During prefill, the model processes the entire input sequence to generate the initial response token. In the subsequent decode stage, it uses the newly generated token and previously cached K and V matrices to produce subsequent tokens until the generation concludes. Although effective, this iterative process increases memory and computation demands due to the expanding KV-Cache.
2.2 Efficient Attention
Efficient methods are proposed to mitigate the computation and memory costs associated with attention. One branch of work uses dynamic sparse attention masks to adaptively skip attention computations during prefill stage [46, 52, 53, 64, 43, 32] or drop KV-Cache during decode stage [3, 75, 20, 54, 41] based on the input sequences. However, due to the complex control and computation flow, dynamic prefill often requires specific hardware to achieve substantial wall-time speedup [52, 64, 43, 25, 24]. Additionally, dynamic KV-Cache pruning in the decode stage may require extensive retraining [3], or additional accumulated attention score computation [54, 75, 41, 20].
Another branch of work uses static sparse attention, where predefined masks are applied consistently across all processed sentences. Thanks to the fixed computation flow, static sparse attention is generally more efficient and GPU-friendly. For language understanding models such as BERT [16], various masks are used [73, 5, 9, 76, 69, 26]. But for generative LLMs, the predominant method is the fixed-span sliding window mask with global attention on a few initial tokens [69, 26]. With the local attention pattern, the KV-Cache beyond the current attention span can be dropped, saving much memory for long sequence scenarios. However, the uniform static masks across different attention heads and input lengths are model- and data-agnostic, which can compromise LLMs’ effective context length and lead to suboptimal performance in long sequence scenarios. Our method falls within this category, benefiting from the efficiency and training-free advantages, while addressing the performance limitations encountered by previous methods.
In addition to sparse attention, alternative mechanisms have been proposed to replace traditional attention for long-sequence modeling [21, 49, 56, 51, 39, 30, 50, 10, 65]. However, these new mechanisms often require different weights compared to vanilla transformers, imposing significant re-training overhead for LLMs.
Previous works also propose LLM acceleration frameworks [22, 2, 54, 33], as well as kernel-level optimizations [13, 12]. These kernel and system optimizations are orthogonal to our work and can be integrated to further enhance efficiency.
3 Mixture of Attention (MoA)
We first illustrate the heterogeneity of the attention patterns in pre-trained LLMs in Section 3.1. Based on this insight, we define the search space for our Mixture-of-Attention (MoA) method in Section 3.2.
Figure 2: Examples of attention matrices from different attention heads of the Vicuna-7B model. Each attention matrix is averaged over 256 data items from the LongEval dataset.
Window/Input Len.
Layers 2k/4k 2k/8k 4k/8k
6, 7, 8 0.83 0.29 0.61
9, 10,11 0.99 0.81 0.96
17,18,19 0.97 0.94 0.97
Table 1: Retrieval accuracy of Vicuna-7B model using sliding-window sparse attention masks across various model layers, window spans, and input lengths.
3.1 Mixture of Attention Patterns and Elastic Rules
Heterogeneous Attention Patterns. Different attention heads in LLMs exhibit heterogeneous attention patterns, as shown in Figure 2. For example, the first head primarily focuses on local contexts with a narrow-span sliding window, while the third head covers nearly the entire input, indicating global attention. Table 1 demonstrates that applying the same sliding-window sparse attention mask across model layers can lead to a 65% variance in retrieval accuracies. It conforms to the multi-head self-attention design principle of capturing varied information [62], as well as the findings from concurrent research that identifies specific attention heads for global text retrieval [67].
Heterogeneous Elastic Rules. In addition to heterogeneity at a certain length, different attention heads also exhibit varying elastic behaviors as the input length changes. Figure 2 illustrates this variability: for shorter inputs (the upper left part of the attention matrix), the second and third heads initially show global attention. However, as input length increases, the second head remains the medium-span local focus, while the third head continues to expand as global attention. Table 1 further evidences the diverse elastic rules. For example, at 4k input length, a 2k sliding-window sparse attention mask on layers 9 to 11 yields better retrieval accuracy than on layers 17 to 19. However, the opposite is true for an 8k input length. This data supports the visual observations from Figure 2, highlighting that attention patterns respond to input length scaling differently. Leveraging these insights, MoA encompasses heterogeneous elastic rules as the search space.
3.2 Heterogeneous Elastic Rule Search Space
In designing the search space for the MoA mask, we consider the inherently heterogeneous and elastic nature of LLM attention patterns. As shown in Figure 3(a), we adopt a hardware-friendly sliding-window mask as our base sparse attention mask [5]. Following previous work [69, 26], the initial few tokens (64 tokens for MoA) are not masked. The attention span equals the sliding-window-span plus the number of initial unmasked tokens. For simplicity, we define the attention span S of head $h$ at input length N using a straightforward linear function:
$S_{h}=\alpha_{h}+\beta_{h}\times N,$ (2)
where $\alpha_{h}$ and $\beta_{h}$ are hyperparameters that control the base span and its expansion rate with input length. To ensure hardware-friendliness, we limit the number of distinct rules to at most two per model layer for MoA unless specified otherwise.
The $\alpha$ and $\beta$ hyperparameters of each attention head can be chosen from multiple discrete options. For LLMs with many layers, the MoA search space can become quite large. For example, for a 7B model consisting of $32$ attention heads and $32$ layers, with $6$ and $9$ different $\alpha$ and $\beta$ options for each head, the potential search space expands to $54^{1024}$ configurations. Thus, we design the automatic pipeline to efficiently pinpoint the optimal $\alpha$s and $\beta$s for any LLM.
4 Automatic Pipeline for MoA Compression
Figure 3: Overview of the MoA. (a) The sparse attention search space includes heterogeneous elastic rules of the attention span on sliding-window masks. (b) The automatic compression pipeline begins with a calibration dataset, which includes long-dependency contexts and supervision texts generated by the original dense LLM. MoA profiles each attention value’s impact on model predictions within this dataset, revealing accuracy losses for different candidate elastic rules across various input lengths. The final optimization step selects elastic rules for each attention head to minimize the total prediction loss while adhering to specified density constraints.
This section outlines the MoA automatic compression pipeline as shown in Figure 3(b). Starting with a trained LLM and a calibration dataset, MoA first profiles the influence of each attention value on the model’s prediction loss for various input sequences from the calibration dataset. The masked sum of the influences represents the accuracy loss associated with each mask at different input lengths, showing the accuracy loss each candidate elastic rule could cause at that length. Then, MoA optimizes the compression plan by selecting the optimal elastic rule for each head, which minimizes the accuracy loss across various lengths while adhering to specified density constraints. The following sections provide detailed discussions of each step in this pipeline.
4.1 Attention Influence Profiling
In the profile step, MoA quantifies the impact of individual attention values on the final prediction loss of a pre-trained LLM. It informs the subsequent step about the influence of masking each attention value, revealing the accuracy trade-offs of the candidate elastic rules for each attention head.
The influence of each attention value is derived from the attention matrix A and its gradient $\partial L/\partial\textbf{A}$, computed over a calibration dataset. When applying sparse attention masks, we approximate the change in the model’s prediction loss, $\Delta L$, using a first-order Taylor expansion based on variations in the attention matrices A: $\Delta L=\sum_{h}\sum_{i}\sum_{j}\partial L/\partial A_{h,i,j}\cdot\Delta A_{h,i,j}$. Here, $h$ indexes the attention heads across all layers, and $i,j$ are the row and column indices within each attention matrix $\textbf{A}_{h}$.
We define the attention influence matrix, $\textit{E}_{h,i,j}$, as the estimated change in loss, $\Delta L$, if the attention value $A_{h,i,j}$ is masked (i.e., set to zero). As shown in Equation 3, this measure considers both the direct and indirect effects of the mask. For notation simplicity, we omit the head index $h$ here. Initially, masking directly reduces the attention value to zero, represented by $\Delta A_{i,j|j}=-A_{i,j}$. Additionally, the softmax function in attention normalizes the sum of each row in the attention matrix to one. Thus, setting one attention value at column $j$ to zero causes an increase in the other attention values, $\Delta A_{i,n|j},n\neq j$, within the same row. These two effects are integrated into the following formulation, whose derivation is provided in Appendix D.1:
$E_{i,j}=\sum_{n}\frac{\partial L}{\partial A_{i,n}}\cdot\Delta A_{i,n|j}=\frac{\partial L}{\partial A_{i,j}}\cdot(-A_{i,j})+\sum_{n\neq j}\frac{\partial L}{\partial A_{i,n}}\cdot A_{i,n}\cdot\frac{A_{i,j}}{1-A_{i,j}}$ (3)
In practice, we use backpropagation on a calibration dataset to calculate the average attention influence $\bar{\textbf{E}}_{h}$ of each head across data items. The average attention influence is calculated respectively for different input lengths. The gradient ${\partial L}/{\partial\textbf{A}_{h}}$ is computed using chain derivative in deep learning frameworks like PyTorch [48]. The detailed calibration dataset setup is discussed in Section 5.
With the average attention influence of each head, MoA can calculate the accuracy loss of applying a candidate elastic rule at a specific input length. The loss is calculated with the sum of masked average attention influence according to the rule. We denote $\textbf{M}_{r_{h}}$ as the binary mask at head $h$ that corresponds to rule $r$, with masked positions marked as 0 and others as 1. We formalize accuracy loss $\Delta L$ as follows:
$\Delta L=\sum_{h}\Delta L_{h,r_{h}}=\sum_{h}\sum_{i}\sum_{j}M_{r_{h},i,j}\bar{E}_{h,i,j}.$ (4)
After the profile stage, MoA acquires the unique accuracy-density trade-offs of elastic rules. It informs the allocation of denser masks to more sensitive heads and lighter masks to less sensitive ones. Profiling at different input lengths enables the identification of the most effective elastic rules, even for unseen lengths.
4.2 Automatic Optimization
MoA automatically selects the optimal elastic rule for each attention head to minimize accuracy losses across various sequence lengths under density budgets. Based on the profiling results, MoA first identifies Pareto front compression plans where any improvement in accuracy loss at one profile length would worsen another. To ensure the best generalization to lengths beyond those profiled, MoA then selects the plan that yields the minimum loss at an unseen length among the Pareto front solutions as the final plan.
Specifically, we utilize multi-objective optimization to search for a set of Pareto optimal compression plans across the profiled lengths. The objective for each length is to minimize the total accuracy loss while conforming to any user-defined density constraints. The objective is formulated as follows:
$\mathop{\arg\min}_{r_{h}\in\mathbb{R}}\Delta L^{(N_{i})},N_{i}\in\mathbb{N}_{\text{profile}}\quad\operatorname{s.t.}\frac{1}{H}\sum_{h=1}^{H}d^{(N_{i})}_{r_{h}}\leq d^{(N_{i})}_{\text{constr}},\forall N_{i}\in\mathbb{N}_{\text{constr}}.$ (5)
Here, superscript $(N)$ denotes values at different lengths; $\mathbb{N}_{\text{profile}}$ and $\mathbb{N}_{\text{constr}}$ denote the sets of lengths for profiling and those subject to density constraints, respectively; $\mathbb{R}$ denotes the set of candidate rules; $\Delta L^{(N_{i})}$ denotes the accuracy loss due to compression; $d_{r_{h}}^{(N_{i})}$ denotes the density of rule $r_{h}$ at head $h$; $d_{\text{constr}}^{(N_{i})}$ denotes the average density constraint; $H$ denotes the total number of attention heads.
Such formulation corresponds to the classic multi-objective mixed-integer-planing problem, which can be effectively solved within minutes using existing linear solvers, like Gurobi [23]. The detailed problem formulation and solving strategies are discussed in Appendix D.2.
Among the Pareto optimal compression plans, we select the one with the minimum loss at the unseen validation length as the optimal solution. This approach allows us to avoid profiling at every possible length while increasing the likelihood that the plan will generalize effectively to unseen lengths.
Thanks to this automatic pipeline, we efficiently get the elastic rules tailored for each attention head. With the pipeline, MoA minimizes the accuracy loss caused by attention sparsification, while conforming to user-defined density constraints.
5 Dataset and Supervision
Table 2: Calibration dataset design choices: dataset content, supervision, and response reference. Calibration dataset with long dependency and model alignment improves MoA performance on retrieval accuracy and perplexity. All tests are done at 25% average density at 8k input length.
Dataset Supervision Reference Long Dep. Align Model Retrieval Acc. $\uparrow$ PPL $\downarrow$
RedPajama Context - 0.25 4.95
MultiNews Context & Summary Human ✗/✓ 0.27 4.62
MultiNews Summary Human 0.87 3.97
MultiNews Summary Model 0.95 3.96
In this section, we highlight the overlooked importance of calibration dataset design and its supervision objective in LLM compression. Calibration datasets are essential for sensitivity analysis across various compression techniques, including weight pruning [44, 34, 42] and quantization [40, 68, 36, 31]. In this work, MoA profiles the attention influence on the calibration dataset, which is crucial for subsequent automatic optimization.
Current Approach. General language modeling datasets, such as human-written text corpus RedPajama [11], are commonly used as the calibration dataset. These datasets, supervised by next-token-prediction on the entire corpus, primarily capture attention patterns coherent with immediately preceding tokens. However, they lack long context dependencies, failing to address the global attention crucial for tasks like long-range retrieval.
Moreover, a notable misalignment exists between the model response and the human-written supervision. Consequently, it leads to inaccuracies when using human responses to compute attention values and gradients during profiling. For example, given the same question, a human might answer ’Blue’, while the model could generate ’The blue color’. Using the human answer for supervision, attention influence is inaccurately quantified based on probability shift for predicting ’Blue’; this diverges from the objective of maintaining crucial attention for the original model prediction, ’The’. These inconsistencies arise from various factors, including mismatched positions, tones, and synonyms.
MoA’s Approach. MoA enhances the calibration dataset by integrating long-range dependencies and model alignment. It utilizes the long-contextual MultiNews dataset [17], which includes summaries heavily dependent on long-range content. The summaries are generated by the original dense model and serve as supervision. Compared to current approaches that adopt human responses as the reference to calculate the loss, using the responses generated by the original model as the supervision can facilitate accurate influence profiling, thus benefiting the compression results.
Approach Comparison. We evaluate our design’s effectiveness by varying dataset choices, supervision types, and summary references, while standardizing data item count and length to 50 and 8k words, respectively. Additional setups and evaluations are in Appendices A and B.3.1.
We show the importance of long-range dependencies by comparing the MoA compression plan generated with different datasets and supervisory methods. In Table 2, RedPajama [11] represents the general language modeling dataset, while MultiNews [17] highlights long-range contexts by aggregating multiple documents on a single incident. Additionally, each MultiNews item includes a human-written summary, providing even stronger long-range dependencies and better performance. Calculating loss on the summary of MultiNews leads to significantly better performance, with a 60% increase in retrieval accuracy and a 0.98 decrease in perplexity.
Furthermore, using summaries generated by the original dense model as supervision promotes higher alignment between its own attention patterns and the text supervision. It improves performance compared to potentially inconsistent human summaries, as shown in the last two rows of Table 2.
6 Experiment
6.1 Setups
We brief the experiment setups here, with more details in Appendix A.
Baselines. We compare MoA with state-of-the-art static and dynamic sparse attention baselines for LLMs: StreamingLLM [69] and H2O [75]. We define the density of an LLM as the average KV-Cache length divided by the sequence length during the sparse decode stage. Notably, in MoA and StreamingLLM, the KV-Cache length matches the attention span during their sparse prefill stages. In contrast, H2O employs a dense prefill and requires additional computations for dynamic mask generation based on accumulative attention scores.
Models and Benchmarks. We evaluate on vicuna-7b-v1.5-16k, vicuna-13b-v1.5-16k [8] from LMSys, and Llama-3-8B-Instruct-262k [1] from Gradient AI. For effective context length evaluation, we use LongEval [35] to test key-value retrieval accuracy with 100 data items per length level. For comprehensive ability evaluation, we use LV-Eval [70] and LongBench [4], which include 11 and 13 sub-tasks, respectively. For coherence testing, we measure perplexity on four subsets of LongBench [15, 17, 38, 28, 45], representing different capabilities.
MoA Settings. We profile MoA on MultiNews [17] with model summaries at 2k, 4k, and 8k lengths. The optimal compression plan is selected with the validation dataset at 12k. Unless specified, the same plan is used across all benchmarks and lengths for each model. The models are not finetuned.
6.2 Performance
Table 3: Comparative analysis of retrieval accuracy, LV-Eval scores, LongBench scores, and perplexity for various models with different attention methods. All methods employ 50% density in both prefill and decode stages, except for H2O, which uses dense prefill.
Retrieve Acc. $\uparrow$ LV-Eval $\uparrow$ LongBench $\uparrow$ PPL $\downarrow$
Model Attention 4k 8k 16k 16k 0-4k 4-8k 8-16k 8-12k
Vicuna-7B Original 1.00 0.98 0.62 5.93 37.91 33.82 32.54 3.79
H2O 0.86 0.68 0.35 5.42 36.23 32.74 31.81 3.94
StreamingLLM 0.43 0.16 0.08 4.72 30.53 33.28 31.70 4.48
MoA 1.00 0.97 0.57 5.61 37.04 32.90 31.94 3.75
Vicuna-13B Original 0.99 0.98 0.44 5.83 42.25 39.52 35.93 3.62
H2O 0.88 0.76 0.28 5.66 41.63 38.02 34.75 3.80
StreamingLLM 0.65 0.49 0.33 5.43 30.65 33.07 32.68 4.10
MoA 0.99 0.93 0.49 7.16 41.73 38.88 35.69 3.62
Llama3-8B Original 0.99 0.99 0.97 17.49 44.27 43.53 43.26 4.52
H2O 0.94 0.89 0.88 16.03 43.46 43.01 42.50 4.63
StreamingLLM 0.68 0.55 0.52 11.16 37.20 38.02 39.43 4.79
MoA 0.99 1.00 1.00 17.46 43.07 42.75 43.09 4.49
Overall Performance. As detailed in Table 3, MoA outperforms state-of-the-art sparse attention methods and achieves comparable performance to the original dense model at 50% density. We calculate the relative performance drop of the compressed model over the original one. For retrieval tasks, MoA demonstrates a maximum of 8% relative accuracy drop (calculated as $\max\{1-\text{Acc.}_{\text{MoA}}/\text{Acc.}_{\text{Original}}\}$ across three lengths and LLMs), significantly less than the 87% and 44% for StreamingLLM and H2O. On average, the relative accuracy drop for MoA is under 1%, while it is much higher for StreamingLLM and H2O at 51% and 20%, respectively. In benchmarks using LV-Eval and LongBench, MoA shows only a 5% and 3% maximum relative drop in scores, respectively, compared to steep declines for StreamingLLM at 36% and 27%. H2O exhibits maximum drops of 9% and 4% with higher efficiency costs. Similar trends show in perplexity tests, where MoA maintains less than 1% relative perplexity increase, while StreamingLLM and H2O exhibit 13% and 4% increases. This trend holds for other densities, as shown in Appendices B.1.1 and B.1.3.
Figure 4: Retrieval accuracy tests on LongEval. With Vicuna-7B: (a) varies input lengths and densities to show effective context lengths across attention spans, (b) sets input length at 8k and shows retrieval accuracy across different densities. With Llama3-8B: (c) sets density at 50% and shows retrieval accuracy across increasing input lengths.
Long-Context Retrieval. We investigate the ultimate effective context length with retrieval tests on LongEval [35]. As shown in Figure 4(a), MoA expands its effective context length to approximately 3.9×3.9$\times$3.9 × its attention span before reaching up to the 12k limit of the original model. Detailed data for effective context length calculation are in Appendix B.1.2. Figure 4(b) shows that at a fixed input length of 8k, MoA reaches over 0.9 retrieval accuracy with just 25% density, whereas StreamingLLM and H2O require 100% and 75% density, respectively. Figure 4(c) further investigates MoA’s ability to expand and generalize to extreme input lengths. By profiling within 8k lengths, MoA extends its capability to 60k inputs while maintaining over 90% retrieval accuracy at 50% density.
Figure 5: (a) LV-Eval and (b) LongBench scores for different attention methods at 50% density, tested on Vicuna-7B and 13B models. Scores normalized against the original dense model.
Long-Context Understanding. Figure 5 illustrates the performance comparison across long-context benchmarks LV-Eval and LongBench. MoA achieves comprehensive performance comparable to the original dense model, as well as H2O that requires higher efficiency cost. In contrast, StreamingLLM displays inconsistent performance: although it occasionally surpasses the original model, it generally suffers noticeable degradation.
Ablation Study. We evaluate the performance impact of different sparse mask search spaces in Table 5. Starting with a basic uniform mask, we observe significant enhancements by sequentially introducing heterogeneity: layers first, then heads, and finally elastic rules.
6.3 Efficiency
Table 4: Efficiency analysis of different frameworks on 7B and 13B models. H2O and MoA use 50% density. GPU memory evaluated with batch sizes 8 (7B model) and 4 (13B model). Decode throughput evaluated at the maximum batch capacity of an A100-80GB GPU.
Memory (GB) Throughput (tok/s)
Size Framework 4k 8k 16k 4k 8k 16k
7B FlashAttn2 28.5 44.4 76.3 134.6 66.9 32.9
H2O 36.9 OOM OOM 754.9 296.3 51.7
MoA 22.7 32.9 53.5 897.7 436.7 206.4
13B FlashAttn2 36.8 49.2 74.0 81.3 40.8 19.8
H2O 40.4 77.9 OOM 330.2 138.2 37.4
MoA 32.0 39.6 55.0 471.5 222.6 108.3
Table 5: Ablation study on search space with consistent 25% density, progressively introducing heterogeneity in layers, heads, and elastic rules. Evaluations are done with retrieval accuracy and perplexity.
Retrieval Acc. PPL
Mask Design 8k 16k 8k 12k
Uniform 0.25 0.15 4.89 5.19
+Hetero. Layers 0.31 0.26 4.55 4.85
+Hetero. Heads 0.95 0.41 3.96 4.30
+Elastic 0.98 0.43 3.96 4.29
We use Huggingface [66] with FlashAttention [13] as the baseline and implement MoA based on it. As shown in Table 5, with an average 50% density, MoA reduces total GPU memory by 1.2×1.2$\times$1.2 × to 1.4×1.4$\times$1.4 × across various model sizes and input lengths. Moreover, MoA boosts decode throughput by $\times5$ to 6.7×6.7$\times$6.7 ×, primarily attributed to a static-size KV-Cache (3.0$\approx$ 3.0$\times$ 3.0 ×), reduced attention computations (1.5$\approx$ 1.5$\times$ 1.5 ×) and increased batch sizes (1.4$\approx$ 1.4$\times$ 1.4 ×). Compared with H2O that uses dense prefill and requires additional cost for dynamic mask generation, MoA achieves $1.2$ to $4.0$ throughput at the same 50% density. Compared to the highly optimized vLLM framework [33], MoA still achieves a 1.4×1.4$\times$1.4 × to 1.5×1.5$\times$1.5 × increase in throughput. More efficiency results are shown in Appendix B.2.
6.4 Rules Discovered by MoA
We investigate MoA’s elastic rules for each head. As shown in Figure 8, masks in the initial and middle layers exhibit high density, aligning with the conclusions from previous research on LLM’s intrinsic dimensions [61] and layer sensitivities [71]. Conversely, in the final layers, most heads require low density, while few need high density. Figure 9 shows that layers with lower average mask density typically display a wider range of densities among heads, confirming the need for heterogeneity within the same layer. Further details and quantified insights are presented in Appendix C.
7 Conclusion and Future Work
MoA automates the selection of heterogeneous elastic masks for each attention head and input length, significantly extending the effective context length of LLMs by 3.9×3.9$\times$3.9 ×. It enhances retrieval accuracy by 1.5×1.5$\times$1.5 × to 7.1×7.1$\times$7.1 × and increases throughput to over 5×5$\times$5 × at 50% average density, maintaining performance on par with dense models in rigorous benchmarks.
Limitations and Future Work. Under an extremely low-density budget, MoA fails to maintain good performance. Designing a dynamic MoA method has the potential to address this issue, which we leave for future work. To further enhance MoA’s efficiency, system, and kernel-level optimizations such as kernel fusion [27] and KV-Cache management [33] could be integrated. Using non-linear elastic rules with bounded attention spans is also worth exploring. Additionally, MoA’s profiling method can be adapted to evaluate the influence of weights and other activations, facilitating other compression methods such as weight and activation quantization [40, 68, 37].
Acknowledgement
This work was supported by National Natural Science Foundation of China (No. 62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164), Tsinghua EE Xilinx AI Research Fund, and Beijing National Research Center for Information Science and Technology (BNRist).
Appendix A Detailed Experiment Setup
A.1 Main Setup
Baselines. In the setup for our experiment, we adhere to specific configurations outlined in the respective papers. In the case of StreamingLLM [69], the initial four tokens remain unmasked, serving as the attention sink. For H2O [75], we ensure the same number of heavy hitter tokens and recent tokens. Note that H2O uses dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction. StreamingLLM and MoA use sparse prefill.
Models and Benchmarks. Since vicuna-7b-v1.5-16k and vicuna-13b-v1.5-16k [8] can only take in 16k context length, we use the 16k split of LV-Eval benchmark [70], truncating the input to 15500 for model input. For the LongBench benchmark [4], we use the LongBench-E split, which features a balanced number of data items at every length level. The LongBench dataset is segmented into ranges of 0-4k, 4-8k, and 8k+ tokens. We test each split using the input length truncation thresholds of 3,500, 7,500, and 15,500 tokens, respectively.
Perplexity Evaluation. We construct a comprehensive yet concise test set by sampling $50\times 4$ data items for each length level from the test split of four long-context understanding datasets: Qasper [15], MultiNew  [17], TREC [38, 28] and LCC [45], representing the question answering, summarization, few-shot learning, and code completion abilities of the LLM. Following LongBench, the data items are organized as question-answer pairs. The questions and answers are written by humans and come with the dataset. The perplexity is calculated solely on the answer part of the data, demonstrating the model’s coherence in responding to user requests.
Validation Dataset. The validation dataset is used to select the optimal compression plan among the Pareto front solutions during the optimization step. The validation dataset is similarly constructed as the perplexity test dataset, but on the respective validation split of the datasets. $50\times 4$ data items are sampled from the same four long-context understanding datasets: Qasper [15], MultiNew  [17], TREC [38, 28] and LCC [45]. The additional $50$ data items from the LongEval [35] dataset are also added to validate the retrieval ability. For the datasets that do not contain the validation split, namely TREC, MultiNews and LCC, we sample from the test split and ensure different data items with the perplexity evaluation dataset.
MoA Settings. MoA uses the block sparse attention pattern with a block size of 64, where each grid depicted in Figure 3(a) represents a block. The first block of tokens is not masked as the attention sink. For the profile stage, we use the MultiNews [17] calibration dataset with model response as supervision, as described in Section 5. We use $50\times 3$ data items at 2k, 4k, 8k lengths. The data items are padded to their corresponding length level in order to ensure a unified shape of attention influence tensors for each length level. We adopt block granularity during the profiling stage, calculating the average attention influence within each block to represent the block’s overall influence. The optimization is done with the multi-objective optimization at the same set of lengths. Among the Pareto front solutions, we select the one with the lowest perplexity on the validation dataset of length 12k.
A.2 Efficiency Experiment Setup
We test the efficiency of different frameworks using a single NVIDIA A100-SXM4-80GB GPU. To improve the runtime profiling accuracy, we first run five forward passes as warmups. Then we use torch.CudaEvent to calculate the runtime for each method. Our experiments are structured around three scenarios: including prefilling 3k tokens and decoding 1k tokens; prefilling 6k tokens and decoding 2k tokens; prefilling 12k tokens and decoding 4k tokens. The labels are marked by the total sequence length, which equals prefill length plus decode length.
For MoA, The implementation is based on Huggingface Transformers. During the prefill stage, we use the sparse triton [58] kernel designed by us with block size 64. During the decode stage, we modify the KV-Cache implementation to support our heterogeneous elastic rules. Thanks to our fixed sliding-window span during the decode stage, we simply replace the old KV-Cache that exceeds the span with the latest KV-Cache. Then, the dense FlashAttention [13] decoding is applied with the updated KV-Cache.
For H2O, we use its official efficient implementation, which is based on Flexgen [54]. Note that H2O uses dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction, which requires the attention matrix to be explicitly calculated. It makes H2O’s prefill stage currently incompatible with kernel optimizations like FlashAttention. Therefore, H2O is easy to get OOM (Out-Of-Memory) with large prefill length and increased batch size.
In our efficiency tests across all frameworks, we implemented a simple optimization at the language modeling head (lm head) during the prefill stage. Specifically, after the final layer of the transformers, we compute the logits—these are the raw outputs that are transformed into probabilities—for only the last token. This selective computation avoids generating these probabilities for preceding tokens, substantially reducing both computational overhead and memory usage. We also set the environment variable PYTORCH_CUDA_ALLOC_CONF to be expandable_segments:True for Hugginface and MoA to mitigate memory fragmentation, allowing larger inference batch size.
Following the performance experiments, we use Vicuna-7B and Vicuna-13B for efficiency tests whenever possible. However, the official efficient implementation of H2O based on Flexgen only supports OPT [74]. Therefore, we use OPT-6.7b and OPT-13b models for H2O in Table 5 for comparison.
A.3 Ablation Study Setup
In the ablation study in Table 2 and Table 5, we use 25% density instead of the 50% used in the main experiment in Table 3. This decision is based on the observation that at a density of 50%, the performance of the various designs is quite similar, making it difficult to discern significant differences. In contrast, a lower density of 25% reveals more pronounced disparities between the designs, providing a clearer basis for comparison.
In the calibration dataset experiments in Table 2, we intentionally exclude the influence of the validation dataset. We avoid using the validation dataset by profile and optimize solely at 8k length, reducing the multi-objective optimization problem to a single-objective one with only one optimal compression plan instead of a set of Pareto fronts.
Appendix B Additional Experiment Results
B.1 Performance
B.1.1 Overall Performance
Table 6: Comparative analysis of retrieval accuracy, LV-Eval scores, LongBench scores, and perplexity for various models with different attention methods. All methods employ 75% density in both prefill and decode stages.
Retrieve Acc. $\uparrow$ LV-Eval $\uparrow$ LongBench $\uparrow$ PPL $\downarrow$
Model Attention 4k 8k 16k 16k 0-4k 4-8k 8-16k 8-12k
Vicuna-7B StreamingLLM 0.91 0.35 0.09 4.30 36.39 32.44 31.04 3.92
MoA 1.00 0.97 0.58 5.67 38.07 33.80 31.75 3.78
Vicuna-13B StreamingLLM 0.73 0.81 0.37 5.65 36.77 34.65 33.43 3.70
MoA 0.99 0.97 0.42 5.57 41.85 39.76 36.06 3.62
Llama3-8B StreamingLLM 1.00 0.83 0.76 14.89 42.45 40.62 42.51 4.51
MoA 0.99 1.00 0.93 15.61 43.51 43.16 43.58 4.53
Table 6 shows the overall performance of MoA at a higher density of 75%. MoA shows improved performance over the baseline with the uniform attention baseline. The progressive change of performance with respect to different densities is also shown in Figure 4(b) and Figure 7
B.1.2 Long-Context Retrieval
Figure 6: Retrieval accuracy of Vicuna-7B model using different attention methods across varying attention spans and input lengths. The X-axis shows different attention spans; the Y-axis shows different input lengths for the retrieval task. Subfigure (a) shows results for StreamingLLM, and subfigure (b) for MoA.
We conduct a detailed experiment to test the retrieval ability of different attention methods across various attention spans and input lengths. Following previous work [7, 60], we quantify effective context length as the maximum input length where retrieval accuracy remains above a 90% threshold. Figure 6 shows the detailed data for effective context length calculation. As shown in the figure, StreamingLLM can hardly maintain retrieval accuracy when the input length is beyond the attention span, while MoA can effectively extend the effective context length.
B.1.3 Long-Context Understanding
Figure 7: LV-Eval score of StreamingLLM and MoA at various densities on Vicuna-7B model.
We conduct experiments with various densities on the LV-Eval benchmark [70]. As shown in Figure 7, MoA constantly outperforms the uniform static attention baseline StreamingLLM at various densities, demonstrating the effectiveness of our heterogeneous elastic rules.
B.2 Efficiency
B.2.1 Throughput Breakdown
Table 7: Improvements in throughput (tokens per second) across various configurations and input sizes for MoA over FlashAttn2 on models with 7b and 13b parameters. The batch sizes for FlashAttn2 are configured to the maximum capacity of an A100-80GB GPU.
4k 8k 16k
Size Configuration Batch Throughput Batch Throughput Batch Throughput
7B FlashAttn2 30 134.6 15 66.9 8 32.9
+Static KV-Cache 30 496.1 15 219.5 8 91.6
+Reduced Attention 30 722.5 15 369.9 8 178.3
+Increased Batch Size 50 897.7 25 436.7 12 206.4
13B FlashAttn2 16 81.3 8 40.8 4 19.8
+Static KV-Cache 16 264.6 8 111.3 4 62.2
+Reduced Attention 16 329.6 8 156.4 4 87.3
+Increased Batch Size 28 471.5 14 222.6 7 108.3
In this section, we conduct a detailed analysis of the throughput breakdown for MoA, compared to the baseline comprising Huggingface with FlashAttention. The observed increase in throughput, as shown in Table 7, primarily stems from three aspects:
Static KV-Cache. MoA only maintains the tokens within the span of each head, thereby preventing growth in the KV-Cache size. This strategy eliminates the need for additional memory allocation.
Reduced Attention Computation. MoA with features reduced density in attention span and KV-Cache. It decreases the computation and memory access required for attention computation.
Increased Batch Size. With the reduced size of KV-Cache, MoA supports a larger batch size, contributing to the increase in throughput.
B.2.2 Comparison with vLLM
Table 8: Throughput comparison of MoA and vLLM on 7B and 13B models.
Throughput (tok/s)
Size Framework 4k 8k 16k
7B vLLM 628.8 323.0 145.5
MoA 897.7 436.7 206.4
13B vLLM 314.8 160.5 71.1
MoA 471.5 222.6 108.3
We compare the throughput of MoA with the highly optimized inference framework vLLM [33] as shown in Table 8. MoA still achieves a 1.4×1.4$\times$1.4 × to 1.5×1.5$\times$1.5 × increase in throughput. Notably, vLLM incorporates extensive system-level optimizations that are orthogonal to MoA, suggesting the potential for MoA to further improve. As vLLM constantly pre-allocates a 71GB memory and automatically manages the actual batch size, we exclude memory metrics from the comparison.
B.2.3 Automatic Optimization Overhead
Table 9: Time overhead for various stages of MoA.
Stage Time
Calibration Dataset Generation 15 min
Profile 25 min
Optimize 25 min
Validate 40 min
Sum. 105 min
We present a detailed breakdown of the time usage of MoA pipeline. Table 9 summarizes the time required for various crucial phases within the MoA framework, encompassing calibration dataset generation, profiling, optimization, and validation, on the Vicuna-13B model.
Profiling is the most resource-demanding part of our pipeline. For a 13b model with an 8k profile length, two A100 GPUs are required. In other cases, we only need one single GPU. Profiling on a 13b model with an 8k profile length and 50 data items takes 15 minutes. Profiling on 4k and 2k lengths takes less than 5 minutes each.
On the Intel(R) Xeon(R) Platinum 8358 2.60 GHz CPU, the optimization concludes within approximately 25 minutes. Typically, this phase generates around 10 compression plans. Validating each one of the compression plans takes about 4 minutes, totaling around 40 minutes.
B.3 Ablation Study
B.3.1 Calibration Dataset
Table 10: Performance comparison on various test sets, using different calibration sets. Tested on Vicuna-7B model. The result is tested with 50% density MoA on LongBench [4] 0-4k split.
Test Score
Dataset Long Dep. & Align Model Qasper MultiNews TREC Avg. Score
Original NA 28.6 28.2 56.0 37.6
RedPajama 20.6 (-8.0) 19.6 (-8.6) 66.0 (+10.0) 35.4 (-2.2)
Qasper 25.6 (-3.0) 27.8 (-0.4) 55.0 (-1.0) 36.1 (-1.5)
MultiNews 29.0 (+0.4) 27.5 (-0.7) 54.0 (-2.0) 36.8 (-0.8)
TREC 27.3 (-1.3) 27.3 (-0.9) 55.0 (-1.0) 36.5 (-1.1)
In this section, we validate the robustness of our calibration dataset design principles. We select three sub-tasks and respective datasets from the LongBench benchmark, including Qasper [15], MultiNews [17], and TREC [38, 28]. We use their training set to construct the calibration dataset, and use their respective test set in LongBench to calculate the score. Following Section 5, all calibration datasets are constructed using the original model’s response to the context and questions as the supervision.
As shown in Table 10, we find that as long as the calibration dataset conforms to the long-range dependency and model alignment highlighted in section 5, the specific choice of the dataset is less important. Calibration datasets with long dependency and model alignment show somewhat similar test results on various datasets. Additionally, they all show strong generalization power to test sets other than their respective calibration dataset.
In contrast, the RedPajama dataset without long-range dependency and model alignment shows large variance on various test sets. It also differs from the performance of the original dense model, which may incur unexpected behaviors after compression. Note that though all datasets exhibit long dependency, the questions in the TREC dataset can be answered without long context. The context in the TREC dataset of LongBench is the many-shot examples, each showing a short sentence and its classification result, while the question is to classify a new short sentence. Although the context helps to determine the complete set of 50 classes, the model can also directly clarify the sentence without any context based on common knowledge. It may contribute to a high score on the TREC test set with the RedPajama calibration dataset.
Appendix C Compression Plan Analysis
C.1 Statistics on Rules Discovered by MoA
Figure 8: The MoA mask density across layers for different LLMs.
Figure 9: The MoA mask’s average density and the density range for each layer for different LLMs.
This subsection provides empirical evidence for rules discovered by MoA as mentioned in Section 6.4. The lines and spans in Figure 8 show that all heads at the first few layers generally need a high KV-Cache density. Following that, a few layers generally only require medium density. Then, in the final layers, most heads require low density, while some outlier heads need high density. This observation conforms to previous findings of the intrinsic dimension of LLM [61]. The geometry of density is similar to the intrinsic dimension of LLM, with two local minima. As observed in Figure 8, layers with lower average density (smaller values on the lines) typically display a wider range of density (wider shades). Figure 9 validates such observation. This observation confirms the need for heterogeneous attention rules within the same layer.
C.2 Connections Between MoA Rule and Semantic
In this section, we invest the masks acquired with MoA and show the interpretable semantics of the masks. Previous works manually restrict the attention pattern of the model, which may harm the semantics learned by the dense model. In contrast, MoA preserves the semantics with statistic analysis and optimization. We use visualization, human interpretation and quantitive methods to analyze the semantics of the original model and to verify whether MoA captures such semantics.
C.2.1 Mask Visualization and Semantic Categorization
Given any token, two kinds of information are used as the model inputs: position encoding and token embedding. Position encoding indicates the absolute [74] or relative positions [59] of tokens in the sentence. Token embedding maps different tokens as different vectors. The attention head $h$ responds to both information and output the corresponding attention value $A_{h}$. As shown in equation 6, we denote the influence of position and token of head $h$ as function $P_{h}$ and $T_{h}$, respectively. The attention value $A_{h,i,j}$ between the $i$th and $j$th token $t_{i}$ and $t_{j}$ is determined by the combination $f_{h}$ of position and token influence functions.
$A_{h,i,j}=\mathbb{A}_{h}(t_{i},t_{j},i,j)=f_{h}\left(P_{h}(i,j),T_{h}(t_{i},t_{j})\right)$ (6)
Figure 2 visualizes two typical heads that are either dominated by position $P$ or token $T$ function. For the first attention head in Figure 2, the local positional attention is clearly observed. In this head, whatever sentences are given, each token pays major attention to the first token and the prior token. As a result, the mean attention matrix accumulates extremely large attention values at the first column and the sub-diagonal. In contrast, the second attention head in Figure 2 lays more emphasis on content-based attention. Since the position distribution of important tokens are generally random, the attention matrix can show large attention values at any position. It results in a mean attention matrix without extreme mean attention values.
In conclusion, the mean attention matrix of different sentences provides a valuable insight of whether attention values of an attention head is more position-based or content-based. Intuitively, the more uneven the attention matrix value distribution is, the more position-based the head is.
C.2.2 Quantitative Semantic Analysis
We quantify how much the attention head is position-based and analyze whether MoA successfully utilizes such semantics through the evaluate-generate-optimization pipeline. We model equation 6 with a linear approximation. $P_{h}$ and $T_{h}$ are random variables with the same expectation $\mu$ and standard variance $\delta$ for all heads. For attention head $h$, the weight factor $\alpha_{h}$ evaluates the relatively influence of position and token to the final attention value.
$A_{h,i,j}=\alpha_{h}P_{h}(i,j)+(1-\alpha_{h})T_{h}(t_{i},t_{j})$ (7)
Given the randomness of token positions in long context, we assume that the token position and its content are irrelevant. For different sentences $s$, the expectation $\mathbb{E}_{t}$ of the attention value between position $i$ and $j$ can be expressed as follows. Note that it excludes the matrix diagonal since $T_{h}(t_{i},t_{j}),i\neq j$ and $T_{h}(t_{i},t_{i})$ may follow different distributions.
$\mathbb{E}_{t}[A_{h,i,j}]$ $=\frac{1}{S}\sum_{s=1}^{S}\left(\alpha_{h}P_{h}(i,j)+(1-\alpha_{h})T_{h}(t_{i}^{(s)},t_{j}^{(s)})\right)$ (8)
$=\alpha_{h}P_{h}(i,j)+(1-\alpha_{h})\frac{1}{S}\sum_{s=1}^{S}T_{h}(t_{i}^{(s)},t_{j}^{(s)})$
$=\alpha_{h}P_{h}(i,j)+(1-\alpha_{h})\mu_{T},\forall i>j$
The standard division $\sigma_{p}$ of $\mathbb{E}_{t}$ over different positions of the attention matrix is
$\sigma_{p}(\mathbb{E}_{t}[A_{h,i,j})$ $=\sqrt{\frac{2}{(1+N)N}\sum_{i,j\in[1,N),i>j}[(\alpha_{h}P_{h}(i,j)+(1-\alpha_{h})\mu_{T})-(\alpha_{h}\mu_{P}+(1-\alpha_{h})\mu_{T})]^{2}}$ (9)
$=\alpha_{h}\delta_{p}$
Figure 10: Positive correlation between MoA’s mask sparsity and head’s dependency on position (SoE).
We name $\sigma_{p}(\mathbb{E}_{t}[A_{h,i,j}])$ the Standard division of Expectation (SoE) of head $h$. Note that the expectation is taken over different sentences, while the standard division is taken over different attention positions. Since $\delta_{p}$ is the same for all heads, we derive that the position impact $\alpha_{h}$ is proportional to the SoE of different heads.
The conclusion quantifies the observation stated in Section C.2.1. Intuitively, SoE shows how uneven the mean attention matrix is, thus showing the influence of position to the attention values. MoA’s generated mask density shows positive relation with SoE, suggesting that MoA successfully captures the semantic information of the dense language model as shown in Figure 10.
Appendix D MoA Automatic Pipeline Details
D.1 Derivation of Attention Influence
We use the first-order Taylor expansion to calculate the influence of each attention value. This approximation approach is supported by methodologies commonly employed in other LLM compression approaches [36, 55, 14, 29].
As discussed in Section 4.1, when masking out attention value $A_{h,i,j}$ at head $h$, row $i$, and column $j$, it also influences the attention values in the same row by $\Delta A_{h,i,n|j}$.
$A_{h,i,n}$ $=\frac{e^{S_{h,i,n}}}{\sum_{j}e^{S_{h,i,j}}}$ (10)
$\Delta A_{h,i,n|j}$ =$=\begin{cases}-A_{h,i,n},&n=j\\ A_{h,i,n}({\sum_{j}e^{S_{h,i,j}}}/{\sum_{j\neq n}e^{S_{h,i,j}}}-1),&n\neq j\\ \end{cases}$
Following the definition, the attention influence $\textbf{E}_{h}$ is calculated as follows:
$E_{h,i,j}=\sum_{n}\frac{\partial L}{\partial A_{h,i,n}}\cdot\Delta A_{h,i,n|j}$ (11)
Given Equation 11 and  10, we derive Equation 3 as follows. For notation simplicity, we omit the head index $h$ here.
$E_{i,j}$ $=\sum_{n}\frac{\partial L}{\partial A_{i,n}}\cdot\Delta A_{i,n|j}$ (12)
$=\frac{\partial L}{\partial A_{i,j}}\cdot(-A_{i,j})+\sum_{n\neq j}\frac{\partial L}{\partial A_{i,n}}\cdot A_{i,n}\cdot\left(\frac{\sum_{k}e^{S_{i,k}}}{\sum_{k\neq j}e^{S_{i,k}}}-1\right)$
$=\frac{\partial L}{\partial A_{i,j}}\cdot(-A_{i,j})+\sum_{n\neq j}\frac{\partial L}{\partial A_{i,n}}\cdot A_{i,n}\cdot\frac{e^{S_{i,j}}}{{\sum_{k}e^{S_{i,k}}}-e^{S_{i,j}}}$
$=\frac{\partial L}{\partial A_{i,j}}\cdot(-A_{i,j})+\sum_{n\neq j}\frac{\partial L}{\partial A_{i,n}}\cdot A_{i,n}\cdot\frac{e^{S_{i,j}}/{\sum_{k}e^{S_{i,k}}}}{1-e^{S_{i,j}}/{\sum_{k}e^{S_{i,k}}}}$
$=\frac{\partial L}{\partial A_{i,j}}\cdot(-A_{i,j})+\sum_{n\neq j}\frac{\partial L}{\partial A_{i,n}}\cdot A_{i,n}\cdot\frac{A_{i,j}}{1-A_{i,j}}$
$=\frac{\partial L}{\partial A_{i,j}}\cdot(-A_{i,j})-\frac{\partial L}{\partial A_{i,j}}\cdot A_{i,j}\cdot\frac{A_{i,j}}{1-A_{i,j}}+\sum_{n}\frac{\partial L}{\partial A_{i,n}}\cdot A_{i,n}\cdot\frac{A_{i,j}}{1-A_{i,j}}$
$=\frac{\partial L}{\partial A_{i,j}}\cdot\left(-\frac{A_{i,j}}{1-A_{i,j}}\right)+\frac{A_{i,j}}{1-A_{i,j}}\cdot\sum_{n}\frac{\partial L}{\partial A_{i,n}}\cdot A_{i,n}$
$=-\frac{A_{i,j}}{1-A_{i,j}}\left(\frac{\partial L}{\partial A_{i,j}}-\sum_{n}\frac{\partial L}{\partial A_{i,n}}\cdot A_{i,n}\right)$
It is worth noting to mention that it can also be formulated as matrix multiplications:
$\textbf{E}_{h}=\frac{\textbf{A}_{h}}{1-\textbf{A}_{h}}\cdot\left(\frac{\partial L}{\partial\textbf{A}_{h}}-\left(\frac{\partial L}{\partial\textbf{A}_{h}}\cdot\textbf{A}_{h}\right)\mathbbm{1}^{N\times N}\right).$ (13)
D.2 Optimization Details
D.2.1 Optimizing at Single Length
The optimization problem is formulated as follows:
$\mathop{\arg\min}\Delta L=\sum_{h}\Delta L_{h,r_{h}},\quad\operatorname{s.t.}\frac{1}{H}\sum_{h}d_{r_{h}}\leq d_{\text{constr}}.$ (14)
To transform the optimization problem into a standard Mixed-Integer Programming (MIP) framework, we introduce the binary variable $X_{h,r_{h}}\in\{0,1\}$. It indicates whether to select rule $r_{h}$ for the attention head $h$. Assume the model has $H$ attention head, and head $h$ has $R_{h}$ elastic rules.
$\mathop{\arg\min}\frac{1}{H}\sum_{h=0}^{H-1}\sum_{r_{h}=0}^{R_{h}-1}\Delta L_{h,r_{h}}X_{h,r_{h}}$ $\quad\operatorname{s.t.}$ (15a)
$\sum_{r_{h}=0}^{R_{h}-1}X_{h,r_{h}}=1,$ $\quad h\in\left\{0,\cdots,H-1\right\}$ (15b)
$\frac{1}{H}\sum_{h=0}^{H-1}\sum_{r_{h}=0}^{R_{h}-1}d_{r_{h}}X_{h,r_{h}}\leq d_{\text{constr}}$ (15c)
$0\leq X_{h,r_{h}}\leq 1,X_{h,r_{h}}\in\mathbb{Z},$ $\quad\forall h\in\left\{0,\cdots,H-1\right\},\forall r_{h}\in\mathbb{R}$ (15d)
In this formulation, (15a) serves as the objective function to minimize the loss, subject to the constraints that each matrix selects exactly one compression plan (15b), and the average density does not exceed $d_{\mathrm{constr}}$ (15c). Finally, (15d) enforces that $X_{h,r_{h}}$ is a binary variable, indicating the selection of plans.
Additionally, to enforce the restriction that each model layer only has a limited number of different plans, we bound the norm of element-wise multiplication of $\mathbf{X}_{h}=\begin{bmatrix}X_{h,0}&X_{h,1}&\cdots&X_{h,R_{h}-1}\end{bmatrix}^{\top}$ in a single layer.
D.2.2 Optimizing at Multiple Lengths
With the ability to optimize at a single length, we utilize the same framework for multi-objective MIP across various lengths. The key is to transform the multi-objective MIP problem into several single-objective MIP problems [47]. We utilize the idea of epsilon-constraint method [72].
We first select each input length as our primary objective to perform the single-objective optimization on it while simultaneously recording the outcomes of other objectives. Specifically, for $N$ distinct objectives, we do single-objective MIP optimization on the $i$-th objective, getting minimum loss $\Delta L^{(N_{i})}_{i}$, and we concurrently collect losses of other objectives $\Delta L^{(N_{j})}_{i}$ for $j\neq i$. This process allows us to establish the range of loss $R^{(N_{j})}=\left[\min_{i}\Delta L^{(N_{j})}_{i},\max_{i}\Delta L^{(N_{j})}_{i}\right]$ for each objective. Then, we iterate through each objective again. Compared with the original multi-objective optimization in Equation 5, we now consider other objectives as constraints. To implement this, we partition each loss range $R^{(N_{j})}$ of other objectives $j\neq i$ into $M$ uniform intervals $S^{(N_{j})}_{k}$, where $0\leq k<M$. We then solve the MIP problems for each objective $i$ and iterating through the constraint intervals:
$\mathop{\arg\min}_{r_{h}\in\mathbb{R}}\Delta L^{(N_{i})}\quad\operatorname{s.t.}\frac{1}{H}\sum_{h=1}^{H}d^{(N_{i})}_{r_{h}}\leq d^{(N_{i})}_{\text{constr}},\forall N_{i}\in\mathbb{N}_{\text{constr}};\quad\Delta L^{(N_{j})}\in S^{(N_{j})}_{k_{j}},\forall j\neq i.$ (16)
where this optimization is performed for each $i$ ranging from 00 to $N$. For each $j$, $k_{j}$ can vary independently from 00 to $M$. For efficiency consideration, we set the number of intervals as five. Finally, the results that do not conform to the Pareto front requirements are removed, resulting in the final Pareto front set of our multi-objective optimization problem.

### v2

@@
 Abstract
-Sparse attention can effectively mitigate the significant memory and throughput demands of Large Language Models (LLMs) in long contexts.
-Existing methods typically employ a uniform sparse attention mask, applying the same sparse pattern across different attention heads and input lengths.
-However, this uniform approach fails to capture the diverse attention patterns inherent in LLMs, ignoring their distinct accuracy-latency trade-offs.
-To address this challenge, we propose the Mixture of Attention (MoA), which automatically tailors distinct sparse attention configurations to different heads and layers.
-MoA constructs and navigates a search space of various attention patterns and their scaling rules relative to input sequence lengths.
-It profiles the model, evaluates potential configurations, and pinpoints the optimal sparse attention compression plan.
-MoA adapts to varying input sizes, revealing that some attention heads expand their focus to accommodate longer sequences, while other heads consistently concentrate on fixed-length local contexts.
-Experiments show that MoA increases the effective context length by 3.9×3.9$\times$3.9 × with the same average attention span, boosting retrieval accuracy by 1.57.1×1.5-7.1$\times$1.5 - 7.1 × over the uniform-attention baseline across Vicuna-7B, Vicuna-13B, and Llama3-8B models.
-Moreover, MoA narrows the capability gaps between sparse and dense models, reducing the maximum relative performance drop from $9\%-36\%$ to within $5\%$ across two long-context understanding benchmarks.
-MoA achieves a 1.21.4×1.2-1.4$\times$1.2 - 1.4 × GPU memory reduction and boosts decode throughput by $\times5$ 6.7 × for 7B and 13B dense models on a single GPU, with minimal impact on performance.
-footnotetext: Corresponding authors: Yu Wang (yu-wang@tsinghua.edu.cn), Xuefei Ning (foxdoraame@gmail.com)
+[…]
+footnotetext: Corresponding author: Xuefei Ning (foxdoraame@gmail.com)
 1 Introduction
-Figure 1: Retrieval accuracy of the Vicuna-7B model using different attention methods across varying input lengths and retrieval positions on the LongEval dataset.
-This "Needle In a Haystack" experiment [19] takes massive key-value pairs as inputs and tests the accuracy to retrieve values based on given keys from diverse positions.
+Figure 1: Retrieval accuracy of the Vicuna-7B model using different attention methods across varying input lengths and retrieval positions on the LongEval benchmark (Li et al., 2023).
+This retrieval benchmark takes massive key-value pairs as inputs and tests the accuracy to retrieve values based on given keys from diverse positions.
 (a) Original model with a full attention span; (b) StreamingLLM with half the attention span, showing reduced effectiveness beyond the span; (c) MoA with half the attention span, maintaining effectiveness beyond the span.
-Large Language Models (LLMs) exhibit remarkable versatility across numerous applications [6, 57, 63].
-Central to LLM is the attention mechanism [62], which computes interactions among tokens within a certain span, thereby enabling context understanding.
-Scaling input length is crucial for enhancing LLM capabilities [7, 60], including fact retrieval, summarization, few-shot learning, question answering and so on [4, 70].
-However, the ever-growing attention computation and Key-Value Cache (KV-Cache) pose significant efficiency challenges [54, 69, 26, 33].
+Large Language Models (LLMs) exhibit remarkable versatility across numerous applications (Brown et al., 2020; Tay et al., 2022; Wan et al., 2023).
+Central to LLM is the attention mechanism (Vaswani et al., 2017), which computes interactions among tokens within a certain span, thereby enabling context understanding.
+Scaling input length is crucial for enhancing LLM capabilities (Chen et al., 2023; Tworkowski et al., 2023), including fact retrieval, summarization, few-shot learning, question answering and so on (Bai et al., 2023; Yuan et al., 2024).
+However, the ever-growing attention computation and Key-Value Cache (KV-Cache) pose significant efficiency challenges (Sheng et al., 2023; Xiao et al., 2024c; Han et al., 2023; Kwon et al., 2023).
 Previous work proposes sparse attention methods to address the efficiency challenges of long contexts in generative LLMs.
-These methods typically employ a uniform, fixed-span sliding window mask across all heads and input lengths, limiting attention to local contexts only [69, 26].
+These methods typically employ a uniform, fixed-span sliding window mask across all heads and input lengths, limiting attention to local contexts only (Xiao et al., 2024c; Han et al., 2023).
 This approach allows the LLM to take long inputs with a fixed attention span, keeping bounded attention computation and KV caching overhead.
-Following previous works [7, 60], we quantify the effective context length as the maximum input length where content retrieval accuracy exceeds a 90% threshold.
-In principle, fixed-span local attention can gradually aggregate global information through multiple model layers, yielding a longer effective context length than each attention span [18].
-Nonetheless, we reveal that uniform masks, like StreamingLLM [69], hardly extend effective context length beyond the span, as shown in Figure 6.
+Following previous works (Chen et al., 2023; Tworkowski et al., 2023), we quantify the effective context length as the maximum input length where content retrieval accuracy exceeds a 90% threshold.
+In principle, fixed-span local attention can gradually aggregate global information through multiple model layers, yielding a longer effective context length than each attention span (Feng et al., 2022; Zaheer et al., 2020).
+Nonetheless, we reveal that uniform masks, like StreamingLLM (Xiao et al., 2024c), hardly extend effective context length beyond the span, as shown in Figure 6.
 Figure 1(b) further illustrates such limitation: with a 50% attention span mask, StreamingLLM fails to accurately retrieve content from the earlier half of the input and performs even worse at longer input lengths.
@@
 Unfortunately, the uniform approaches do not include heterogeneous rules to scale the attention spans differently for various heads.
-Besides, existing model compression methods [44, 40, 68, 36, 31, 37] use general language modeling corpora to decide the compression plan, which cannot accurately profile the influence of compression on long-context tasks.
+Besides, existing model compression methods (Men et al., 2024; Lin et al., 2023; Xiao et al., 2024b; Li et al., 2024a; Kim et al., 2023; Li et al., 2024b) use general language modeling corpora to decide the compression plan, which cannot accurately profile the influence of compression on long-context tasks.
 In this work, we propose Mixture of Attention (MoA), a training-free sparse attention method.
@@
   • Calibration Dataset Construction We emphasize the importance of data engineering in LLM compression.
-Our findings demonstrate that using datasets with long-range dependencies and referencing the original LLM’s responses are crucial for accurately profiling the influences of compression.
+Our findings demonstrate that, instead of relying on general language modeling datasets and human responses, using datasets with long-range dependencies and referencing the original LLM’s responses is essential for accurately profiling the effects of compression.
   • Automatic Optimization.
@@
 This pipeline can efficiently find the optimal plan within several hours, for example, two hours for compressing Vicuna-13B.
-Experiments show that MoA achieves $\times5$ to 6.7×6.7$\times$6.7 × throughput improvements on 7B and 13B dense LLMs at a 50% density (the average of KV-Cache length / input length), with only 1% average relative degradation in retrieval accuracy.
+Experiments show that MoA achieves $\times6$ 8.2 × throughput improvements over FlashAttention2, 1.71.9×1.7-1.9$\times$1.7 - 1.9 × over vLLM framework on 7B and 13B dense LLMs at a 50% density (the average of KV-Cache length / input length), with only a 1% average relative degradation in retrieval accuracy.
+In Section 6.4, we show that this significant throughput improvements of MoA over FlashAttention2 can be attributed to four factors: (1) the static size of the KV-cache, (2) reduced attention computations, (3) increased batch size enabled by reduced memory usage, and (4) a specialized kernel implementation.
 Additionally, MoA achieves over 90% retrieval accuracy with just 25% average density, far surpassing sparse attention baselines that need a density of 75% to 100% for similar performance.
@@
 2.1 Attention mechanism
-The Multi-Head Self Attention (MHA) mechanism [62] is crucial to the functionality of LLMs.
+The Multi-Head Self Attention (MHA) mechanism (Vaswani et al., 2017) is crucial to the functionality of LLMs.
 It starts with an input sequence transformed into query (Q), key (K), and value (V) matrices through linear projections.
@@
 Efficient methods are proposed to mitigate the computation and memory costs associated with attention.
-One branch of work uses dynamic sparse attention masks to adaptively skip attention computations during prefill stage [46, 52, 53, 64, 43, 32] or drop KV-Cache during decode stage [3, 75, 20, 54, 41] based on the input sequences.
-However, due to the complex control and computation flow, dynamic prefill often requires specific hardware to achieve substantial wall-time speedup [52, 64, 43, 25, 24].
-Additionally, dynamic KV-Cache pruning in the decode stage may require extensive retraining [3], or additional accumulated attention score computation [54, 75, 41, 20].
+One branch of work uses dynamic sparse attention masks to adaptively skip attention computations during prefill stage (Pagliardini et al., 2023; Qu et al., 2022; Roy et al., 2021; Wang et al., 2021; Lu et al., 2021; Kitaev et al., 2020) or drop KV-Cache during decode stage (Anagnostidis et al., 2023; Zhang et al., 2023; Ge et al., 2023; Sheng et al., 2023; Liu et al., 2023a) based on the input sequences.
+However, due to the complex control and computation flow, dynamic prefill often requires specific hardware to achieve substantial wall-time speedup (Qu et al., 2022; Wang et al., 2021; Lu et al., 2021; Ham et al., 2021; 2020).
+Additionally, dynamic KV-Cache pruning in the decode stage may require extensive retraining (Anagnostidis et al., 2023), additional accumulated attention score computation (Sheng et al., 2023; Zhang et al., 2023; Liu et al., 2023a; Ge et al., 2023), or extensive memory swap for KV-Cache retrieval (Tang et al., 2024; Xiao et al., 2024a).
 Another branch of work uses static sparse attention, where predefined masks are applied consistently across all processed sentences.
 Thanks to the fixed computation flow, static sparse attention is generally more efficient and GPU-friendly.
-For language understanding models such as BERT [16], various masks are used [73, 5, 9, 76, 69, 26].
-But for generative LLMs, the predominant method is the fixed-span sliding window mask with global attention on a few initial tokens [69, 26].
+For language understanding models such as BERT (Devlin et al., 2018), various masks are used (Zaheer et al., 2020; Beltagy et al., 2020; Child et al., 2019; Zhou et al., 2024; Xiao et al., 2024c; Han et al., 2023).
+But for generative LLMs, the predominant method is the fixed-span sliding window mask with global attention on a few initial tokens (Xiao et al., 2024c; Han et al., 2023).
 With the local attention pattern, the KV-Cache beyond the current attention span can be dropped, saving much memory for long sequence scenarios.
@@
 Our method falls within this category, benefiting from the efficiency and training-free advantages, while addressing the performance limitations encountered by previous methods.
-In addition to sparse attention, alternative mechanisms have been proposed to replace traditional attention for long-sequence modeling [21, 49, 56, 51, 39, 30, 50, 10, 65].
+In addition to sparse attention, alternative mechanisms have been proposed to replace traditional attention for long-sequence modeling (Gu & Dao, 2023; Peng et al., 2023; Sun et al., 2023; Poli et al., 2023; Li et al., 2022; Kacham et al., 2023; Peng et al., 2021; Choromanski et al., 2020; Wang et al., 2020).
 However, these new mechanisms often require different weights compared to vanilla transformers, imposing significant re-training overhead for LLMs.
-Previous works also propose LLM acceleration frameworks [22, 2, 54, 33], as well as kernel-level optimizations [13, 12].
+Previous works also propose LLM acceleration frameworks (Gugger et al., 2022; Aminabadi et al., 2022; Sheng et al., 2023; Kwon et al., 2023), as well as kernel-level optimizations (Dao et al., 2022; Dao, 2023; Shah et al., 2024).
 These kernel and system optimizations are orthogonal to our work and can be integrated to further enhance efficiency.
@@
 17,18,19 0.97 0.94 0.97
-Table 1: Retrieval accuracy of Vicuna-7B model using sliding-window sparse attention masks across various model layers, window spans, and input lengths.
+Table 1: Retrieval accuracy of Vicuna-7B with sliding-window sparse attention across various model layers, window spans, and input lengths.
 3.1 Mixture of Attention Patterns and Elastic Rules
@@
 For example, the first head primarily focuses on local contexts with a narrow-span sliding window, while the third head covers nearly the entire input, indicating global attention.
+The attention spans of different heads mostly remain constant across various tasks and datasets, as shown in Appendix D.1.
 Table 1 demonstrates that applying the same sliding-window sparse attention mask across model layers can lead to a 65% variance in retrieval accuracies.
-It conforms to the multi-head self-attention design principle of capturing varied information [62], as well as the findings from concurrent research that identifies specific attention heads for global text retrieval [67].
+It conforms to the multi-head self-attention design principle of capturing varied information (Vaswani et al., 2017), as well as the findings from concurrent research that identifies specific attention heads for global text retrieval (Wu et al., 2024).
 Heterogeneous Elastic Rules.
@@
 In designing the search space for the MoA mask, we consider the inherently heterogeneous and elastic nature of LLM attention patterns.
-As shown in Figure 3(a), we adopt a hardware-friendly sliding-window mask as our base sparse attention mask [5].
-Following previous work [69, 26], the initial few tokens (64 tokens for MoA) are not masked.
+As shown in Figure 3(a), we adopt a hardware-friendly sliding-window mask as our base sparse attention mask (Beltagy et al., 2020).
+Following previous work (Xiao et al., 2024c; Han et al., 2023), the initial few tokens (64 tokens for MoA) are not masked.
 The attention span equals the sliding-window-span plus the number of initial unmasked tokens.
-For simplicity, we define the attention span S of head $h$ at input length N using a straightforward linear function:
-$S_{h}=\alpha_{h}+\beta_{h}\times N,$ (2)
-where $\alpha_{h}$ and $\beta_{h}$ are hyperparameters that control the base span and its expansion rate with input length.
-To ensure hardware-friendliness, we limit the number of distinct rules to at most two per model layer for MoA unless specified otherwise.
+We define the attention span S of head $h$ at input length N using a straightforward linear function:
+$S_{h}=\alpha_{h}+\beta_{h}\cdot N,$ (2)
+where $\alpha_{h}$ and $\beta_{h}$ are hyperparameters that control the base span and its expansion rate with input length of a specific attention head.
 The $\alpha$ and $\beta$ hyperparameters of each attention head can be chosen from multiple discrete options.
@@
 The average attention influence is calculated respectively for different input lengths.
-The gradient ${\partial L}/{\partial\textbf{A}_{h}}$ is computed using chain derivative in deep learning frameworks like PyTorch [48].
+The gradient ${\partial L}/{\partial\textbf{A}_{h}}$ is computed using chain derivative in deep learning frameworks like PyTorch (Paszke et al., 2019).
 The detailed calibration dataset setup is discussed in Section 5.
@@
 The loss is calculated with the sum of masked average attention influence according to the rule.
-We denote $\textbf{M}_{r_{h}}$ as the binary mask at head $h$ that corresponds to rule $r$, with masked positions marked as 0 and others as 1.
+We denote $\textbf{M}_{r_{h}}$ as the binary mask at head $h$ that corresponds to rule $r$, with masked positions marked as 1 and others as 0.
 We formalize accuracy loss $\Delta L$ as follows:
-$\Delta L=\sum_{h}\Delta L_{h,r_{h}}=\sum_{h}\sum_{i}\sum_{j}M_{r_{h},i,j}\bar{E}_{h,i,j}.$ (4)
+$\Delta L=\sum_{h}\Delta L_{h,r_{h}}=\sum_{h}\sum_{i}\sum_{j}M_{r_{h},i,j}\cdot\bar{E}_{h,i,j}.$ (4)
 After the profile stage, MoA acquires the unique accuracy-density trade-offs of elastic rules.
@@
 Here, superscript $(N)$ denotes values at different lengths; $\mathbb{N}_{\text{profile}}$ and $\mathbb{N}_{\text{constr}}$ denote the sets of lengths for profiling and those subject to density constraints, respectively; $\mathbb{R}$ denotes the set of candidate rules; $\Delta L^{(N_{i})}$ denotes the accuracy loss due to compression; $d_{r_{h}}^{(N_{i})}$ denotes the density of rule $r_{h}$ at head $h$; $d_{\text{constr}}^{(N_{i})}$ denotes the average density constraint; $H$ denotes the total number of attention heads.
-Such formulation corresponds to the classic multi-objective mixed-integer-planing problem, which can be effectively solved within minutes using existing linear solvers, like Gurobi [23].
+Such formulation corresponds to the classic multi-objective mixed-integer-planing problem, which can be effectively solved within minutes using existing linear solvers, like Gurobi (Gurobi Optimization, LLC, 2023).
 The detailed problem formulation and solving strategies are discussed in Appendix D.3.
@@
 In this section, we highlight the overlooked importance of calibration dataset design and its supervision objective in LLM compression.
-Calibration datasets are essential for sensitivity analysis across various compression techniques, including weight pruning [44, 34, 42] and quantization [40, 68, 36, 31].
+Calibration datasets are essential for sensitivity analysis across various compression techniques, including weight pruning (Men et al., 2024; Lee et al., 2024; Liu et al., 2023b) and quantization (Lin et al., 2023; Xiao et al., 2024b; Li et al., 2024a; Kim et al., 2023).
 In this work, MoA profiles the attention influence on the calibration dataset, which is crucial for subsequent automatic optimization.
 Current Approach.
-General language modeling datasets, such as human-written text corpus RedPajama [11], are commonly used as the calibration dataset.
+General language modeling datasets, such as human-written text corpus RedPajama (Computer, 2023), are commonly used as the calibration dataset.
 These datasets, supervised by next-token-prediction on the entire corpus, primarily capture attention patterns coherent with immediately preceding tokens.
@@
 Consequently, it leads to inaccuracies when using human responses to compute attention values and gradients during profiling.
-For example, given the same question, a human might answer ’Blue’, while the model could generate ’The blue color’.
-Using the human answer for supervision, attention influence is inaccurately quantified based on probability shift for predicting ’Blue’; this diverges from the objective of maintaining crucial attention for the original model prediction, ’The’.
+For example, given the same question, a human might answer ‘Blue’, while the model could generate ‘The blue color’.
+Using the human answer for supervision, attention influence is inaccurately quantified based on probability shift for predicting ‘Blue’; this diverges from the objective of maintaining crucial attention for the original model prediction, ‘The’.
 These inconsistencies arise from various factors, including mismatched positions, tones, and synonyms.
@@
 MoA enhances the calibration dataset by integrating long-range dependencies and model alignment.
-It utilizes the long-contextual MultiNews dataset [17], which includes summaries heavily dependent on long-range content.
+Specifically, we utilize the long-contextual MultiNews dataset (Fabbri et al., 2019), which includes summaries heavily dependent on long-range content.
 The summaries are generated by the original dense model and serve as supervision.
@@
 We show the importance of long-range dependencies by comparing the MoA compression plan generated with different datasets and supervisory methods.
-In Table 2, RedPajama [11] represents the general language modeling dataset, while MultiNews [17] highlights long-range contexts by aggregating multiple documents on a single incident.
+In Table 2, RedPajama (Computer, 2023) represents the general language modeling dataset, while MultiNews (Fabbri et al., 2019) highlights long-range contexts by aggregating multiple documents on a single incident.
 Additionally, each MultiNews item includes a human-written summary, providing even stronger long-range dependencies and better performance.
@@
 Baselines.
-We compare MoA with state-of-the-art static and dynamic sparse attention baselines for LLMs: StreamingLLM [69] and H2O [75].
-We define the density of an LLM as the average KV-Cache length divided by the sequence length during the sparse decode stage.
-Notably, in MoA and StreamingLLM, the KV-Cache length matches the attention span during their sparse prefill stages.
-In contrast, H2O employs a dense prefill and requires additional computations for dynamic mask generation based on accumulative attention scores.
+We compare MoA with state-of-the-art static and dynamic sparse attention methods, including StreamingLLM (Xiao et al., 2024c), InfLLM (Xiao et al., 2024a) and H2O (Zhang et al., 2023).
+We define the density of an LLM as the ratio of the average in-memory KV-Cache length to the sequence length during the sparse decode stage.
+Notably, in MoA and StreamingLLM, KV-Cache length equals the attention span during the sparse prefill stage.
+In contrast, H2O use dense prefill.
+Besides, H2O and InfLLM require additional computations to dynamically determine the KV-Cache.
 Models and Benchmarks.
-We evaluate on vicuna-7b-v1.5-16k, vicuna-13b-v1.5-16k [8] from LMSys, and Llama-3-8B-Instruct-262k [1] from Gradient AI.
-For effective context length evaluation, we use LongEval [35] to test key-value retrieval accuracy with 100 data items per length level.
-For comprehensive ability evaluation, we use LV-Eval [70] and LongBench [4], which include 11 and 13 sub-tasks, respectively.
-For coherence testing, we measure perplexity on four subsets of LongBench [15, 17, 38, 28, 45], representing different capabilities.
+We evaluate on vicuna-{7b, 13b}-v1.5-16k models (Chiang et al., 2023) from LMSys and Llama-3-{8b, 70b}-Instruct-262k models (AI, 2024) from Gradient AI.
+For long-context retrieval, we use LongEval (Li et al., 2023) to test key-value retrieval accuracy with 100 data items per length level.
+For long-context understanding, we use LV-Eval (Yuan et al., 2024) and LongBench (Bai et al., 2023), which include 11 and 13 sub-datasets, respectively.
+For coherence testing, we measure perplexity on four long datasets (Dasigi et al., 2021; Fabbri et al., 2019; Li & Roth, 2002; Hovy et al., 2001; Mohler et al., 2016) with diverse tasks.
+Unless otherwise specified, performance experiments are restricted to eight A100-80GB GPUs over a 24-hour period, with OOM (Out-Of-Memory) and OOT (Out-Of-Time) conditions noted.
+Efficiency experiments measure the decode throughput on a single A100-80GB GPU at maximum batch sizes of respective methods.
 MoA Settings.
-We profile MoA on MultiNews [17] with model summaries at 2k, 4k, and 8k lengths.
+We restrict the number of distinct rules to at most two per model layer to ensure inference-time efficiency.
+We profile MoA on MultiNews (Fabbri et al., 2019) with model summaries at 2k, 4k, and 8k lengths.
 The optimal compression plan is selected with the validation dataset at 12k.
-Unless specified, the same plan is used across all benchmarks and lengths for each model.
-The models are not finetuned.
+Each model uses the same plan across all benchmarks and lengths.
+The models are not fine-tuned.
+6.2 Accuracy-Throughput Trade-off
+Figure 4 shows that MoA advances the Pareto Front in context retrieval accuracy and decode throughput across varied densities.
+At the same densities, MoA notably enhances throughput by 1.6×1.6$\times$1.6 × to 18.1×18.1$\times$18.1 × compared to H2O and InfLLM, due to its efficient static attention design.
+The throughput even outperforms StreamingLLM thanks to our customized GPU kernel.
+Additionally, MoA achieves notably higher retrieval accuracies across all densities.
+We conduct extensive evaluations for MoA ’s performance and efficiency on various benchmarks across context lengths from 4k to 256k and model sizes ranging from 7B to 70B in subsequent sections.
 6.3 Performance
-Table 3: Comparative analysis of retrieval accuracy, LV-Eval scores, LongBench scores, and perplexity for various models with different attention methods.
-All methods employ 50% density in both prefill and decode stages, except for H2O, which uses dense prefill.
-Retrieve Acc. $\uparrow$ LV-Eval $\uparrow$ LongBench $\uparrow$ PPL $\downarrow$
-Model Attention 4k 8k 16k 16k 0-4k 4-8k 8-16k 8-12k
-Vicuna-7B Original 1.00 0.98 0.62 5.93 37.91 33.82 32.54 3.79
-H2O 0.86 0.68 0.35 5.42 36.23 32.74 31.81 3.94
-StreamingLLM 0.43 0.16 0.08 4.72 30.53 33.28 31.70 4.48
-MoA 1.00 0.97 0.57 5.61 37.04 32.90 31.94 3.75
-Vicuna-13B Original 0.99 0.98 0.44 5.83 42.25 39.52 35.93 3.62
-H2O 0.88 0.76 0.28 5.66 41.63 38.02 34.75 3.80
-StreamingLLM 0.65 0.49 0.33 5.43 30.65 33.07 32.68 4.10
-MoA 0.99 0.93 0.49 7.16 41.73 38.88 35.69 3.62
-Llama3-8B Original 0.99 0.99 0.97 17.49 44.27 43.53 43.26 4.52
-H2O 0.94 0.89 0.88 16.03 43.46 43.01 42.50 4.63
-StreamingLLM 0.68 0.55 0.52 11.16 37.20 38.02 39.43 4.79
-MoA 0.99 1.00 1.00 17.46 43.07 42.75 43.09 4.49
-Overall Performance.
-As detailed in Table 3, MoA outperforms state-of-the-art sparse attention methods and achieves comparable performance to the original dense model at 50% density.
-We calculate the relative performance drop of the compressed model over the original one.
-For retrieval tasks, MoA demonstrates a maximum of 8% relative accuracy drop (calculated as $\max\{1-\text{Acc.}_{\text{MoA}}/\text{Acc.}_{\text{Original}}\}$ across three lengths and LLMs), significantly less than the 87% and 44% for StreamingLLM and H2O.
-On average, the relative accuracy drop for MoA is under 1%, while it is much higher for StreamingLLM and H2O at 51% and 20%, respectively.
-In benchmarks using LV-Eval and LongBench, MoA shows only a 5% and 3% maximum relative drop in scores, respectively, compared to steep declines for StreamingLLM at 36% and 27%.
-H2O exhibits maximum drops of 9% and 4% with higher efficiency costs.
-Similar trends show in perplexity tests, where MoA maintains less than 1% relative perplexity increase, while StreamingLLM and H2O exhibit 13% and 4% increases.
-This trend holds for other densities, as shown in Appendices B.1.1 and B.1.3.
-Figure 4: Retrieval accuracy tests on LongEval.
-With Vicuna-7B: (a) varies input lengths and densities to show effective context lengths across attention spans, (b) sets input length at 8k and shows retrieval accuracy across different densities.
-With Llama3-8B: (c) sets density at 50% and shows retrieval accuracy across increasing input lengths.
-Long-Context Retrieval.
-We investigate the ultimate effective context length with retrieval tests on LongEval [35].
-As shown in Figure 4(a), MoA expands its effective context length to approximately 3.9×3.9$\times$3.9 × its attention span before reaching up to the 12k limit of the original model.
-Detailed data for effective context length calculation are in Appendix B.1.2.
-Figure 4(b) shows that at a fixed input length of 8k, MoA reaches over 0.9 retrieval accuracy with just 25% density, whereas StreamingLLM and H2O require 100% and 75% density, respectively.
-Figure 4(c) further investigates MoA’s ability to expand and generalize to extreme input lengths.
-By profiling within 8k lengths, MoA extends its capability to 60k inputs while maintaining over 90% retrieval accuracy at 50% density.
-Figure 5: (a) LV-Eval and (b) LongBench scores for different attention methods at 50% density, tested on Vicuna-7B and 13B models. Scores normalized against the original dense model.
-Long-Context Understanding.
-Figure 5 illustrates the performance comparison across long-context benchmarks LV-Eval and LongBench.
-MoA achieves comprehensive performance comparable to the original dense model, as well as H2O that requires higher efficiency cost.
-In contrast, StreamingLLM displays inconsistent performance: although it occasionally surpasses the original model, it generally suffers noticeable degradation.
-Ablation Study.
-We evaluate the performance impact of different sparse mask search spaces in Table 5.
-Starting with a basic uniform mask, we observe significant enhancements by sequentially introducing heterogeneity: layers first, then heads, and finally elastic rules.
-6.3 Efficiency
-Table 4: Efficiency analysis of different frameworks on 7B and 13B models.
-H2O and MoA use 50% density.
-GPU memory evaluated with batch sizes 8 (7B model) and 4 (13B model).
-Decode throughput evaluated at the maximum batch capacity of an A100-80GB GPU.
-Memory (GB) Throughput (tok/s)
-Size Framework 4k 8k 16k 4k 8k 16k
-7B FlashAttn2 28.5 44.4 76.3 134.6 66.9 32.9
-H2O 36.9 OOM OOM 754.9 296.3 51.7
-MoA 22.7 32.9 53.5 897.7 436.7 206.4
-13B FlashAttn2 36.8 49.2 74.0 81.3 40.8 19.8
-H2O 40.4 77.9 OOM 330.2 138.2 37.4
-MoA 32.0 39.6 55.0 471.5 222.6 108.3
-Table 5: Ablation study on search space with consistent 25% density, progressively introducing heterogeneity in layers, heads, and elastic rules.
-Evaluations are done with retrieval accuracy and perplexity.
+Figure 4: Accuracy-throughput trade-offs of various attention methods at different densities, tested on Vicuna-7B with 8k input length using one A100-80GB GPU on the LongEval dataset.
 Retrieval Acc. PPL
@@
 +Elastic 0.98 0.43 3.96 4.29
-We use Huggingface [66] with FlashAttention [13] as the baseline and implement MoA based on it.
-As shown in Table 5, with an average 50% density, MoA reduces total GPU memory by 1.2×1.2$\times$1.2 × to 1.4×1.4$\times$1.4 × across various model sizes and input lengths.
-Moreover, MoA boosts decode throughput by $\times5$ to 6.7×6.7$\times$6.7 ×, primarily attributed to a static-size KV-Cache (3.0$\approx$ 3.0$\times$ 3.0 ×), reduced attention computations (1.5$\approx$ 1.5$\times$ 1.5 ×) and increased batch sizes (1.4$\approx$ 1.4$\times$ 1.4 ×).
-Compared with H2O that uses dense prefill and requires additional cost for dynamic mask generation, MoA achieves $1.2$ to $4.0$ throughput at the same 50% density.
-Compared to the highly optimized vLLM framework [33], MoA still achieves a 1.4×1.4$\times$1.4 × to 1.5×1.5$\times$1.5 × increase in throughput.
-More efficiency results are shown in Appendix B.2.
+Table 3: Ablation study on search space with consistent 25% density, progressively introducing heterogeneity in layers, heads, and elastic rules.
+Evaluations are done with retrieval accuracy and perplexity.
+Table 4: Comparative analysis of retrieval accuracy, LV-Eval scores, LongBench scores, and perplexity for various models with different attention methods.
+All sparse methods employ 50% density in decode stage.
+H2O uses dense prefill, while StreamingLLM, InfLLM and MoA use sparse prefill.
+InfLLM for 70B model is excluded due to OOT issues.
+Retrieve Acc. $\uparrow$ LV-Eval $\uparrow$ LongBench $\uparrow$ PPL $\downarrow$
+Model Attention 4k 8k 16k 16k 0-16k 8-12k
+Vicuna-7B Original 1.00 0.98 0.62 5.93 34.76 3.79
+H2O 0.86 0.68 0.35 5.42 33.59 3.94
+InfLLM 0.67 0.57 0.26 5.13 32.97 4.07
+StreamingLLM 0.43 0.16 0.08 4.72 31.84 4.48
+MoA 1.00 0.97 0.57 5.61 33.96 3.75
+Vicuna-13B Original 0.99 0.98 0.44 5.83 39.23 3.62
+H2O 0.88 0.76 0.28 5.66 38.13 3.80
+InfLLM 0.70 0.53 0.27 6.80 37.13 4.07
+StreamingLLM 0.65 0.49 0.33 5.43 32.13 4.10
+MoA 0.99 0.93 0.49 7.16 38.77 3.62
+Llama3-8B Original 0.99 0.99 0.97 17.49 43.69 4.52
+H2O 0.94 0.89 0.88 16.03 42.99 4.63
+InfLLM 0.65 0.59 0.37 14.44 42.43 4.68
+StreamingLLM 0.68 0.55 0.52 11.16 38.22 4.79
+MoA 0.99 1.00 1.00 17.46 42.97 4.49
+Llama3-70B Original 1.00 0.99 0.93 24.51 49.10 3.67
+H2O 0.93 0.91 OOM OOM OOM OOM
+StreamingLLM 0.20 0.15 0.04 17.45 42.53 4.26
+MoA 1.00 1.00 0.94 23.65 47.79 3.75
+MoA outperforms state-of-the-art sparse attention methods across various model sizes and benchmarks, achieving comparable performance to the original dense model at 50% density.
+Long-Context Retrieval.
+As shown in Table 4, MoA demonstrates a maximum of 8% relative accuracy drop (calculated as $\max\{1-\text{Acc.}_{\text{MoA}}/\text{Acc.}_{\text{Original}}\}$ across three lengths and LLMs), significantly less than the 87%, 58% and 44% for StreamingLLM, InfLLM and H2O.
+On average, the relative accuracy drop for MoA is under 1%, much less than others at 51%, 41% and 20%, respectively.
+Figure 5(a) shows that MoA retains over 90% retrieval accuracy up to 60k lengths, equaling the dense model’s effective context length.
+Note that it is done within 8k profiling and 12k validation.
+In contrast, the effective context lengths for H2O, InfLLM, and StreamingLLM are only 8k, <4k, and <4k, respectively.
+Appendix B.1.2 shows that MoA extends its effective context to approximately 3.9×3.9$\times$3.9 × the average KV-Cache length.
+Long-Context Understanding.
+As shown in Table 4, MoA minimizes the maximum relative performance drop in LV-Eval and LongBench benchmarks to only 5% and 3%, respectively—much lower than the 36% and 18% experienced by StreamingLLM.
+H2O and InfLLM show maximum relative drops of 9%-17% and 3%-5% with higher efficiency costs.
+Similar trends show in perplexity tests, where MoA maintains less than 1% relative perplexity increase, while others exhibit 4%-13% increases.
+This trend holds for other densities, as shown in Appendices B.1.1 and B.1.3.
+Figure 9 and Table 8 further details the score with different tasks.
+MoA achieves comprehensive performance comparable to the original dense model, as well as H2O that requires higher efficiency cost.
+In contrast, StreamingLLM and InfLLM display inconsistent performance: it sometimes surpasses the original model in some tasks, while suffering noticeable degradation in others.
+Longer-Context Generalization.
+By compressing within 12k lengths, MoA effectively generalizes to lengths of 32k-256k, as shown in Figure 5(b).
+At the extended lengths, MoA outperforms both InfLLM and StreamingLLM by 1.93.3×1.9-3.3$\times$1.9 - 3.3 × in retrieval accuracy and 1.21.4×1.2-1.4$\times$1.2 - 1.4 × in LV-Eval scores, demonstrating comparable performance to the original dense model.
+(a) Retrieval accuracy and the effective context length (arrow).
+Retrieve Acc. $\uparrow$ LV-Eval $\uparrow$
+Attention 32k 64k 128k 256k 32k 64k 128k
+Original 0.98 0.93 0.76 0.37 16.74 15.39 14.71
+InfLLM 0.43 0.32 0.25 OOT 14.22 12.17 OOT
+StreamingLLM 0.52 0.48 0.41 0.25 12.38 11.45 11.94
+MoA 1.00 0.92 0.83 0.46 17.07 15.13 14.14
+(b) Retrieval accuracy and LV-Eval score at longer lengths
+Figure 5: Comparative analysis at extended sequence lengths with different attention methods using Llama3-8B model. All methods employ 50% density in both prefill and decode stages.
+Ablation Study.
+We evaluate the performance impact of different sparse mask search spaces in Table 3.
+Starting with a basic uniform mask, we observe significant enhancements by sequentially introducing heterogeneity: layers first, then heads, and finally elastic rules.
+6.4 Efficiency
+Table 5: Runtime efficiency of different methods on Vicuna-7B and 13B models.
+Efficiency improvements of MoA are ablated with four factors.
+All sparse attention methods use 50% density.
+Decode throughput (tokens per second) evaluated at the maximum batch capacity of an A100-80GB GPU.
+4k 8k 16k
+Model Framework Attention Batch Throughput Batch Throughput Batch Throughput
+7B vLLM PagedAttention 30 628.8 15 323.0 8 145.5
+FlexGen H2O 20 754.9 6 296.3 1 51.7
+HuggingFace InfLLM 15 62.0 10 37.5 6 19.2
+HuggingFace StreamingLLM 50 945.1 25 467.3 12 232.0
+HuggingFace FlashAttention2 30 134.6 15 66.9 8 32.9
++Static KV-Cache 30 496.1 15 219.5 8 91.6
++Reduced Attention 30 722.5 15 369.9 8 178.3
++Increased Batch 50 897.7 25 436.7 12 206.4
++Kernel (=MoA) 50 1099.0 25 535.7 12 257.3
+13B vLLM PagedAttention 16 314.8 8 160.5 4 71.1
+FlexGen H2O 12 330.2 4 138.2 1 37.4
+HuggingFace InfLLM 8 30.3 5 17.63 3 11.3
+HuggingFace StreamingLLM 28 478.4 14 241.2 7 116.5
+HuggingFace FlashAttention2 16 81.3 8 40.8 4 19.8
++Static KV-Cache 16 264.6 8 111.3 4 62.2
++Reduced Attention 16 329.6 8 156.4 4 87.3
++Increased Batch 28 471.5 14 222.6 7 108.3
++Kernel (=MoA) 28 550.9 14 267.6 7 132.3
+MoA shows high runtime efficiency with a manageable one-time compression overhead.
+Runtime Efficiency.
+Table 5 compares the runtime efficiency of MoA over various attention methods and LLM frameworks, with the ablation of efficiency improvements brought by each design factor of MoA.
+At 50% density, MoA boosts the decode throughput by 6.6$\times$ to 8.2$\times$ compared to FlashAttention2.
+It outperforms H2O and InfLLM with 1.2×1.2$\times$1.2 × to 4.0×4.0$\times$4.0 × decode throughput improvements.
+Even compared to the highly system-level optimized vLLM framework (Kwon et al., 2023), MoA still achieves a 1.7×1.7$\times$1.7 × to 1.9×1.9$\times$1.9 × throughput increase.
+MoA also reduces total GPU memory by 1.2×1.2$\times$1.2 × to 1.4×1.4$\times$1.4 ×, are detailed in Appendix B.2.
+This throughput gain results from four main factors: static-sized KV-Cache during generation(3.0$\approx$ 3.0$\times$ 3.0 ×); reduced attention computations due to sparsity (1.5$\approx$ 1.5$\times$ 1.5 ×); increased batch sizes enabled by smaller KV-Cache memory (1.4$\approx$ 1.4$\times$ 1.4 ×); and our CUDA-implemented GPU kernel for MoA heterogeneous attention (1.2$\approx$ 1.2$\times$ 1.2 ×).
+Compression Pipeline Efficiency.
+MoA completes the automatic compression pipeline for the Vicuna-7B and 13B models within two hours.
+For the larger Llama3-70B model, the process requires 8.5 hours of real-time and 34.7 hours of GPU time.
+Details for each compression step overhead are in Appendix B.2.2.
 6.5 Rules Discovered by MoA
 We investigate MoA’s elastic rules for each head.
-As shown in Figure 8, masks in the initial and middle layers exhibit high density, aligning with the conclusions from previous research on LLM’s intrinsic dimensions [61] and layer sensitivities [71].
+As shown in Figure 10, masks in the initial and middle layers exhibit high density, aligning with the conclusions from previous research on LLM’s intrinsic dimensions (Valeriani et al., 2023) and layer sensitivities (Yuan et al., 2023).
 Conversely, in the final layers, most heads require low density, while few need high density.
@@
 MoA automates the selection of heterogeneous elastic masks for each attention head and input length, significantly extending the effective context length of LLMs by 3.9×3.9$\times$3.9 ×.
-It enhances retrieval accuracy by 1.5×1.5$\times$1.5 × to 7.1×7.1$\times$7.1 × and increases throughput to over 5×5$\times$5 × at 50% average density, maintaining performance on par with dense models in rigorous benchmarks.
+It enhances retrieval accuracy by 1.5×1.5$\times$1.5 × to 7.1×7.1$\times$7.1 × over uniform sparse attention method and increases throughput to over 6×6$\times$6 × at 50% average density, maintaining performance on par with dense models in rigorous benchmarks.
 Limitations and Future Work.
@@
 Designing a dynamic MoA method has the potential to address this issue, which we leave for future work.
-To further enhance MoA’s efficiency, system, and kernel-level optimizations such as kernel fusion [27] and KV-Cache management [33] could be integrated.
 Using non-linear elastic rules with bounded attention spans is also worth exploring.
-Additionally, MoA’s profiling method can be adapted to evaluate the influence of weights and other activations, facilitating other compression methods such as weight and activation quantization [40, 68, 37].
+Additionally, MoA’s profiling method can be adapted to evaluate the influence of weights and other activations, facilitating other compression methods such as weight and activation quantization (Lin et al., 2023; Xiao et al., 2024b; Li et al., 2024b).
 Acknowledgement
@@
 In the setup for our experiment, we adhere to specific configurations outlined in the respective papers.
-In the case of StreamingLLM [69], the initial four tokens remain unmasked, serving as the attention sink.
-For H2O [75], we ensure the same number of heavy hitter tokens and recent tokens.
+In the case of In StreamingLLM (Xiao et al., 2024c), the initial four tokens remain unmasked, serving as the attention sink, except for the 70b model in Table 4 and the super long setting in Figure 5, where we use 64 tokens as the attention sink.
+For InfLLM (Xiao et al., 2024a), we adhere to the original configuration by maintaining the same local window size and selected memory size, using 128 initial tokens as specified in their setup.
+For H2O (Zhang et al., 2023), we ensure the same number of heavy hitter tokens and recent tokens.
 Note that H2O uses dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction.
-StreamingLLM and MoA use sparse prefill.
+StreamingLLM, InfLLM and MoA use sparse prefill.
 Models and Benchmarks.
-Since vicuna-7b-v1.5-16k and vicuna-13b-v1.5-16k [8] can only take in 16k context length, we use the 16k split of LV-Eval benchmark [70], truncating the input to 15500 for model input.
-For the LongBench benchmark [4], we use the LongBench-E split, which features a balanced number of data items at every length level.
+Since vicuna-7b-v1.5-16k and vicuna-13b-v1.5-16k (Chiang et al., 2023) can only take in 16k context length, we use the 16k split of LV-Eval benchmark (Yuan et al., 2024), truncating the input to 15500 for model input in Table 4.
+For the LongBench benchmark (Bai et al., 2023), we use the LongBench-E split, which features a balanced number of data items at every length level.
 The LongBench dataset is segmented into ranges of 0-4k, 4-8k, and 8k+ tokens.
@@
 Perplexity Evaluation.
-We construct a comprehensive yet concise test set by sampling $50\times 4$ data items for each length level from the test split of four long-context understanding datasets: Qasper [15], MultiNew  [17], TREC [38, 28] and LCC [45], representing the question answering, summarization, few-shot learning, and code completion abilities of the LLM.
+We construct a comprehensive yet concise test set by sampling $50\times 4$ data items for each length level from the test split of four long-context understanding datasets: Qasper (Dasigi et al., 2021), MultiNew  (Fabbri et al., 2019), TREC (Li & Roth, 2002; Hovy et al., 2001) and LCC (Mohler et al., 2016), representing the question answering, summarization, few-shot learning, and code completion abilities of the LLM.
 Following LongBench, the data items are organized as question-answer pairs.
@@
 The validation dataset is similarly constructed as the perplexity test dataset, but on the respective validation split of the datasets.
-$50\times 4$ data items are sampled from the same four long-context understanding datasets: Qasper [15], MultiNew  [17], TREC [38, 28] and LCC [45].
-The additional $50$ data items from the LongEval [35] dataset are also added to validate the retrieval ability.
+$50\times 4$ data items are sampled from the same four long-context understanding datasets: Qasper (Dasigi et al., 2021), MultiNew  (Fabbri et al., 2019), TREC (Li & Roth, 2002; Hovy et al., 2001) and LCC (Mohler et al., 2016).
+The additional $50$ data items from the LongEval (Li et al., 2023) dataset are also added to validate the retrieval ability.
 For the datasets that do not contain the validation split, namely TREC, MultiNews and LCC, we sample from the test split and ensure different data items with the perplexity evaluation dataset.
@@
 The first block of tokens is not masked as the attention sink.
-For the profile stage, we use the MultiNews [17] calibration dataset with model response as supervision, as described in Section 5.
+For the profile stage, we use the MultiNews (Fabbri et al., 2019) calibration dataset with model response as supervision, as described in Section 5.
 We use $50\times 3$ data items at 2k, 4k, 8k lengths.
@@
 The optimization is done with the multi-objective optimization at the same set of lengths.
+We limit the number of distinct rules to at most two per model layer to ensure inference-time efficiency.
 Among the Pareto front solutions, we select the one with the lowest perplexity on the validation dataset of length 12k.
@@
 For MoA, The implementation is based on Huggingface Transformers.
-During the prefill stage, we use the sparse triton [58] kernel designed by us with block size 64.
+During the prefill stage, we use the sparse CUDA kernel designed by us with block size 64.
 During the decode stage, we modify the KV-Cache implementation to support our heterogeneous elastic rules.
 Thanks to our fixed sliding-window span during the decode stage, we simply replace the old KV-Cache that exceeds the span with the latest KV-Cache.
-Then, the dense FlashAttention [13] decoding is applied with the updated KV-Cache.
-For H2O, we use its official efficient implementation, which is based on Flexgen [54].
+Our custom decoding CUDA kernel then handles KV-Cache with varying lengths across different attention heads during the decoding process.
+For H2O, we use its official efficient implementation, which is based on Flexgen (Sheng et al., 2023).
 Note that H2O uses dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction, which requires the attention matrix to be explicitly calculated.
@@
 Following the performance experiments, we use Vicuna-7B and Vicuna-13B for efficiency tests whenever possible.
-However, the official efficient implementation of H2O based on Flexgen only supports OPT [74].
+However, the official efficient implementation of H2O based on Flexgen only supports OPT (Zhang et al., 2022).
 Therefore, we use OPT-6.7b and OPT-13b models for H2O in Table 9 for comparison.
@@
 Subfigure (a) shows results for StreamingLLM, and subfigure (b) for MoA.
+Figure 7: Retrieval accuracy tests on LongEval with Vicuna-7B.
+(a) Varies input lengths and densities to show effective context lengths across attention spans, (b) Set input length at 8k and show retrieval accuracy across different densities.
 We conduct a detailed experiment to test the retrieval ability of different attention methods across various attention spans and input lengths.
-Following previous work [7, 60], we quantify effective context length as the maximum input length where retrieval accuracy remains above a 90% threshold.
 Figure 6 shows the detailed data for effective context length calculation.
 As shown in the figure, StreamingLLM can hardly maintain retrieval accuracy when the input length is beyond the attention span, while MoA can effectively extend the effective context length.
+Following previous work (Chen et al., 2023; Tworkowski et al., 2023), we quantify effective context length as the maximum input length where retrieval accuracy remains above a 90% threshold.
+As shown in Figure 7(a), StreamingLLM and H2O achieve effective context lengths of no more than 2k tokens beyond their attention spans.
+In contrast, MoA expands its effective context length to approximately 3.9×3.9$\times$3.9 × its attention span before reaching up to the 12k limit of the original model.
+Figure 7(b) further shows that at a fixed input length of 8k, MoA reaches over 0.9 retrieval accuracy with just 25% density, whereas StreamingLLM and H2O require 100% and 75% density, respectively.
 B.1.3 Long-Context Understanding
 Figure 8: LV-Eval score of StreamingLLM and MoA at various densities on Vicuna-7B model.
-We conduct experiments with various densities on the LV-Eval benchmark [70].
+We conduct experiments with various densities on the LV-Eval benchmark (Yuan et al., 2024).
 As shown in Figure 8, MoA constantly outperforms the uniform static attention baseline StreamingLLM at various densities, demonstrating the effectiveness of our heterogeneous elastic rules.
+Table 7: LongBench scores for various models with different attention methods. All methods employ 50% density in the decode stage.
+LongBench $\uparrow$
+Model Attention 0-4k 4-8k 8-16k
+Vicuna-7B Original 37.91 33.82 32.54
+H2O 36.23 32.74 31.81
+InfLLM 35.23 33.54 30.15
+StreamingLLM 30.53 33.28 31.70
+MoA 37.04 32.90 31.94
+Vicuna-13B Original 42.25 39.52 35.93
+H2O 41.63 38.02 34.75
+InfLLM 39.36 37.66 34.36
+StreamingLLM 30.65 33.07 32.68
+MoA 41.73 38.88 35.69
+Llama3-8B Original 44.27 43.53 43.26
+H2O 43.46 43.01 42.50
+InfLLM 42.78 42.69 41.81
+StreamingLLM 37.20 38.02 39.43
+MoA 43.07 42.75 43.09
+Llama3-70B Original 50.70 48.05 48.55
+H2O 50.16 47.77 OOM
+StreamingLLM 45.14 42.40 40.04
+MoA 49.74 46.80 46.84
+Table 8: Performance comparison across different models and attention methods with the LV-Eval dataset. The numbers in brackets indicate the number of sub-datasets for the category.
+Single-QA Multi-QA Retrieval
+Model Attention w/o. Conf (2) w. Conf (2) w/o. Conf (3) w. Conf (2) w. Conf (2)
+Vicuna-7B Original 10.49 6.29 6.83 5.60 0.00
+H20 9.16 6.20 6.44 4.80 0.00
+InfLLM 7.11 6.70 6.07 4.80 0.00
+StreamingLLM 7.54 5.90 5.98 3.56 0.00
+MoA 9.98 6.27 6.16 5.31 0.09
+Vicuna-13B Original 10.64 7.28 5.32 5.07 1.08
+H20 9.53 6.54 5.25 5.36 1.83
+InfLLM 10.21 9.35 6.03 3.19 2.08
+StreamingLLM 9.05 5.86 5.37 3.19 3.70
+MoA 11.04 6.93 5.79 5.84 6.88
+Llama3-8B Original 34.05 19.51 11.41 17.70 7.84
+H20 28.52 17.05 11.11 15.98 9.95
+InfLLM 24.94 17.75 10.61 14.80 6.04
+StreamingLLM 20.21 9.57 8.14 9.36 10.03
+MoA 32.98 20.53 10.65 17.57 8.98
+Llama3-70B Original 44.44 25.02 16.71 22.86 17.43
+StreamingLLM 26.63 14.22 14.04 14.70 19.38
+MoA 42.44 23.58 15.75 21.27 19.19
+Figure 9: (a) LV-Eval and (b) LongBench scores for different attention methods at 50% density, tested on Vicuna-7B, 13B and Llama3-70B models. Scores normalized against the original dense model.
+We detailed the respective scores for LongBench and LV-Eval in Table 7 and Table 8. The number in the bracket of Table 8 indicates the number of sub-datasets for the category.
 B.2 Efficiency
 B.2.1 Throughput Breakdown
-Table 7: Improvements in throughput (tokens per second) across various configurations and input sizes for MoA over FlashAttn2 on models with 7b and 13b parameters.
-The batch sizes for FlashAttn2 are configured to the maximum capacity of an A100-80GB GPU.
-4k 8k 16k
-Size Configuration Batch Throughput Batch Throughput Batch Throughput
-7B FlashAttn2 30 134.6 15 66.9 8 32.9
-+Static KV-Cache 30 496.1 15 219.5 8 91.6
-+Reduced Attention 30 722.5 15 369.9 8 178.3
-+Increased Batch Size 50 897.7 25 436.7 12 206.4
-13B FlashAttn2 16 81.3 8 40.8 4 19.8
-+Static KV-Cache 16 264.6 8 111.3 4 62.2
-+Reduced Attention 16 329.6 8 156.4 4 87.3
-+Increased Batch Size 28 471.5 14 222.6 7 108.3
-In this section, we conduct a detailed analysis of the throughput breakdown for MoA, compared to the baseline comprising Huggingface with FlashAttention.
+Table 9: Efficiency analysis of different frameworks on 7B and 13B models. H2O and MoA use 50% density. GPU memory evaluated with batch sizes 8 (7B model) and 4 (13B model).
+Memory (GB)
+Size Framework 4k 8k 16k
+7B FlashAttn2 28.5 44.4 76.3
+H2O 36.9 OOM OOM
+MoA 22.7 32.9 53.5
+13B FlashAttn2 36.8 49.2 74.0
+H2O 40.4 77.9 OOM
+MoA 32.0 39.6 55.0
+In this section, we explain the decode throughput breakdown for MoA, compared to the baseline comprising Huggingface with FlashAttention2.
 The observed increase in throughput, as shown in Table 5, primarily stems from three aspects:
@@
 Increased Batch Size. With the reduced size of KV-Cache, MoA supports a larger batch size, contributing to the increase in throughput.
-B.2.2 Comparison with vLLM
-Table 8: Throughput comparison of MoA and vLLM on 7B and 13B models.
-Throughput (tok/s)
-Size Framework 4k 8k 16k
-7B vLLM 628.8 323.0 145.5
-MoA 897.7 436.7 206.4
-13B vLLM 314.8 160.5 71.1
-MoA 471.5 222.6 108.3
-We compare the throughput of MoA with the highly optimized inference framework vLLM [33] as shown in Table 8.
-MoA still achieves a 1.4×1.4$\times$1.4 × to 1.5×1.5$\times$1.5 × increase in throughput.
-Notably, vLLM incorporates extensive system-level optimizations that are orthogonal to MoA, suggesting the potential for MoA to further improve.
-As vLLM constantly pre-allocates a 71GB memory and automatically manages the actual batch size, we exclude memory metrics from the comparison.
-B.2.3 Automatic Optimization Overhead
-Table 9: Time overhead for various stages of MoA.
-Stage Time
-Calibration Dataset Generation 15 min
-Profile 25 min
-Optimize 25 min
-Validate 40 min
-Sum. 105 min
+B.2.2 Automatic Compression Pipeline Overhead
+Table 10: Compression overhead for various stages of MoA across models with differing parameter sizes, reported as the amount of GPU $\times$ latency, except when only one GPU is used.
+Larger models necessitate more GPUs due to model parallelism.
+All stages utilize GPUs, except for the Optimize stage, which uses the CPU.
+Stage 7B LLM 13B LLM 70B LLM
+Calibration Data Gen. 10min 15min 2 $\times$ 60min
+Profile 20min 2 $\times$ 25min 8 $\times$ 210min
+Optimize (CPU) 30min 25min 100min
+Validate 35min 40min 2 $\times$ 140min
+Total Latency 1h 35min 1h 45min 8h 30min
+Total GPU Time 1h 5min 1h 45min 34h 40min
+Table 11: Progressive compression overhead for various stages of MoA with respect to different parameter sizes and calibration (validation) dataset sizes.
+Stage Complexity w.r.t parameter size Complexity w.r.t dataset size
+Calibration Dataset Gen. Linear Linear
+Profile Linear Linear
+Optimize Polynomial $\sim$ Exponential for #Head Irrelevant
+Validate Linear Linear
+Empirical Latency Almost Linear Linear
 We present a detailed breakdown of the time usage of MoA pipeline.
@@
 Validating each one of the compression plans takes about 4 minutes, totaling around 40 minutes.
+We also show the progressive compression overhead for MoA in Table 11.
 B.3 Ablation Study
 B.3.1 Calibration Dataset
-Table 10: Performance comparison on various test sets, using different calibration sets. Tested on Vicuna-7B model. The result is tested with 50% density MoA on LongBench [4] 0-4k split.
+Table 12: Performance comparison on various test sets, using different calibration sets.
+Tested on Vicuna-7B model.
+The result is tested with 50% density MoA on LongBench (Bai et al., 2023) 0-4k split.
 Test Score
@@
 In this section, we validate the robustness of our calibration dataset design principles.
-We select three sub-tasks and respective datasets from the LongBench benchmark, including Qasper [15], MultiNews [17], and TREC [38, 28].
+We select three sub-tasks and respective datasets from the LongBench benchmark, including Qasper (Dasigi et al., 2021), MultiNews (Fabbri et al., 2019), and TREC (Li & Roth, 2002; Hovy et al., 2001).
 We use their training set to construct the calibration dataset, and use their respective test set in LongBench to calculate the score.
@@
 Then, in the final layers, most heads require low density, while some outlier heads need high density.
-This observation conforms to previous findings of the intrinsic dimension of LLM [61].
+This observation conforms to previous findings of the intrinsic dimension of LLM Valeriani et al.
+(2023).
 The geometry of density is similar to the intrinsic dimension of LLM, with two local minima.
@@
 Given any token, two kinds of information are used as the model inputs: position encoding and token embedding.
-Position encoding indicates the absolute [74] or relative positions [59] of tokens in the sentence.
+Position encoding indicates the absolute (Zhang et al., 2022) or relative positions (Touvron et al., 2023) of tokens in the sentence.
 Token embedding maps different tokens as different vectors.
@@
 MoA’s generated mask density shows positive relation with SoE, suggesting that MoA successfully captures the semantic information of the dense language model as shown in Figure 12.
-Appendix D MoA Automatic Pipeline Details
-D.1 Derivation of Attention Influence
+Appendix D Automatic Pipeline Details
+D.1 Addtional Oracle on Elastic Pattern Design
+Figure 13: Examples of attention matrices from different attention heads (columns) and tasks (rows) of the Vicuna-7B model.
+The attention matrices were averaged over 256 data items per dataset.
+The same head shows a similar attention span across different tasks, explaining the robust cross-dataset generalizability of our method.
+We visualize the attention matrix of the same attention heads across three additional tasks in Figure 13, as an extension of Figure 2.
+The consistent attention span across tasks sheds light on the strong cross-dataset generalization ability of our MoA method.
+D.2 Derivation of Attention Influence
 We use the first-order Taylor expansion to calculate the influence of each attention value.
-This approximation approach is supported by methodologies commonly employed in other LLM compression approaches [36, 55, 14, 29].
+This approximation approach is supported by methodologies commonly employed in other LLM compression approaches (Li et al., 2024a; Shi et al., 2021; Das et al., 2023; Jiang et al., 2023).
 As discussed in Section 4.1, when masking out attention value $A_{h,i,j}$ at head $h$, row $i$, and column $j$, it also influences the attention values in the same row by $\Delta A_{h,i,n|j}$.
@@
 $\textbf{E}_{h}=\frac{\textbf{A}_{h}}{1-\textbf{A}_{h}}\cdot\left(\frac{\partial L}{\partial\textbf{A}_{h}}-\left(\frac{\partial L}{\partial\textbf{A}_{h}}\cdot\textbf{A}_{h}\right)\mathbbm{1}^{N\times N}\right).$ (13)
-D.2 Optimization Details
-D.2.1 Optimizing at Single Length
+D.3 Optimization Details
+D.3.1 Optimizing at Single Length
 The optimization problem is formulated as follows:
@@
 Additionally, to enforce the restriction that each model layer only has a limited number of different plans, we bound the norm of element-wise multiplication of $\mathbf{X}_{h}=\begin{bmatrix}X_{h,0}&X_{h,1}&\cdots&X_{h,R_{h}-1}\end{bmatrix}^{\top}$ in a single layer.
-D.2.2 Optimizing at Multiple Lengths
+D.3.2 Optimizing at Multiple Lengths
 With the ability to optimize at a single length, we utilize the same framework for multi-objective MIP across various lengths.
-The key is to transform the multi-objective MIP problem into several single-objective MIP problems [47].
-We utilize the idea of epsilon-constraint method [72].
+The key is to transform the multi-objective MIP problem into several single-objective MIP problems (Paria et al., 2018).
+We utilize the idea of epsilon-constraint method (Yv et al., 1971).
 We first select each input length as our primary objective to perform the single-objective optimization on it while simultaneously recording the outcomes of other objectives.

### v3

@@
-MoA: Mixture of Sparse Attention for Automatic Large Language Model Compression
+Mixture of Attention Spans: Optimizing LLM Inference Efficiency with Heterogeneous Sliding-Window Lengths
 Abstract
 […]
-footnotetext: Corresponding author: Xuefei Ning (foxdoraame@gmail.com)
 1 Introduction
-Figure 1: Retrieval accuracy of the Vicuna-7B model using different attention methods across varying input lengths and retrieval positions on the LongEval benchmark (Li et al., 2023).
-This retrieval benchmark takes massive key-value pairs as inputs and tests the accuracy to retrieve values based on given keys from diverse positions.
+Large Language Models (LLMs) exhibit remarkable versatility across numerous applications (brown2020gpt3; tay2022efficientSurvey; Wan2023EfficientLLM).
+Central to LLM is the attention mechanism (vaswani2017attention), which computes interactions among tokens within a certain span, thereby enabling context understanding.
+Scaling input length is crucial for enhancing LLM capabilities (Chen2023PI; Tworkowski2023FocusedTrans), including fact retrieval, summarization, few-shot learning, question answering and so on (bai2023longbench; yuan2024lveval).
+However, the ever-growing attention computation and Key-Value Cache (KV-Cache) pose significant efficiency challenges (xiao2023streamingLLM).
+Figure 1: Retrieval accuracy of the Vicuna-7B model using different attention methods across varying input lengths and retrieval positions on the LongEval benchmark (lmsys2023longeval).
+The benchmark takes massive key-value pairs as inputs and tests the accuracy to retrieve values based on given keys from diverse positions.
 (a) Original model with a full attention span; (b) StreamingLLM with half the attention span, showing reduced effectiveness beyond the span; (c) MoA with half the attention span, maintaining effectiveness beyond the span.
-Large Language Models (LLMs) exhibit remarkable versatility across numerous applications (Brown et al., 2020; Tay et al., 2022; Wan et al., 2023).
-Central to LLM is the attention mechanism (Vaswani et al., 2017), which computes interactions among tokens within a certain span, thereby enabling context understanding.
-Scaling input length is crucial for enhancing LLM capabilities (Chen et al., 2023; Tworkowski et al., 2023), including fact retrieval, summarization, few-shot learning, question answering and so on (Bai et al., 2023; Yuan et al., 2024).
-However, the ever-growing attention computation and Key-Value Cache (KV-Cache) pose significant efficiency challenges (Sheng et al., 2023; Xiao et al., 2024c; Han et al., 2023; Kwon et al., 2023).
-Previous work proposes sparse attention methods to address the efficiency challenges of long contexts in generative LLMs.
-These methods typically employ a uniform, fixed-span sliding window mask across all heads and input lengths, limiting attention to local contexts only (Xiao et al., 2024c; Han et al., 2023).
-This approach allows the LLM to take long inputs with a fixed attention span, keeping bounded attention computation and KV caching overhead.
-Following previous works (Chen et al., 2023; Tworkowski et al., 2023), we quantify the effective context length as the maximum input length where content retrieval accuracy exceeds a 90% threshold.
-In principle, fixed-span local attention can gradually aggregate global information through multiple model layers, yielding a longer effective context length than each attention span (Feng et al., 2022; Zaheer et al., 2020).
-Nonetheless, we reveal that uniform masks, like StreamingLLM (Xiao et al., 2024c), hardly extend effective context length beyond the span, as shown in Figure 6.
-Figure 1(b) further illustrates such limitation: with a 50% attention span mask, StreamingLLM fails to accurately retrieve content from the earlier half of the input and performs even worse at longer input lengths.
+Previous work has proposed sliding-window attention to address the efficiency challenges of long contexts in generative LLMs.
+These methods typically employ a uniform, fixed-span sliding-window mask across all heads and input lengths, limiting attention to local span only (xiao2023streamingLLM; han2023lmInfinite).
+This design allows the LLM to process long inputs within a bounded attention computation and KV caching overhead.
+Following previous works (Chen2023PI; Tworkowski2023FocusedTrans), we quantify the effective context length as the maximum input length where content retrieval accuracy exceeds a 90% threshold.
+In principle, fixed-span local attention can gradually aggregate global information through multiple model layers, yielding a longer effective context length than each attention span (feng2022diffuser; zaheer2020bigbird).
+Nonetheless, we reveal that uniform sliding-window methods, such as StreamingLLM (xiao2023streamingLLM), hardly extend effective context length beyond the span.
+As shown in Figure 1(b), with a 50% attention span, StreamingLLM fails to accurately retrieve earlier half of the input and performs even worse at longer input lengths.
 Figure 2 reveals one possible explanation for the problem: while some attention heads focus on local contexts, others encompass the broad span of the entire input sequence.
-Consequently, the uniform approach fails to achieve a long effective context length as it limits the attention span of the global-context heads, while excessively allocates compute and memory budget for local-context heads.
+Consequently, the uniform approach fails to achieve a long effective context length, as it limits the attention span of global-context heads while excessively allocating compute and memory budgets to local-context heads.
 Additionally, as the input length increases, some attention heads need a faster increase in attention span than others to avoid serious performance degradation, as shown in Table 1.
 Unfortunately, the uniform approaches do not include heterogeneous rules to scale the attention spans differently for various heads.
-Besides, existing model compression methods (Men et al., 2024; Lin et al., 2023; Xiao et al., 2024b; Li et al., 2024a; Kim et al., 2023; Li et al., 2024b) use general language modeling corpora to decide the compression plan, which cannot accurately profile the influence of compression on long-context tasks.
-In this work, we propose Mixture of Attention (MoA), a training-free sparse attention method.
-As illustrated in Figure 3, MoA constructs the search space of heterogeneous elastic rules of attention spans.
-For automatic LLM compression, MoA first utilizes gradient-based profiling to inspect the influences of each attention position on the prediction loss.
-Based on the profiling results, MoA tailors heterogeneous sparse attention configurations for each model layer and attention head.
-During profiling, MoA employs a calibration dataset with long-range dependencies and uses the original dense model’s response instead of the human-written response as the reference to calculate the loss.
-This ensures an accurate profiling of the attention influences to facilitate better compression results.
+In addition, existing model compression methods, including quantization and sparse attention (smooth_quant; llm-mq; squeezellm; li2024evaluating), typically rely on general language modeling corpora to determine compression configurations, which fail to accurately capture their impact on long-context tasks.
+In this work, we propose Mixture of Attention Spans (MoA), a training-free heterogeneous sliding-window attention method.
+As shown in Figure 3, MoA constructs a search space of heterogeneous and elastic rules that scale each head’s window length with input length.
+To automate rule selection, MoA first utilizes gradient-based profiling to inspect the influences of each attention position on the prediction loss.
+Based on the profiling results, MoA tailors heterogeneous window length for each attention head and input length.
+During profiling, MoA employs a calibration dataset with long-range summaries from a full-attention model, instead of the human-written summaries, as the reference to calculate the loss.
+This ensures an accurate profiling of attention influences to facilitate better performance.
 Our contributions are summarized as follows.
   • Heterogeneous Elastic Rules.
-We propose heterogeneous elastic rules for masks of each attention head.
-We formulate MoA compression search space to include a diverse range of elastic rules that tailor the local attention span relative to the input length for each attention head.
-The heterogeneous elastic rules improve the fact retrieval accuracy of MoA from 25% to 98% compared with masks with uniform span and scaling function for each head.
+We propose heterogeneous elastic rules for sliding-window length of each attention head.
+We formulate MoA configuration search space to include a diverse range of elastic rules that tailor the local attention span relative to the input length for each attention head.
+The heterogeneous elastic rules improve content retrieval accuracy from 25% to 98% over the uniform counterpart.
   • Calibration Dataset Construction We emphasize the importance of data engineering in LLM compression.
 Our findings demonstrate that, instead of relying on general language modeling datasets and human responses, using datasets with long-range dependencies and referencing the original LLM’s responses is essential for accurately profiling the effects of compression.
-  • Automatic Optimization.
-We propose an automatic pipeline to find the optimal compression plan encompassing heterogeneous elastic rules for various attention heads.
-This pipeline can efficiently find the optimal plan within several hours, for example, two hours for compressing Vicuna-13B.
-Experiments show that MoA achieves $\times6$ 8.2 × throughput improvements over FlashAttention2, 1.71.9×1.7-1.9$\times$1.7 - 1.9 × over vLLM framework on 7B and 13B dense LLMs at a 50% density (the average of KV-Cache length / input length), with only a 1% average relative degradation in retrieval accuracy.
-In Section 6.4, we show that this significant throughput improvements of MoA over FlashAttention2 can be attributed to four factors: (1) the static size of the KV-cache, (2) reduced attention computations, (3) increased batch size enabled by reduced memory usage, and (4) a specialized kernel implementation.
-Additionally, MoA achieves over 90% retrieval accuracy with just 25% average density, far surpassing sparse attention baselines that need a density of 75% to 100% for similar performance.
-On long-context understanding benchmarks, MoA performs comparably to dense models, with a maximum relative performance drop of less than 5%, which is about one-sixth of that observed with the uniform sparse attention baseline.
-Our code is available at https://github.com/thu-nics/MoA.
+  • Automatic Configuration Search.
+We propose an automatic pipeline to find the optimal configuration encompassing heterogeneous elastic rules for various attention heads.
+This pipeline efficiently finds the optimal configuration within several hours—for example, two hours for Vicuna-13B.
+[…]
 2 Preliminary and Related Work
 2.1 Attention Mechanism
-The Multi-Head Self Attention (MHA) mechanism (Vaswani et al., 2017) is crucial to the functionality of LLMs.
+The Multi-Head Self-Attention (MHA) mechanism (vaswani2017attention) is crucial to the functionality of LLMs.
 It starts with an input sequence transformed into query (Q), key (K), and value (V) matrices through linear projections.
@@
 During prefill, the model processes the entire input sequence to generate the initial response token.
-In the subsequent decode stage, it uses the newly generated token and previously cached K and V matrices to produce subsequent tokens until the generation concludes.
+In the subsequent decode stage, it iteratively uses the newly generated token and previously cached K, V matrices to produce subsequent tokens.
 Although effective, this iterative process increases memory and computation demands due to the expanding KV-Cache.
-2.2 Efficient Attention
-Efficient methods are proposed to mitigate the computation and memory costs associated with attention.
-One branch of work uses dynamic sparse attention masks to adaptively skip attention computations during prefill stage (Pagliardini et al., 2023; Qu et al., 2022; Roy et al., 2021; Wang et al., 2021; Lu et al., 2021; Kitaev et al., 2020) or drop KV-Cache during decode stage (Anagnostidis et al., 2023; Zhang et al., 2023; Ge et al., 2023; Sheng et al., 2023; Liu et al., 2023a) based on the input sequences.
-However, due to the complex control and computation flow, dynamic prefill often requires specific hardware to achieve substantial wall-time speedup (Qu et al., 2022; Wang et al., 2021; Lu et al., 2021; Ham et al., 2021; 2020).
-Additionally, dynamic KV-Cache pruning in the decode stage may require extensive retraining (Anagnostidis et al., 2023), additional accumulated attention score computation (Sheng et al., 2023; Zhang et al., 2023; Liu et al., 2023a; Ge et al., 2023), or extensive memory swap for KV-Cache retrieval (Tang et al., 2024; Xiao et al., 2024a).
-Another branch of work uses static sparse attention, where predefined masks are applied consistently across all processed sentences.
+2.2 Efficient Attention Methods
+Efficient attention methods have been proposed to mitigate the computation and memory costs.
+One branch of work dynamically skip attention computations during the prefill stage (Pagliardini2023SparseFlash; qu2022dota; roy2021routingTransformer; wang2021spatten; lu2021sanger; kitaev2020reformer) or drop the KV-Cache during the decode stage (Anagnostidis2023DynamicPruning; Zhang2023H2O; Ge2023FastGen; sheng2023flexgen; Liu2023Scissorhands) based on the input sequence.
+However, due to the complex control and computation flow, dynamic sparse prefill often requires specific hardware to achieve substantial wall-time speedups (qu2022dota; wang2021spatten; lu2021sanger; ham2021elsa; ham2020A3).
+Additionally, dynamic KV-Cache pruning in the decode stage may require extensive retraining (Anagnostidis2023DynamicPruning), additional pruning score computation (sheng2023flexgen; Zhang2023H2O; Liu2023Scissorhands; Ge2023FastGen; li2024snapkv; cai2024pyramidkv), or extensive memory swapping for KV-Cache retrieval (tang2024quest; xiao2024infllm).
+Another branch of work uses static sparse attention, where predefined masks are applied consistently across all input sequences.
 Thanks to the fixed computation flow, static sparse attention is generally more efficient and GPU-friendly.
-For language understanding models such as BERT (Devlin et al., 2018), various masks are used (Zaheer et al., 2020; Beltagy et al., 2020; Child et al., 2019; Zhou et al., 2024; Xiao et al., 2024c; Han et al., 2023).
-But for generative LLMs, the predominant method is the fixed-span sliding window mask with global attention on a few initial tokens (Xiao et al., 2024c; Han et al., 2023).
-With the local attention pattern, the KV-Cache beyond the current attention span can be dropped, saving much memory for long sequence scenarios.
-However, the uniform static masks across different attention heads and input lengths are model- and data-agnostic, which can compromise LLMs’ effective context length and lead to suboptimal performance in long sequence scenarios.
+For language understanding models such as BERT (devlin2018bert), various masks are used (zaheer2020bigbird; beltagy2020longformer; child2019spTrans; Zhou2024EfficientSurvey; xiao2023streamingLLM; han2023lmInfinite).
+But for generative LLMs, the predominant method is the uniform sliding-window mask with global attention on a few initial tokens (xiao2023streamingLLM; han2023lmInfinite).
+For these approaches, the KV-Cache beyond the local attention span can be dropped, saving much memory for long sequence scenarios.
+However, uniform sliding-window masks across different attention heads and input lengths are model- and data-agnostic, which can compromise LLMs’ effective context length and lead to suboptimal performance in long-sequence scenarios.
 Our method falls within this category, benefiting from the efficiency and training-free advantages, while addressing the performance limitations encountered by previous methods.
-In addition to sparse attention, alternative mechanisms have been proposed to replace traditional attention for long-sequence modeling (Gu & Dao, 2023; Peng et al., 2023; Sun et al., 2023; Poli et al., 2023; Li et al., 2022; Kacham et al., 2023; Peng et al., 2021; Choromanski et al., 2020; Wang et al., 2020).
-However, these new mechanisms often require different weights compared to vanilla transformers, imposing significant re-training overhead for LLMs.
-Previous works also propose LLM acceleration frameworks (Gugger et al., 2022; Aminabadi et al., 2022; Sheng et al., 2023; Kwon et al., 2023), as well as kernel-level optimizations (Dao et al., 2022; Dao, 2023; Shah et al., 2024).
+Previous works also propose LLM acceleration frameworks (huggingface2022accelerate; Aminabadi2022DeepSpeed; sheng2023flexgen; Kwon2023vllm), as well as kernel-level optimizations (dao2022flashattention; dao2023flashattention2; shah2024flashattention3).
 These kernel and system optimizations are orthogonal to our work and can be integrated to further enhance efficiency.
-3 Mixture of Attention (MoA)
-We first illustrate the heterogeneity of the attention patterns in pre-trained LLMs in Section 3.1.
-Based on this insight, we define the search space for our Mixture-of-Attention (MoA) method in Section 3.2.
+3 Mixture of Attention Spans (MoA)
+We first illustrate the heterogeneity of attention patterns in pre-trained LLMs in Section 3.1. Based on this insight, we define the search space for our MoA method in Section 3.2.
 Figure 2: Examples of attention matrices from different attention heads of the Vicuna-7B model. Each attention matrix is averaged over 256 data items from the LongEval dataset.
@@
 17,18,19 0.97 0.94 0.97
-Table 1: Retrieval accuracy of Vicuna-7B with sliding-window sparse attention across various model layers, window spans, and input lengths.
+Table 1: Retrieval accuracy of Vicuna-7B with sliding-window attention across various model layers, window spans, and input lengths.
 3.1 Mixture of Attention Patterns and Elastic Rules
@@
 For example, the first head primarily focuses on local contexts with a narrow-span sliding window, while the third head covers nearly the entire input, indicating global attention.
-The attention spans of different heads mostly remain constant across various tasks and datasets, as shown in Appendix D.1.
-Table 1 demonstrates that applying the same sliding-window sparse attention mask across model layers can lead to a 65% variance in retrieval accuracies.
-It conforms to the multi-head self-attention design principle of capturing varied information (Vaswani et al., 2017), as well as the findings from concurrent research that identifies specific attention heads for global text retrieval (Wu et al., 2024).
+The attention spans of different heads mostly remain constant across various tasks and datasets, as shown in Appendix E.1.
+Table 1 demonstrates that applying the same sliding-window attention mask across model layers can lead to a 65% variance in retrieval accuracies.
+This aligns with the multi-head self-attention design principle of capturing varied information (vaswani2017attention), as well as findings from concurrent research that identify specific attention heads for global text retrieval (wu2024retrievalHead).
 Heterogeneous Elastic Rules.
 In addition to heterogeneity at a certain length, different attention heads also exhibit varying elastic behaviors as the input length changes.
-Figure 2 illustrates this variability: for shorter inputs (the upper left part of the attention matrix), the second and third heads initially show global attention.
-However, as input length increases, the second head remains the medium-span local focus, while the third head continues to expand as global attention.
+Figure 2 illustrates this variability: for shorter inputs (the upper-left part of the attention matrix), the second and third heads initially show global attention.
+However, as input length increases, the second head maintains a medium-span local focus, while the third head continues to expand as global attention.
 Table 1 further evidences the diverse elastic rules.
-For example, at 4k input length, a 2k sliding-window sparse attention mask on layers 9 to 11 yields better retrieval accuracy than on layers 17 to 19.
-However, the opposite is true for an 8k input length.
-This data supports the visual observations from Figure 2, highlighting that attention patterns respond to input length scaling differently.
+For example, at a 4k input length, a 2k sliding-window attention mask on layers 9 to 11 yields better retrieval accuracy than on layers 17 to 19; however, the opposite holds for an 8k input length.
+These data support the visual observations from Figure 2, highlighting that attention patterns respond differently to input-length scaling.
 Leveraging these insights, MoA encompasses heterogeneous elastic rules as the search space.
 3.2 Heterogeneous Elastic Rule Search Space
-In designing the search space for the MoA mask, we consider the inherently heterogeneous and elastic nature of LLM attention patterns.
-As shown in Figure 3(a), we adopt a hardware-friendly sliding-window mask as our base sparse attention mask (Beltagy et al., 2020).
-Following previous work (Xiao et al., 2024c; Han et al., 2023), the initial few tokens (64 tokens for MoA) are not masked.
-The attention span equals the sliding-window-span plus the number of initial unmasked tokens.
+In designing the search space for MoA, we consider the inherently heterogeneous and elastic nature of LLM attention patterns.
+As shown in Figure 3(a), we adopt a hardware-friendly sliding-window mask as our base attention mask (beltagy2020longformer).
+Following previous work (xiao2023streamingLLM; han2023lmInfinite), the initial few tokens (64 tokens for MoA) are not masked.
+The attention span equals the sliding-window span plus the number of initially unmasked tokens.
 We define the attention span S of head $h$ at input length N using a straightforward linear function:
 $S_{h}=\alpha_{h}+\beta_{h}\cdot N,$ (2)
-where $\alpha_{h}$ and $\beta_{h}$ are hyperparameters that control the base span and its expansion rate with input length of a specific attention head.
-The $\alpha$ and $\beta$ hyperparameters of each attention head can be chosen from multiple discrete options.
-For LLMs with many layers, the MoA search space can become quite large.
-For example, for a 7B model consisting of $32$ attention heads and $32$ layers, with $6$ and $9$ different $\alpha$ and $\beta$ options for each head, the potential search space expands to $54^{1024}$ configurations.
-Thus, we design the automatic pipeline to efficiently pinpoint the optimal $\alpha$s and $\beta$s for any LLM.
-4 Automatic Pipeline for MoA Compression
-Figure 3: Overview of the MoA.
-(a) The sparse attention search space includes heterogeneous elastic rules of the attention span on sliding-window masks.
-(b) The automatic compression pipeline begins with a calibration dataset, which includes long-dependency contexts and supervision texts generated by the original dense LLM.
+where $\alpha_{h}$ and $\beta_{h}$ are hyperparameters that control the base span and its expansion rate with input length for a specific attention head.
+[…]
+4 Automatic MoA Configuration Search Pipeline
+Figure 3: Overview of MoA.
+(a) The search space includes heterogeneous elastic rules of the attention span on sliding-window masks.
+(b) The automatic rule search pipeline begins with a calibration dataset, which includes long-dependency contexts and supervision texts generated by the original dense LLM.
 MoA profiles each attention value’s impact on model predictions within this dataset, revealing accuracy losses for different candidate elastic rules across various input lengths.
 The final optimization step selects elastic rules for each attention head to minimize the total prediction loss while adhering to specified density constraints.
-This section outlines the MoA automatic compression pipeline as shown in Figure 3(b).
+This section outlines the MoA automatic configuration search pipeline, as shown in Figure 3(b).
 Starting with a trained LLM and a calibration dataset, MoA first profiles the influence of each attention value on the model’s prediction loss for various input sequences from the calibration dataset.
 The masked sum of the influences represents the accuracy loss associated with each mask at different input lengths, showing the accuracy loss each candidate elastic rule could cause at that length.
-Then, MoA optimizes the compression plan by selecting the optimal elastic rule for each head, which minimizes the accuracy loss across various lengths while adhering to specified density constraints.
+Then, MoA optimizes the window-lengths by selecting the optimal elastic rule for each head, which minimizes the accuracy loss across various input lengths while adhering to specified density constraints.
 The following sections provide detailed discussions of each step in this pipeline.
+Detailed definitions of key notions are in Appendix A.
+The user-friendly code interface is described in Appendix E.4.
 4.1 Attention Influence Profiling
-In the profile step, MoA quantifies the impact of individual attention values on the final prediction loss of a pre-trained LLM.
+In the profiling step, MoA quantifies the impact of individual attention values on the final prediction loss of a pre-trained LLM.
 It informs the subsequent step about the influence of masking each attention value, revealing the accuracy trade-offs of the candidate elastic rules for each attention head.
 The influence of each attention value is derived from the attention matrix A and its gradient $\partial L/\partial\textbf{A}$, computed over a calibration dataset.
-When applying sparse attention masks, we approximate the change in the model’s prediction loss, $\Delta L$, using a first-order Taylor expansion based on variations in the attention matrices A: $\Delta L=\sum_{h}\sum_{i}\sum_{j}\partial L/\partial A_{h,i,j}\cdot\Delta A_{h,i,j}$.
+When applying sliding-window attention masks, we approximate the change in the model’s prediction loss, $\Delta L$, using a first-order Taylor expansion based on variations in the attention matrices A: $\Delta L=\sum_{h}\sum_{i}\sum_{j}\partial L/\partial A_{h,i,j}\cdot\Delta A_{h,i,j}$.
 Here, $h$ indexes the attention heads across all layers, and $i,j$ are the row and column indices within each attention matrix $\textbf{A}_{h}$.
+Details on the calibration dataset and the prediction loss $L$ are provided in Section 5.
 We define the attention influence matrix, $\textit{E}_{h,i,j}$, as the estimated change in loss, $\Delta L$, if the attention value $A_{h,i,j}$ is masked (i.e., set to zero).
@@
 Thus, setting one attention value at column $j$ to zero causes an increase in the other attention values, $\Delta A_{i,n|j},n\neq j$, within the same row.
-These two effects are integrated into the following formulation, whose derivation is provided in Appendix D.2:
+These two effects are integrated into the following formulation, whose derivation is provided in Appendix E.2:
 $E_{i,j}=\sum_{n}\frac{\partial L}{\partial A_{i,n}}\cdot\Delta A_{i,n|j}=\frac{\partial L}{\partial A_{i,j}}\cdot(-A_{i,j})+\sum_{n\neq j}\frac{\partial L}{\partial A_{i,n}}\cdot A_{i,n}\cdot\frac{A_{i,j}}{1-A_{i,j}}$ (3)
@@
 The average attention influence is calculated respectively for different input lengths.
-The gradient ${\partial L}/{\partial\textbf{A}_{h}}$ is computed using chain derivative in deep learning frameworks like PyTorch (Paszke et al., 2019).
+The gradient ${\partial L}/{\partial\textbf{A}_{h}}$ is computed using chain derivative rule in deep learning frameworks like PyTorch (paszke2019pytorch).
 The detailed calibration dataset setup is discussed in Section 5.
 With the average attention influence of each head, MoA can calculate the accuracy loss of applying a candidate elastic rule at a specific input length.
-The loss is calculated with the sum of masked average attention influence according to the rule.
+The loss is calculated as the sum of masked average attention influence according to the rule.
 We denote $\textbf{M}_{r_{h}}$ as the binary mask at head $h$ that corresponds to rule $r$, with masked positions marked as 1 and others as 0.
@@
 $\Delta L=\sum_{h}\Delta L_{h,r_{h}}=\sum_{h}\sum_{i}\sum_{j}M_{r_{h},i,j}\cdot\bar{E}_{h,i,j}.$ (4)
-After the profile stage, MoA acquires the unique accuracy-density trade-offs of elastic rules.
-It informs the allocation of denser masks to more sensitive heads and lighter masks to less sensitive ones.
+After the profiling stage, MoA acquires the estimated accuracy impact of elastic rules across attention heads.
+This informs the allocation of longer windows to more sensitive heads and shorter windows to less sensitive ones.
 Profiling at different input lengths enables the identification of the most effective elastic rules, even for unseen lengths.
@@
 MoA automatically selects the optimal elastic rule for each attention head to minimize accuracy losses across various sequence lengths under density budgets.
-Based on the profiling results, MoA first identifies Pareto front compression plans where any improvement in accuracy loss at one profile length would worsen another.
-To ensure the best generalization to lengths beyond those profiled, MoA then selects the plan that yields the minimum loss at an unseen length among the Pareto front solutions as the final plan.
-Specifically, we utilize multi-objective optimization to search for a set of Pareto optimal compression plans across the profiled lengths.
+Based on the profiling results, MoA first identifies Pareto-front configurations where any improvement in accuracy loss at one profiled length would worsen another.
+To ensure the best generalization to lengths beyond those profiled, MoA then selects the configuration that yields the minimum loss at an unseen length among the Pareto front solutions as the final configuration.
+Specifically, we utilize multi-objective optimization to search for a set of Pareto optimal configurations across the profiled lengths.
 The objective for each length is to minimize the total accuracy loss while conforming to any user-defined density constraints.
@@
 $\mathop{\arg\min}_{r_{h}\in\mathbb{R}}\Delta L^{(N_{i})},N_{i}\in\mathbb{N}_{\text{profile}}\quad\operatorname{s.t.}\frac{1}{H}\sum_{h=1}^{H}d^{(N_{i})}_{r_{h}}\leq d^{(N_{i})}_{\text{constr}},\forall N_{i}\in\mathbb{N}_{\text{constr}}.$ (5)
-Here, superscript $(N)$ denotes values at different lengths; $\mathbb{N}_{\text{profile}}$ and $\mathbb{N}_{\text{constr}}$ denote the sets of lengths for profiling and those subject to density constraints, respectively; $\mathbb{R}$ denotes the set of candidate rules; $\Delta L^{(N_{i})}$ denotes the accuracy loss due to compression; $d_{r_{h}}^{(N_{i})}$ denotes the density of rule $r_{h}$ at head $h$; $d_{\text{constr}}^{(N_{i})}$ denotes the average density constraint; $H$ denotes the total number of attention heads.
-Such formulation corresponds to the classic multi-objective mixed-integer-planing problem, which can be effectively solved within minutes using existing linear solvers, like Gurobi (Gurobi Optimization, LLC, 2023).
-The detailed problem formulation and solving strategies are discussed in Appendix D.3.
-Among the Pareto optimal compression plans, we select the one with the minimum loss at the unseen validation length as the optimal solution.
-This approach allows us to avoid profiling at every possible length while increasing the likelihood that the plan will generalize effectively to unseen lengths.
-Thanks to this automatic pipeline, we efficiently get the elastic rules tailored for each attention head.
-With the pipeline, MoA minimizes the accuracy loss caused by attention sparsification, while conforming to user-defined density constraints.
+Here, superscript $(N)$ denotes values at different lengths; $\mathbb{N}_{\text{profile}}$ and $\mathbb{N}_{\text{constr}}$ denote the sets of lengths for profiling and those subject to density constraints, respectively; $\mathbb{R}$ denotes the set of candidate rules; $\Delta L^{(N_{i})}$ denotes the accuracy loss due to applying attention mask; $d_{r_{h}}^{(N_{i})}$ denotes the density of rule $r_{h}$ at head $h$; $d_{\text{constr}}^{(N_{i})}$ denotes the average density constraint; $H$ denotes the total number of attention heads.
+This formulation corresponds to the classic multi-objective mixed-integer-programming problem, which can be effectively solved within minutes using existing linear solvers, like Gurobi (gurobi).
+The detailed formulation and solving strategies are discussed in Appendix E.3.
+Among the Pareto-optimal MoA configurations, we select the one with the minimum loss at the unseen validation length as the optimal solution.
+This avoids profiling at every possible length and increases the likelihood of effective generalization to unseen input lengths.
 5 Dataset and Supervision
-Table 2: Calibration dataset design choices: dataset content, supervision, and response reference.
-Calibration dataset with long dependency and model alignment improves MoA performance on retrieval accuracy and perplexity.
-All tests are done at 25% average density at 8k input length.
 Dataset Supervision Reference Long Dep. Align Model Retrieval Acc. $\uparrow$ PPL $\downarrow$
@@
 MultiNews Summary Model 0.95 3.96
-In this section, we highlight the overlooked importance of calibration dataset design and its supervision objective in LLM compression.
-Calibration datasets are essential for sensitivity analysis across various compression techniques, including weight pruning (Men et al., 2024; Lee et al., 2024; Liu et al., 2023b) and quantization (Lin et al., 2023; Xiao et al., 2024b; Li et al., 2024a; Kim et al., 2023).
+Table 2: Calibration dataset design choices: dataset content, supervision, and response reference.
+Calibration dataset with long dependency and model alignment improves MoA performance on retrieval accuracy and perplexity.
+All tests are done at 25% average density at 8k input length.
+We emphasize the often-overlooked role of calibration dataset design and supervision objectives in LLM compression.
+Calibration datasets enable effective sensitivity analysis across compression methods like weight pruning (Men2024ShortGPT; lee2024cats) and quantization (awq; smooth_quant).
 In this work, MoA profiles the attention influence on the calibration dataset, which is crucial for subsequent automatic optimization.
 Current Approach.
-General language modeling datasets, such as human-written text corpus RedPajama (Computer, 2023), are commonly used as the calibration dataset.
-These datasets, supervised by next-token-prediction on the entire corpus, primarily capture attention patterns coherent with immediately preceding tokens.
-However, they lack long context dependencies, failing to address the global attention crucial for tasks like long-range retrieval.
-Moreover, a notable misalignment exists between the model response and the human-written supervision.
-Consequently, it leads to inaccuracies when using human responses to compute attention values and gradients during profiling.
+General language modeling datasets, such as the human-written text corpus RedPajama (together2023redpajama), are commonly used as calibration datasets.
+These datasets, supervised by next-token prediction on the entire corpus, primarily capture attention patterns coherent with immediately preceding tokens.
+However, they lack long context dependencies, failing to address the global attention crucial for long-context tasks.
+Moreover, a notable misalignment exists between the model response and human-written supervision.
+Consequently, this leads to inaccuracies when using human responses to compute attention values and gradients during profiling.
 For example, given the same question, a human might answer ‘Blue’, while the model could generate ‘The blue color’.
-Using the human answer for supervision, attention influence is inaccurately quantified based on probability shift for predicting ‘Blue’; this diverges from the objective of maintaining crucial attention for the original model prediction, ‘The’.
+When using the human answer for supervision, attention influence is inaccurately quantified based on the probability shift for predicting ‘Blue’; this diverges from the objective of maintaining crucial attention for the original model prediction, ‘The’.
 These inconsistencies arise from various factors, including mismatched positions, tones, and synonyms.
@@
 MoA enhances the calibration dataset by integrating long-range dependencies and model alignment.
-Specifically, we utilize the long-contextual MultiNews dataset (Fabbri et al., 2019), which includes summaries heavily dependent on long-range content.
+Specifically, we utilize the long-context MultiNews dataset (Fabbri2019MultiNews), which includes summaries that depend heavily on long-range content.
 The summaries are generated by the original dense model and serve as supervision.
-Compared to current approaches that adopt human responses as the reference to calculate the loss, using the responses generated by the original model as the supervision can facilitate accurate influence profiling, thus benefiting the compression results.
+Compared to current approaches that adopt human responses as the reference to calculate the cross-entropy loss $L$, using the responses generated by the original model as the supervision facilitates accurate profiling, thus benefiting the MoA configuration search.
 Approach Comparison.
-We evaluate our design’s effectiveness by varying dataset choices, supervision types, and summary references, while standardizing data item count and length to 50 and 8k words, respectively.
-Additional setups and evaluations are in Appendices A and B.3.1.
-We show the importance of long-range dependencies by comparing the MoA compression plan generated with different datasets and supervisory methods.
-In Table 2, RedPajama (Computer, 2023) represents the general language modeling dataset, while MultiNews (Fabbri et al., 2019) highlights long-range contexts by aggregating multiple documents on a single incident.
-Additionally, each MultiNews item includes a human-written summary, providing even stronger long-range dependencies and better performance.
-Calculating loss on the summary of MultiNews leads to significantly better performance, with a 60% increase in retrieval accuracy and a 0.98 decrease in perplexity.
+We validate our design with varied dataset choices, supervision types, and summary references, while standardizing the data item count and length to 50 and 8k tokens, respectively.
+Detailed setups and evaluations are in Appendices B and C.3.1.
+We show the importance of long-range dependencies by comparing the MoA configuration generated with different datasets and supervisory methods.
+In Table 2, RedPajama (together2023redpajama) represents the general language modeling dataset, while MultiNews (Fabbri2019MultiNews) highlights long-range contexts by aggregating multiple documents on a single incident.
+MultiNews additionally provides human-written summaries, enhancing long-range dependencies.
+Using MultiNews summaries for loss calculation significantly improves retrieval accuracy by 60% and reduces perplexity by 0.98.
 Furthermore, using summaries generated by the original dense model as supervision promotes higher alignment between its own attention patterns and the text supervision.
@@
 6 Experiment
-6.1 Setups
-We brief the experiment setups here, with more details in Appendix A.
-Baselines.
-We compare MoA with state-of-the-art static and dynamic sparse attention methods, including StreamingLLM (Xiao et al., 2024c), InfLLM (Xiao et al., 2024a) and H2O (Zhang et al., 2023).
-We define the density of an LLM as the ratio of the average in-memory KV-Cache length to the sequence length during the sparse decode stage.
-Notably, in MoA and StreamingLLM, KV-Cache length equals the attention span during the sparse prefill stage.
-In contrast, H2O use dense prefill.
-Besides, H2O and InfLLM require additional computations to dynamically determine the KV-Cache.
-Models and Benchmarks.
-We evaluate on vicuna-{7b, 13b}-v1.5-16k models (Chiang et al., 2023) from LMSys and Llama-3-{8b, 70b}-Instruct-262k models (AI, 2024) from Gradient AI.
-For long-context retrieval, we use LongEval (Li et al., 2023) to test key-value retrieval accuracy with 100 data items per length level.
-For long-context understanding, we use LV-Eval (Yuan et al., 2024) and LongBench (Bai et al., 2023), which include 11 and 13 sub-datasets, respectively.
-For coherence testing, we measure perplexity on four long datasets (Dasigi et al., 2021; Fabbri et al., 2019; Li & Roth, 2002; Hovy et al., 2001; Mohler et al., 2016) with diverse tasks.
-Unless otherwise specified, performance experiments are restricted to eight A100-80GB GPUs over a 24-hour period, with OOM (Out-Of-Memory) and OOT (Out-Of-Time) conditions noted.
-Efficiency experiments measure the decode throughput on a single A100-80GB GPU at maximum batch sizes of respective methods.
-MoA Settings.
-We restrict the number of distinct rules to at most two per model layer to ensure inference-time efficiency.
-We profile MoA on MultiNews (Fabbri et al., 2019) with model summaries at 2k, 4k, and 8k lengths.
-The optimal compression plan is selected with the validation dataset at 12k.
-Each model uses the same plan across all benchmarks and lengths.
-The models are not fine-tuned.
-6.2 Accuracy-Throughput Trade-off
-Figure 4 shows that MoA advances the Pareto Front in context retrieval accuracy and decode throughput across varied densities.
-At the same densities, MoA notably enhances throughput by 1.6×1.6$\times$1.6 × to 18.1×18.1$\times$18.1 × compared to H2O and InfLLM, due to its efficient static attention design.
-The throughput even outperforms StreamingLLM thanks to our customized GPU kernel.
-Additionally, MoA achieves notably higher retrieval accuracies across all densities.
-We conduct extensive evaluations for MoA ’s performance and efficiency on various benchmarks across context lengths from 4k to 256k and model sizes ranging from 7B to 70B in subsequent sections.
-6.3 Performance
-Figure 4: Accuracy-throughput trade-offs of various attention methods at different densities, tested on Vicuna-7B with 8k input length using one A100-80GB GPU on the LongEval dataset.
+Figure 4: Accuracy-throughput trade-offs of seven attention methods at different densities, tested on Vicuna-7B with 8k input length using one A100-80GB GPU on the LongEval dataset.
 Retrieval Acc. PPL
@@
 Evaluations are done with retrieval accuracy and perplexity.
-Table 4: Comparative analysis of retrieval accuracy, LV-Eval scores, LongBench scores, and perplexity for various models with different attention methods.
-All sparse methods employ 50% density in decode stage.
-H2O uses dense prefill, while StreamingLLM, InfLLM and MoA use sparse prefill.
-InfLLM for 70B model is excluded due to OOT issues.
+6.1 Setups
+We briefly summarize the experimental setup here; more details appear in Appendix B.
+Baselines.
+We compare MoA with state-of-the-art static and dynamic efficient attention methods, including StreamingLLM (xiao2023streamingLLM), InfLLM (xiao2024infllm) and H2O (Zhang2023H2O).
+We define the density as the ratio of the average in-memory KV-Cache length to sequence length during decoding.
+Notably, MoA and StreamingLLM use efficient prefill, while H2O and InfLLM use the original dense prefill, with additional computations to dynamically determine the KV-Cache for decoding.
+Models and Benchmarks.
+We use vicuna-{7b, 13b}-v1.5-16k models (vicuna2023) from LMSys and Llama-3-{8b, 70b}-Instruct-262k models (meta2024llama3) from Gradient AI.
+We test long-context retrieval, understanding, and coherence abilities, using LongEval (lmsys2023longeval) retrieval accuracy, LV-Eval (yuan2024lveval), and LongBench (bai2023longbench) scores, as well as average perplexity across four datasets (Dasigi2021Qasper; Fabbri2019MultiNews; li-roth-2002-trec1; hovy-etal-2001-trec2; lcc).
+Performance experiments are restricted to eight A100-80GB GPUs over a 24-hour period, with Out-Of-Memory (OOM) and Out-Of-Time (OOT) conditions noted.
+MoA Settings.
+We restrict the number of distinct rules to at most two per model layer to ensure inference-time efficiency.
+We profile MoA on MultiNews (Fabbri2019MultiNews) with model summaries at 2k, 4k, and 8k lengths.
+The optimal MoA configuration is selected using the validation dataset at 12k.
+Each model uses the same plan across all benchmarks and lengths.
+The models are not fine-tuned.
+6.2 Accuracy-Throughput Trade-off
+[…]
+6.3 Performance
 Retrieve Acc. $\uparrow$ LV-Eval $\uparrow$ LongBench $\uparrow$ PPL $\downarrow$
@@
 MoA 1.00 1.00 0.94 23.65 47.79 3.75
-MoA outperforms state-of-the-art sparse attention methods across various model sizes and benchmarks, achieving comparable performance to the original dense model at 50% density.
+Table 4: Comparative analysis of retrieval accuracy, LV-Eval scores, LongBench scores, and perplexity for various models with different attention methods.
+All methods employ 50% density in decode stage.
+H2O uses dense prefill; StreamingLLM, InfLLM and MoA use sparse prefill.
+InfLLM for 70B model is excluded due to OOT issues.
+MoA outperforms state-of-the-art efficient attention methods across various model sizes and benchmarks, achieving comparable performance to full-attention model at 50% density.
 Long-Context Retrieval.
-As shown in Table 4, MoA demonstrates a maximum of 8% relative accuracy drop (calculated as $\max\{1-\text{Acc.}_{\text{MoA}}/\text{Acc.}_{\text{Original}}\}$ across three lengths and LLMs), significantly less than the 87%, 58% and 44% for StreamingLLM, InfLLM and H2O.
-On average, the relative accuracy drop for MoA is under 1%, much less than others at 51%, 41% and 20%, respectively.
+As shown in Table 4, MoA demonstrates a maximum of 8% relative accuracy drop (calculated as $\max\{1-\text{Acc.}_{\text{MoA}}/\text{Acc.}_{\text{Original}}\}$ across three lengths and LLMs), significantly lower than StreamingLLM (87%), InfLLM (58%), and H2O (44%).
+On average, MoA ’s relative accuracy drop is below 1%, substantially better than the other methods (51%, 41%, and 20%).
 Figure 5(a) shows that MoA retains over 90% retrieval accuracy up to 60k lengths, equaling the dense model’s effective context length.
@@
 In contrast, the effective context lengths for H2O, InfLLM, and StreamingLLM are only 8k, <4k, and <4k, respectively.
-Appendix B.1.2 shows that MoA extends its effective context to approximately 3.9×3.9$\times$3.9 × the average KV-Cache length.
+Appendix C.1.2 shows that MoA extends its effective context to approximately $3.9\times$ the average KV-Cache length.
 Long-Context Understanding.
-As shown in Table 4, MoA minimizes the maximum relative performance drop in LV-Eval and LongBench benchmarks to only 5% and 3%, respectively—much lower than the 36% and 18% experienced by StreamingLLM.
-H2O and InfLLM show maximum relative drops of 9%-17% and 3%-5% with higher efficiency costs.
-Similar trends show in perplexity tests, where MoA maintains less than 1% relative perplexity increase, while others exhibit 4%-13% increases.
-This trend holds for other densities, as shown in Appendices B.1.1 and B.1.3.
-Figure 9 and Table 8 further details the score with different tasks.
-MoA achieves comprehensive performance comparable to the original dense model, as well as H2O that requires higher efficiency cost.
-In contrast, StreamingLLM and InfLLM display inconsistent performance: it sometimes surpasses the original model in some tasks, while suffering noticeable degradation in others.
-Longer-Context Generalization.
-By compressing within 12k lengths, MoA effectively generalizes to lengths of 32k-256k, as shown in Figure 5(b).
-At the extended lengths, MoA outperforms both InfLLM and StreamingLLM by 1.93.3×1.9-3.3$\times$1.9 - 3.3 × in retrieval accuracy and 1.21.4×1.2-1.4$\times$1.2 - 1.4 × in LV-Eval scores, demonstrating comparable performance to the original dense model.
+Table 4 indicates that MoA reduces the maximum relative performance drop in LV-Eval and LongBench benchmarks to 5% and 3%, significantly less than StreamingLLM’s 36% and 18%.
+H2O and InfLLM incur maximum relative drops of 9%-17% and 3%-5%, respectively, but with greater efficiency costs.
+Figure 10 and Table 8 further show MoA’s consistent performance across subtasks, in contrast to the inconsistent performance of the baselines.
+In perplexity tests, MoA maintains less than a 1% increase in relative perplexity, whereas other methods see increases of 4%-13%.
+These trends persist at other densities (Appendices C.1.1, C.1.3).
+[…]
 (a) Retrieval accuracy and the effective context length (arrow).
@@
 (b) Retrieval accuracy and LV-Eval score at longer lengths
-Figure 5: Comparative analysis at extended sequence lengths with different attention methods using Llama3-8B model. All methods employ 50% density in both prefill and decode stages.
+Figure 5: Experiments at extended sequence lengths with different attention methods using Llama3-8B model. All methods employ 50% density in both prefill and decode stages.
 Ablation Study.
-We evaluate the performance impact of different sparse mask search spaces in Table 3.
-Starting with a basic uniform mask, we observe significant enhancements by sequentially introducing heterogeneity: layers first, then heads, and finally elastic rules.
+We evaluate the performance impact of different sliding-window attention search spaces in Table 3.
+Starting with a basic uniform sliding-window method, we observe significant enhancements by sequentially introducing heterogeneity: layers first, then heads, and finally elastic rules.
 6.4 Efficiency
-Table 5: Runtime efficiency of different methods on Vicuna-7B and 13B models.
-Efficiency improvements of MoA are ablated with four factors.
-All sparse attention methods use 50% density.
-Decode throughput (tokens per second) evaluated at the maximum batch capacity of an A100-80GB GPU.
 4k 8k 16k
@@
 +Kernel (=MoA) 28 550.9 14 267.6 7 132.3
-MoA shows high runtime efficiency with a manageable one-time compression overhead.
-Runtime Efficiency.
-Table 5 compares the runtime efficiency of MoA over various attention methods and LLM frameworks, with the ablation of efficiency improvements brought by each design factor of MoA.
-At 50% density, MoA boosts the decode throughput by 6.6$\times$ to 8.2$\times$ compared to FlashAttention2.
-It outperforms H2O and InfLLM with 1.2×1.2$\times$1.2 × to 4.0×4.0$\times$4.0 × decode throughput improvements.
-Even compared to the highly system-level optimized vLLM framework (Kwon et al., 2023), MoA still achieves a 1.7×1.7$\times$1.7 × to 1.9×1.9$\times$1.9 × throughput increase.
-MoA also reduces total GPU memory by 1.2×1.2$\times$1.2 × to 1.4×1.4$\times$1.4 ×, are detailed in Appendix B.2.
-This throughput gain results from four main factors: static-sized KV-Cache during generation(3.0$\approx$ 3.0$\times$ 3.0 ×); reduced attention computations due to sparsity (1.5$\approx$ 1.5$\times$ 1.5 ×); increased batch sizes enabled by smaller KV-Cache memory (1.4$\approx$ 1.4$\times$ 1.4 ×); and our CUDA-implemented GPU kernel for MoA heterogeneous attention (1.2$\approx$ 1.2$\times$ 1.2 ×).
-Compression Pipeline Efficiency.
-MoA completes the automatic compression pipeline for the Vicuna-7B and 13B models within two hours.
-For the larger Llama3-70B model, the process requires 8.5 hours of real-time and 34.7 hours of GPU time.
-Details for each compression step overhead are in Appendix B.2.2.
-6.5 Rules Discovered by MoA
-We investigate MoA’s elastic rules for each head.
-As shown in Figure 10, masks in the initial and middle layers exhibit high density, aligning with the conclusions from previous research on LLM’s intrinsic dimensions (Valeriani et al., 2023) and layer sensitivities (Yuan et al., 2023).
-Conversely, in the final layers, most heads require low density, while few need high density.
-Figure 11 shows that layers with lower average mask density typically display a wider range of densities among heads, confirming the need for heterogeneity within the same layer.
-Further details and quantified insights are presented in Appendix C.
-7 Conclusion and Future Work
-MoA automates the selection of heterogeneous elastic masks for each attention head and input length, significantly extending the effective context length of LLMs by 3.9×3.9$\times$3.9 ×.
-It enhances retrieval accuracy by 1.5×1.5$\times$1.5 × to 7.1×7.1$\times$7.1 × over uniform sparse attention method and increases throughput to over 6×6$\times$6 × at 50% average density, maintaining performance on par with dense models in rigorous benchmarks.
-Limitations and Future Work.
-Under an extremely low-density budget, MoA fails to maintain good performance.
-Designing a dynamic MoA method has the potential to address this issue, which we leave for future work.
-Using non-linear elastic rules with bounded attention spans is also worth exploring.
-Additionally, MoA’s profiling method can be adapted to evaluate the influence of weights and other activations, facilitating other compression methods such as weight and activation quantization (Lin et al., 2023; Xiao et al., 2024b; Li et al., 2024b).
-Acknowledgement
-This work was supported by National Natural Science Foundation of China (No.
-62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164), Tsinghua EE Xilinx AI Research Fund, and Beijing National Research Center for Information Science and Technology (BNRist).
-Appendix A Detailed Experiment Setup
-A.1 Main Setup
+Table 5: Decode throughput (tokens/s) of Vicuna-7B,13B, evaluated at the maximum batch capacity of an A100-80GB GPU.
+All efficient attention methods use 50% density.
+Efficiency improvements of MoA are ablated with four factors.
+MoA shows high runtime efficiency with a manageable one-time search overhead.
+[…]
+Search Pipeline.
+MoA completes the automatic configuration search for the Vicuna-7B and 13B models within two hours.
+For the larger Llama3-70B model, the process requires 8.5 hours of wall time and 34.7 hours of GPU time.
+See Appendix C.2.4 for more details.
+7 Conclusion
+[…]
+Acknowledgments
+This work was supported by the National Natural Science Foundation of China (No.
+62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164), the Tsinghua EE Xilinx AI Research Fund, and the Beijing National Research Center for Information Science and Technology (BNRist).
+We thank Yi Ge for valuable discussions.
+Appendix A Notations and Definitions
+To clarify the terminology and concepts introduced in our approach, we present the key notations and definitions used throughout the paper.
+A.1 Attention Span
+The attention span of each attention head is defined as the sum of a sliding-window length and a fixed prefix of always-visible tokens (attention sink). In the attention mask:
+  • Sliding Window: A contiguous diagonal region indicating recent tokens each head can attend to.
+  • Prefix: The first 64 tokens of the input sequence, which remain visible to all attention heads regardless of the sliding-window length.
+A.2 Elastic Rule
+An elastic rule is a function $f(\cdot)$ parameterized by $\alpha$ and $\beta$, specifying how the attention span scales with the input sequence length $N$ for a single head.
+Formally, the attention span $\mathcal{S}$ given by an elastic rule is:
+$\mathcal{S}(N;\alpha,\beta)=\alpha\times N+\beta$ (6)
+Different attention heads employ different elastic rules by adopting different sets of $\alpha$ and $\beta$, resulting in heterogeneous attention spans optimized per head.
+A.3 MoA Configuration
+An MoA configuration is the collection of elastic rules of all attention heads of all layers in a model.
+A.4 Attention Mask
+A.4.1 Visualization
+An attention mask is a binary mask applied to each head individually to gate its attention computation.
+It is defined by the elastic rule and indicates which tokens are visible (unmasked) or pruned (masked).
+In Figure 3, each attention mask has three visual elements:
+Visual Element Meaning Contribution to Span
+Vertical purple stripe Initial unmasked tokens (prefix) 64 tokens (fixed)
+Diagonal purple band Sliding window region Window length
+White cells Pruned tokens 0
+As shown in the figure, each grid represents a candidate attention mask:
+  • Vertically aligned masks represent masks generated from different elastic rules at the same input length.
+  • Horizontally aligned masks represent masks generated from the same elastic rule at different input lengths.
+A.4.2 Application of Attention Masks
+Attention masks for all heads are applied in parallel—each mask independently gates attention computation and the corresponding key-value (KV) cache entries.
+There is no overlap or sequential application between masks across different heads.
+A.5 Correlations Between Notions
+The following list summarizes the correlations between notions. Specifically:
+  1. 1.
+    MoA configuration and elastic rule: An MoA configuration is the elastic rules of all attention heads of all layers in an LLM.
+  2. 2.
+    Elastic rule and attention Span: Given an elastic rule parameterized by $(\alpha,\beta)$ and input length $N$, the attention span is computed using Equation 6.
+  3. 3.
+    Attention span and attention mask: The computed span directly determines the shape and extent of the sliding-window portion of the attention mask.
+Appendix B Detailed Experiment Setup
+B.1 Main Setup
 Baselines.
-In the setup for our experiment, we adhere to specific configurations outlined in the respective papers.
-In the case of In StreamingLLM (Xiao et al., 2024c), the initial four tokens remain unmasked, serving as the attention sink, except for the 70b model in Table 4 and the super long setting in Figure 5, where we use 64 tokens as the attention sink.
-For InfLLM (Xiao et al., 2024a), we adhere to the original configuration by maintaining the same local window size and selected memory size, using 128 initial tokens as specified in their setup.
-For H2O (Zhang et al., 2023), we ensure the same number of heavy hitter tokens and recent tokens.
-Note that H2O uses dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction.
-StreamingLLM, InfLLM and MoA use sparse prefill.
+In our experimental setup, we adhere to the specific configurations outlined in the respective papers.
+In the case of StreamingLLM (xiao2023streamingLLM), the initial four tokens remain unmasked, serving as the attention sink, except for the 70b model in Table 4 and the super long setting in Figure 5, where we use 64 tokens as the attention sink.
+For InfLLM (xiao2024infllm), we adhere to the original configuration by maintaining the same local window size and selected memory size, using 128 initial tokens as specified in their setup.
+For H2O (Zhang2023H2O), we ensure the same number of heavy hitter tokens and recent tokens.
+Note that H2O uses original dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction.
+StreamingLLM, InfLLM and MoA use efficient prefill.
+Notably, in MoA and StreamingLLM, the KV-Cache length equals the attention span during the efficient prefill stage.
+In contrast, H2O uses the original dense prefill.
+Besides, H2O and InfLLM require additional computations to dynamically determine the KV-Cache.
 Models and Benchmarks.
-Since vicuna-7b-v1.5-16k and vicuna-13b-v1.5-16k (Chiang et al., 2023) can only take in 16k context length, we use the 16k split of LV-Eval benchmark (Yuan et al., 2024), truncating the input to 15500 for model input in Table 4.
-For the LongBench benchmark (Bai et al., 2023), we use the LongBench-E split, which features a balanced number of data items at every length level.
+For long-context retrieval, we use LongEval (lmsys2023longeval) to test key-value retrieval accuracy with 100 data items per length level.
+For long-context understanding, we use LV-Eval (yuan2024lveval) and LongBench (bai2023longbench), which include 11 and 13 sub-datasets, respectively.
+Since vicuna-7b-v1.5-16k and vicuna-13b-v1.5-16k (vicuna2023) can only take in 16k context length, we use the 16k split of LV-Eval benchmark (yuan2024lveval), truncating the input to 15500 for model input in Table 4.
+For the LongBench benchmark (bai2023longbench), we use the LongBench-E split, which features a balanced number of data items at every length level.
 The LongBench dataset is segmented into ranges of 0-4k, 4-8k, and 8k+ tokens.
 We test each split using the input length truncation thresholds of 3,500, 7,500, and 15,500 tokens, respectively.
+Efficiency experiments measure the decode throughput on a single A100-80GB GPU at maximum batch sizes of respective methods.
 Perplexity Evaluation.
-We construct a comprehensive yet concise test set by sampling $50\times 4$ data items for each length level from the test split of four long-context understanding datasets: Qasper (Dasigi et al., 2021), MultiNew  (Fabbri et al., 2019), TREC (Li & Roth, 2002; Hovy et al., 2001) and LCC (Mohler et al., 2016), representing the question answering, summarization, few-shot learning, and code completion abilities of the LLM.
+We construct a comprehensive yet concise test set by sampling $50\times 4$ data items for each length level from the test split of four long-context understanding datasets: Qasper (Dasigi2021Qasper), MultiNews (Fabbri2019MultiNews), TREC (li-roth-2002-trec1; hovy-etal-2001-trec2) and LCC (lcc), representing the question answering, summarization, few-shot learning, and code completion abilities of the LLM.
 Following LongBench, the data items are organized as question-answer pairs.
@@
 Validation Dataset.
-The validation dataset is used to select the optimal compression plan among the Pareto front solutions during the optimization step.
+The validation dataset is used to select the MoA configuration among the Pareto front solutions during the optimization step.
 The validation dataset is similarly constructed as the perplexity test dataset, but on the respective validation split of the datasets.
-$50\times 4$ data items are sampled from the same four long-context understanding datasets: Qasper (Dasigi et al., 2021), MultiNew  (Fabbri et al., 2019), TREC (Li & Roth, 2002; Hovy et al., 2001) and LCC (Mohler et al., 2016).
-The additional $50$ data items from the LongEval (Li et al., 2023) dataset are also added to validate the retrieval ability.
+$50\times 4$ data items are sampled from the same four long-context understanding datasets: Qasper (Dasigi2021Qasper), MultiNews (Fabbri2019MultiNews), TREC (li-roth-2002-trec1; hovy-etal-2001-trec2) and LCC (lcc).
+The additional $50$ data items from the LongEval (lmsys2023longeval) dataset are also added to validate the retrieval ability.
 For the datasets that do not contain the validation split, namely TREC, MultiNews and LCC, we sample from the test split and ensure different data items with the perplexity evaluation dataset.
 MoA Settings.
-MoA uses the block sparse attention pattern with a block size of 64, where each grid depicted in Figure 3(a) represents a block.
-The first block of tokens is not masked as the attention sink.
-For the profile stage, we use the MultiNews (Fabbri et al., 2019) calibration dataset with model response as supervision, as described in Section 5.
-We use $50\times 3$ data items at 2k, 4k, 8k lengths.
-The data items are padded to their corresponding length level in order to ensure a unified shape of attention influence tensors for each length level.
-We adopt block granularity during the profiling stage, calculating the average attention influence within each block to represent the block’s overall influence.
+MoA uses the block sliding-window attention pattern with a block size of 64, where each grid depicted in Figure 3(a) represents a block.
+The first block of tokens is not masked and serves as the attention sink.
+For the profiling stage, we use the MultiNews (Fabbri2019MultiNews) calibration dataset with the model response as supervision, as described in Section 5.
+We use $50\times 3$ data items at 2k, 4k, and 8k lengths.
+The data items are padded to their corresponding length level to ensure a unified shape of attention-influence tensors for each level.
+We adopt block granularity during profiling, calculating the average attention influence within each block to represent the block’s overall influence.
+For hyperparameter search space $\alpha$ and $\beta$, we use 6 values for $\alpha$ and 9 values for $\beta$, creating a search space of 54 pairs for each attention head.
+$\alpha$ is uniformly sampled from the range [2048,8192][-2048,8192], and $\beta$ is uniformly sampled from $[0,1]$.
+The resulting attention span lengths are clipped to the range between 0 and the current input length.
 The optimization is done with the multi-objective optimization at the same set of lengths.
@@
 Among the Pareto front solutions, we select the one with the lowest perplexity on the validation dataset of length 12k.
-A.2 Efficiency Experiment Setup
+B.2 Efficiency Experiment Setup
 We test the efficiency of different frameworks using a single NVIDIA A100-SXM4-80GB GPU.
@@
 For MoA, the implementation is based on HuggingFace Transformers.
-During the prefill stage, we use the sparse CUDA kernel designed by us with block size 64.
+During the prefill stage, we use the sliding-window attention CUDA kernel that we designed with a block size of 64.
 During the decode stage, we modify the KV-Cache implementation to support our heterogeneous elastic rules.
-Thanks to our fixed sliding-window span during the decode stage, we simply replace the old KV-Cache that exceeds the span with the latest KV-Cache.
-Our custom decoding CUDA kernel then handles KV-Cache with varying lengths across different attention heads during the decoding process.
-For H2O, we use its official efficient implementation, which is based on Flexgen (Sheng et al., 2023).
+Thanks to our fixed sliding-window span during decoding, we simply replace the old KV-Cache that exceeds the span with the latest KV-Cache.
+Our custom decoding CUDA kernel then handles KV-Cache with varying lengths across different attention heads during decoding.
+For H2O, we use its official efficient implementation, which is based on FlexGen (sheng2023flexgen).
 Note that H2O uses dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction, which requires the attention matrix to be explicitly calculated.
-It makes H2O’s prefill stage currently incompatible with kernel optimizations like FlashAttention.
-Therefore, H2O is easy to get OOM (Out-Of-Memory) with large prefill length and increased batch size.
+This makes H2O’s prefill stage currently incompatible with kernel optimizations like FlashAttention.
+Therefore, H2O is prone to OOM (Out-Of-Memory) with large prefill lengths and increased batch sizes.
 In our efficiency tests across all frameworks, we implemented a simple optimization at the language modeling head (lm head) during the prefill stage.
@@
 This selective computation avoids generating these probabilities for preceding tokens, substantially reducing both computational overhead and memory usage.
-We also set the environment variable PYTORCH_CUDA_ALLOC_CONF to be expandable_segments:True for Hugginface and MoA to mitigate memory fragmentation, allowing larger inference batch size.
+We also set the environment variable PYTORCH_CUDA_ALLOC_CONF to expandable_segments:True for HuggingFace and MoA to mitigate memory fragmentation, allowing larger inference batch sizes.
 Following the performance experiments, we use Vicuna-7B and Vicuna-13B for efficiency tests whenever possible.
-However, the official efficient implementation of H2O based on Flexgen only supports OPT (Zhang et al., 2022).
+However, the official efficient implementation of H2O based on Flexgen only supports OPT (zhang2205opt).
 Therefore, we use OPT-6.7b and OPT-13b models for H2O in Table 11 for comparison.
-A.3 Ablation Study Setup
+B.3 Ablation Study Setup
 In the ablation study in Table 2 and Table 3, we use 25% density instead of the 50% used in the main experiment in Table 4.
@@
 In the calibration dataset experiments in Table 2, we intentionally exclude the influence of the validation dataset.
-We avoid using the validation dataset by profile and optimize solely at 8k length, reducing the multi-objective optimization problem to a single-objective one with only one optimal compression plan instead of a set of Pareto fronts.
-Appendix B Additional Experiment Results
-B.1 Performance
-B.1.1 Overall Performance
-Table 6: Comparative analysis of retrieval accuracy, LV-Eval scores, LongBench scores, and perplexity for various models with different attention methods.
-All methods employ 75% density in both prefill and decode stages.
+We avoid using the validation dataset by profile and optimize solely at 8k length, reducing the multi-objective optimization problem to a single-objective one with only one optimal MoA configuration instead of a set of Pareto fronts.
+B.4 Input Format And Examples
+We list the prompt format and input examples used in our primary experiments and datasets. Dashed lines are included only for illustration clarity and are not part of the texts given to the LLMs.
+Format 1.
+LongEval Below is a record of lines I want you to remember.
+Each line begins with ’line <line index>’ and contains a ’<REGISTER_CONTENT>’ at the end of the line as a numerical value.
+For each line index, memorize its corresponding <REGISTER_CONTENT>.
+At the end of the record, I will ask you to retrieve the corresponding <REGISTER_CONTENT> of a certain line index.
+Now the record start: line delightful-incandescence: REGISTER_CONTENT is <19147> line cloistered-presence: REGISTER_CONTENT is <8862> Now the record is over.
+Tell me what is the <REGISTER_CONTENT> in line cloistered-presence?
+I need the number.
+Format B.4 illustrates the input format for the LongEval (lmsys2023longeval) retrieval benchmark.
+The instruction indicating which line to retrieve is provided after a lengthy context containing massive lines of register contents to remember.
+Format 2.
+Needle-In-A-Haystack (NIAH) People who are powerful but uncharismatic will tend to be disliked.
+Their power makes them a target for criticism that they don’t have the charisma to disarm.
+That was Hillary Clinton’s problem.
+The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day.
+It also tends to be a problem for any CEO who is more of a builder than a schmoozer.
+What is the best thing to do in San Francisco?
+Format B.4 depicts the input format for another common retrieval benchmark, Needle-In-A-Haystack (NIAH) (kamradt2024NIAH).
+The NIAH test comprises a single "needle" sentence that commonly does not fit into an irrelevant context.
+The model tries to answer the question based on this needle sentence.
+Format 3. MultiNews Calibration Dataset You are given several news passages. Write a one-page summary of all news. <News1> <News2> Now, write a one-page summary of all the news. Summarization
+Format B.4 demonstrates the input format for our calibration dataset.
+The long-context MultiNews dataset (Fabbri2019MultiNews) consists of multiple news documents.
+The context includes a prompt instructing the original dense model to generate a summarization for these news articles, reflecting long-range dependencies and model alignment.
+The generated summarization serves as supervision during the cross-entropy loss calculation at the profiling stage.
+Appendix C Additional Experiment Results
+C.1 Performance
+C.1.1 Overall Performance
 Retrieve Acc. $\uparrow$ LV-Eval $\uparrow$ LongBench $\uparrow$ PPL $\downarrow$
@@
 MoA 0.99 1.00 0.93 15.61 43.51 43.16 43.58 4.53
+Table 6: Comparative analysis of retrieval accuracy, LV-Eval scores, LongBench scores, and perplexity for various models with different attention methods.
+All methods employ 75% density in both prefill and decode stages.
 Table 6 shows the overall performance of MoA at a higher density of 75%.
@@
 The progressive change of performance with respect to different densities is also shown in Figure 7(b) and Figure 9
-B.1.2 Long-Context Retrieval
+C.1.2 Long-Context Retrieval
 Figure 6: Retrieval accuracy of Vicuna-7B model using different attention methods across varying attention spans and input lengths.
@@
 (a) Varies input lengths and densities to show effective context lengths across attention spans, (b) Set input length at 8k and show retrieval accuracy across different densities.
-We conduct a detailed experiment to test the retrieval ability of different attention methods across various attention spans and input lengths.
+LongEval Retrieval.
+We conduct a detailed experiment to test the retrieval ability of different attention methods across various attention spans and input lengths with the LongEval (lmsys2023longeval) dataset.
 Figure 6 shows the detailed data for effective context length calculation.
 As shown in the figure, StreamingLLM can hardly maintain retrieval accuracy when the input length is beyond the attention span, while MoA can effectively extend the effective context length.
-Following previous work (Chen et al., 2023; Tworkowski et al., 2023), we quantify effective context length as the maximum input length where retrieval accuracy remains above a 90% threshold.
+Following previous work (Chen2023PI; Tworkowski2023FocusedTrans), we quantify effective context length as the maximum input length where retrieval accuracy remains above a 90% threshold.
 As shown in Figure 7(a), StreamingLLM and H2O achieve effective context lengths of no more than 2k tokens beyond their attention spans.
-In contrast, MoA expands its effective context length to approximately 3.9×3.9$\times$3.9 × its attention span before reaching up to the 12k limit of the original model.
+In contrast, MoA expands its effective context length to approximately $3.9\times$ its attention span before reaching up to the 12k limit of the original model.
 Figure 7(b) further shows that at a fixed input length of 8k, MoA reaches over 0.9 retrieval accuracy with just 25% density, whereas StreamingLLM and H2O require 100% and 75% density, respectively.
-B.1.3 Long-Context Understanding
+Figure 8: The Needle-In-A-Haystack (NIAH) retrieval accuracy using different attention methods across 8k to 256k input lengths on Llama-3-8B model.
+All efficient attention methods employ a 50% density.
+Needle-In-A-Haystack (NIAH) Retrieval.
+We also conduct the retrieval task using the Needle-In-A-Haystack (NIAH) dataset (kamradt2024NIAH).
+As shown in Figure 8, MoA achieves perfect retrieval accuracy across input lengths ranging from 8k to 256k.
+In comparison, StreamingLLM demonstrates a limited effective context length, while InfLLM exhibits reduced retrieval accuracy within 64k input lengths.
+Notably, H2O and InfLLM are unable to complete tests at extreme lengths due to Out-Of-Memory and Out-Of-Time errors.
+These findings align with the results observed in the LongEval benchmark throughout the paper.
+C.1.3 Long-Context Understanding
 Figure 9: LV-Eval score of StreamingLLM and MoA at various densities on Vicuna-7B model.
-We conduct experiments with various densities on the LV-Eval benchmark (Yuan et al., 2024).
+We conduct experiments with various densities on the LV-Eval benchmark (yuan2024lveval).
 As shown in Figure 9, MoA constantly outperforms the uniform static attention baseline StreamingLLM at various densities, demonstrating the effectiveness of our heterogeneous elastic rules.
-Table 7: LongBench scores for various models with different attention methods. All methods employ 50% density in the decode stage.
 LongBench $\uparrow$
@@
 MoA 49.74 46.80 46.84
-Table 8: Performance comparison across different models and attention methods with the LV-Eval dataset. The numbers in brackets indicate the number of sub-datasets for the category.
+Table 7: LongBench scores for various models with different attention methods. All methods employ 50% density in the decode stage.
 Single-QA Multi-QA Retrieval
@@
 MoA 42.44 23.58 15.75 21.27 19.19
+Table 8: Performance comparison across different models and attention methods with the LV-Eval dataset. The numbers in brackets indicate the number of sub-datasets for the category.
 Figure 10: (a) LV-Eval and (b) LongBench scores for different attention methods at 50% density, tested on Vicuna-7B, 13B and Llama3-70B models. Scores normalized against the original dense model.
-We detailed the respective scores for LongBench and LV-Eval in Table 7 and Table 8. The number in the bracket of Table 8 indicates the number of sub-datasets for the category.
-B.2 Efficiency
-B.2.1 Throughput Breakdown
-Table 9: Efficiency analysis of different frameworks on 7B and 13B models. H2O and MoA use 50% density. GPU memory evaluated with batch sizes 8 (7B model) and 4 (13B model).
+We detailed the respective scores for LongBench and LV-Eval in Table 7 and Table 8.
+The number in the bracket of Table 8 indicates the number of sub-datasets for the category.
+MoA achieves comprehensive performance comparable to the original dense model, as well as H2O that requires higher efficiency cost.
+In contrast, StreamingLLM and InfLLM display inconsistent performance: it sometimes surpasses the original model in some tasks, while suffering noticeable degradation in others.
+C.1.4 Longer-Context Generalization
+Retrieve Acc. $\uparrow$
+Attention 32k 64k 128k 256k
+SnapKV 1.00 0.88 0.71 0.33
+PyramidKV 1.00 0.85 0.62 0.37
+MoA 1.00 0.92 0.83 0.46
+Table 9: Retrieval accuracy at longer lengths for more recent baselines, tested at 50% density.
+We compare the retrieval accuracy with more recent works SnapKV (li2024snapkv) and PyramidKV (cai2024pyramidkv) on context lengths of 32K to 256K.
+As shown in Table 9, MoA constantly outperforms the two latest baselines at longer contexts.
+C.1.5 Instruction-following Generation
+Attention Length-controlled Win Rate $\uparrow$ Standard Error
+Original 8.84 0.53
+H2O 9.66 0.55
+InfLLM 5.76 0.42
+StreamingLLM 7.96 0.49
+MoA 9.83 0.57
+Table 10: Length-controlled win rate and its standard error of Vicuna-7B with different attention mechanisms on AlpacaEval 2.0 benchmark.
+All efficient attention methods employ 50% density during decoding.
+We evaluate MoA ’s performance on general instruction-following tasks using the AlpacaEval 2.0 benchmark (alpacaEvalRepo; dubois2024alphacaEvalLength).
+Following the official setup, we compare the model’s output with gpt4_turbo using the standard weighted_alpaca_eval_gpt4_turbo evaluator, which leverages the gpt-4-1106-preview model.
+The benchmark consists of inputs and outputs with average lengths of approximately 50 and 450 tokens, respectively.
+To accommodate the short input lengths while maintaining a density of around 50% during generation, we set the expected total token length to 512 and adjust hyperparameters across all methods accordingly.
+Thanks to its elastic design, MoA employs the same configuration used in experiments with input lengths ranging from 4k to 256k.
+As shown in Table 10, MoA achieves the highest length-controlled win rate, outperforming both efficient attention baselines and the original model.
+C.2 Efficiency
+C.2.1 Memory and Throughput Breakdown
 Memory (GB)
@@
 MoA 32.0 39.6 55.0
-In this section, we explain the decode throughput breakdown for MoA, compared to the baseline comprising Huggingface with FlashAttention2.
-The observed increase in throughput, as shown in Table 5, primarily stems from three aspects:
+Table 11: Efficiency analysis of different frameworks on 7B and 13B models. H2O and MoA use 50% density. GPU memory evaluated with batch sizes 8 (7B model) and 4 (13B model).
+Table 11 highlights the memory efficiency of MoA compared to H2O and FlashAttention2 on 7B and 13B models.
+Notably, H2O runs into Out-Of-Memory (OOM) issues at longer input lengths.
+In contrast, MoA achieves a significant reduction in memory consumption, using $1.2$ to $1.4\times$ less memory compared to FlashAttenion2.
+We further explain the decode throughput breakdown in Table 5, compared to the baseline comprising Huggingface with FlashAttention2.
+The observed increase in throughput primarily stems from four aspects:
 Static KV-Cache. MoA only maintains the tokens within the span of each head, thereby preventing growth in the KV-Cache size. This strategy eliminates the need for additional memory allocation.
@@
 Increased Batch Size. With the reduced size of KV-Cache, MoA supports a larger batch size, contributing to the increase in throughput.
-B.2.2 Automatic Compression Pipeline Overhead
-Table 10: Compression overhead for various stages of MoA across models with differing parameter sizes, reported as the amount of GPU $\times$ latency, except when only one GPU is used.
-Larger models necessitate more GPUs due to model parallelism.
-All stages utilize GPUs, except for the Optimize stage, which uses the CPU.
+GPU Kernel Design. We customize MoA GPU kernel using CUDA to support heterogeneous attention patterns with high efficiency.
+C.2.2 Efficiency Results for Longer Input
+Min. Total Total Throughput
+Model Size Framework Attention #GPU Throughput Memory (GB) per GPU
+7B vLLM PagedAttention 2 30.2 142.0 15.1
+FlexGen H2O >>8 - OOM -
+HuggingFace InfLLM 1 6.1 47.7 6.1
+HuggingFace StreamingLLM 1 19.8 43.9 19.8
+HuggingFace FlashAttention2 2 4.3 85.6 2.2
+HuggingFace MoA 1 20.3 44.0 20.3
+13B vLLM PagedAttention 2 21.5 142.0 10.8
+FlexGen H2O >>8 - OOM -
+HuggingFace InfLLM 1 4.3 78.6 4.3
+HuggingFace StreamingLLM 1 14.0 64.6 14.0
+HuggingFace FlashAttention2 2 3.0 130.6 1.5
+HuggingFace MoA 1 14.7 63.4 14.7
+Table 12: Runtime efficiency at 128k input length across different methods on Vicuna-7B and 13B models.
+All efficient attention methods use 50% density.
+Decode throughput (tokens per second) is measured with a batch size of 1, using the minimum number of A100-80GB GPUs required for testing.
+H2O encounters OOM error with 8 GPUs.
+We evaluate the runtime efficiency of Vicuna-7B and 13B models at a 128k input length with a single batch size.
+Thanks to the reduced KV-Cache, MoA efficiently processes 128k input using only one A100 GPU, whereas FlashAttention2 and vLLM baselines require at least two GPUs to handle a single request.
+As shown in Table 12, MoA achieves a $4.7$-$4.9\times$ decode speedup compared to FlashAttention2, while using half the number of GPUs.
+Additionally, it demonstrates a $1.9$-$2.1\times$ reduction in GPU memory usage.
+Compared to vLLM, which utilizes tensor parallelism, MoA delivers $1.3$-$1.4\times$ higher throughput per GPU, alongside significant memory savings.
+C.2.3 Energy Efficiency Result
+[…]
+Energy Per Token (J) Power (W)
+Framework Attention 4k 8k 16k 4k 8k 16k
+Huggingface FlashAttention2 2.98 5.93 12.1 350 354 359
+Huggingface MoA 0.34 0.62 1.21 330 322 315
+Table 13: Per token energy consumption and average GPU power with different attention mechanisms at various sequence lengths, using Vicuna-7B on a single A100 GPU.
+C.2.4 Automatic Search Pipeline Overhead
 Stage 7B LLM 13B LLM 70B LLM
@@
 Total GPU Time 1h 5min 1h 45min 34h 40min
-Table 11: Progressive compression overhead for various stages of MoA with respect to different parameter sizes and calibration (validation) dataset sizes.
+Table 14: Search overhead for various stages of MoA across models with differing parameter sizes, reported as the amount of GPU $\times$ latency, except when only one GPU is used.
+Larger models necessitate more GPUs due to model parallelism.
+All stages utilize GPUs, except for the Optimize stage, which uses the CPU.
 Stage Complexity w.r.t parameter size Complexity w.r.t dataset size
@@
 Empirical Latency Almost Linear Linear
+Table 15: Overheads for various stages of MoA with respect to different parameter sizes and calibration (validation) dataset sizes.
 We present a detailed breakdown of the time usage of MoA pipeline.
@@
 On the Intel(R) Xeon(R) Platinum 8358 2.60 GHz CPU, the optimization concludes within approximately 25 minutes.
-Typically, this phase generates around 10 compression plans.
-Validating each one of the compression plans takes about 4 minutes, totaling around 40 minutes.
-We also show the progressive compression overhead for MoA in Table 11.
-B.3 Ablation Study
-B.3.1 Calibration Dataset
-Table 12: Performance comparison on various test sets, using different calibration sets.
-Tested on Vicuna-7B model.
-The result is tested with 50% density MoA on LongBench (Bai et al., 2023) 0-4k split.
+Typically, this phase generates around 10 MoA configurations.
+Validating each one of the configuration takes about 4 minutes, totaling around 40 minutes.
+We also show the overhead of each stages in Table 15.
+C.3 Ablation Study
+C.3.1 Calibration Dataset
 Test Score
@@
 TREC 27.3 (-1.3) 27.3 (-0.9) 55.0 (-1.0) 36.5 (-1.1)
+Table 16: Performance comparison on various test sets, using different calibration sets.
+Tested on Vicuna-7B model.
+The result is tested with 50% density MoA on LongBench (bai2023longbench) 0-4k split.
 In this section, we validate the robustness of our calibration dataset design principles.
-We select three sub-tasks and respective datasets from the LongBench benchmark, including Qasper (Dasigi et al., 2021), MultiNews (Fabbri et al., 2019), and TREC (Li & Roth, 2002; Hovy et al., 2001).
+We select three sub-tasks and respective datasets from the LongBench benchmark, including Qasper (Dasigi2021Qasper), MultiNews (Fabbri2019MultiNews), and TREC (li-roth-2002-trec1; hovy-etal-2001-trec2).
 We use their training set to construct the calibration dataset, and use their respective test set in LongBench to calculate the score.
@@
 In contrast, the RedPajama dataset without long-range dependency and model alignment shows large variance on various test sets.
-It also differs from the performance of the original dense model, which may incur unexpected behaviors after compression.
+It also differs from the performance of the original dense model, which may incur unexpected behaviors after the MoA search pipeline.
 Note that though all datasets exhibit long dependency, the questions in the TREC dataset can be answered without long context.
@@
 It may contribute to a high score on the TREC test set with the RedPajama calibration dataset.
-Appendix C Compression Plan Analysis
-C.1 Statistics on Rules Discovered by MoA
+Appendix D MoA Configuration Analysis
+D.1 Overview
+We investigate MoA’s elastic rules for each head.
+As shown in Figure 11, masks in the initial and middle layers exhibit high density, aligning with conclusions from previous research on LLMs’ intrinsic dimensions (Valeriani2023hiddenRepresent) and layer sensitivities (Yuan2023ASVD).
+Conversely, in the final layers, most heads require low density, while few need high density.
+Figure 12 shows that layers with lower average density typically display more diverse densities among heads, confirming the need for heterogeneity within the same layer.
+D.2 Statistics on Rules Discovered by MoA
 Figure 11: The MoA mask density across layers for different LLMs.
 Figure 12: The MoA mask’s average density and the density range for each layer for different LLMs.
-This subsection provides empirical evidence for rules discovered by MoA as mentioned in Section 6.5.
-The lines and spans in Figure 10 show that all heads at the first few layers generally need a high KV-Cache density.
-Following that, a few layers generally only require medium density.
-Then, in the final layers, most heads require low density, while some outlier heads need high density.
-This observation conforms to previous findings of the intrinsic dimension of LLM Valeriani et al.
-(2023).
-The geometry of density is similar to the intrinsic dimension of LLM, with two local minima.
+This subsection provides empirical evidence for rules discovered by MoA, as mentioned in Section D.1.
+The lines and spans in Figure 11 show that all heads in the first few layers generally need a high KV-Cache density and are thus assigned larger attention spans.
+Subsequently, a few layers generally require only medium density.
+In the final layers, most heads require low density, while some outlier heads need high density.
+This observation conforms to previous findings about the intrinsic dimension of LLMs Valeriani2023hiddenRepresent.
+The geometry of density is similar to the intrinsic dimension, with two local minima.
 As observed in Figure 11, layers with lower average density (smaller values on the lines) typically display a wider range of density (wider shades).
-Figure 11 validates such observation.
-This observation confirms the need for heterogeneous attention rules within the same layer.
-C.2 Connections Between MoA Rule and Semantic
-In this section, we invest the masks acquired with MoA and show the interpretable semantics of the masks.
-Previous works manually restrict the attention pattern of the model, which may harm the semantics learned by the dense model.
-In contrast, MoA preserves the semantics with statistic analysis and optimization.
-We use visualization, human interpretation and quantitive methods to analyze the semantics of the original model and to verify whether MoA captures such semantics.
-C.2.1 Mask Visualization and Semantic Categorization
-Given any token, two kinds of information are used as the model inputs: position encoding and token embedding.
-Position encoding indicates the absolute (Zhang et al., 2022) or relative positions (Touvron et al., 2023) of tokens in the sentence.
+Figure 12 validates this observation and confirms the need for heterogeneous attention rules within the same layer.
+D.3 Connections Between MoA Rule and Semantic
+In this section, we investigate the masks acquired with MoA and show their interpretable semantics.
+Previous works manually restrict the model’s attention pattern, which may harm the semantics learned by the dense model.
+In contrast, MoA preserves semantics with statistical analysis and optimization.
+We use visualization, human interpretation, and quantitative methods to analyze the semantics of the original model and verify whether MoA captures them.
+D.3.1 Mask Visualization and Semantic Categorization
+Given any token, two kinds of information are used as the model inputs: positional encoding and token embedding.
+Position encoding indicates the absolute (zhang2205opt) or relative positions (touvron2023llama2) of tokens in the sentence.
 Token embedding maps different tokens as different vectors.
-The attention head $h$ responds to both information and output the corresponding attention value $A_{h}$.
+Attention head $h$ responds to both types of information and outputs the corresponding attention value $A_{h}$.
 As shown in equation 7, we denote the influence of position and token of head $h$ as function $P_{h}$ and $T_{h}$, respectively.
@@
 Figure 2 visualizes two typical heads that are either dominated by position $P$ or token $T$ function.
-For the first attention head in Figure 2, the local positional attention is clearly observed.
-In this head, whatever sentences are given, each token pays major attention to the first token and the prior token.
-As a result, the mean attention matrix accumulates extremely large attention values at the first column and the sub-diagonal.
-In contrast, the second attention head in Figure 2 lays more emphasis on content-based attention.
-Since the position distribution of important tokens are generally random, the attention matrix can show large attention values at any position.
-It results in a mean attention matrix without extreme mean attention values.
-In conclusion, the mean attention matrix of different sentences provides a valuable insight of whether attention values of an attention head is more position-based or content-based.
-Intuitively, the more uneven the attention matrix value distribution is, the more position-based the head is.
-C.2.2 Quantitative Semantic Analysis
-We quantify how much the attention head is position-based and analyze whether MoA successfully utilizes such semantics through the evaluate-generate-optimization pipeline.
+For the first attention head in Figure 2, local positional attention is clearly observed.
+In this head, regardless of the sentence, each token pays major attention to the first token and the prior token.
+As a result, the mean attention matrix accumulates extremely large attention values in the first column and on the sub-diagonal.
+In contrast, the second attention head in Figure 2 places more emphasis on content-based attention.
+Since the positional distribution of important tokens is generally random, the attention matrix can show large attention values at any position, resulting in a mean attention matrix without extreme values.
+In conclusion, the mean attention matrix across different sentences provides valuable insight into whether an attention head is more position-based or content-based.
+Intuitively, the more uneven the distribution of attention matrix values is, the more position-based the head is.
+D.3.2 Quantitative Semantic Analysis
+We quantify how position-based an attention head is and analyze whether MoA successfully utilizes such semantics through the evaluate–generate–optimize pipeline.
 We model Equation 7 with a linear approximation.
-$P_{h}$ and $T_{h}$ are random variables with the same expectation $\mu$ and standard variance $\delta$ for all heads.
-For attention head $h$, the weight factor $\alpha_{h}$ evaluates the relatively influence of position and token to the final attention value.
+$P_{h}$ and $T_{h}$ are random variables with the same expectation $\mu$ and standard deviation $\delta$ for all heads.
+For attention head $h$, the weight factor $\alpha_{h}$ evaluates the relative influence of position and token on the final attention value.
 $A_{h,i,j}=\alpha_{h}P_{h}(i,j)+(1-\alpha_{h})T_{h}(t_{i},t_{j})$ (8)
@@
 We name $\sigma_{p}(\mathbb{E}_{t}[A_{h,i,j}])$ the Standard division of Expectation (SoE) of head $h$.
-Note that the expectation is taken over different sentences, while the standard division is taken over different attention positions.
+Note that the expectation is taken over different sentences, while the standard deviation is taken over different attention positions.
 Since $\delta_{p}$ is the same for all heads, we derive that the position impact $\alpha_{h}$ is proportional to the SoE of different heads.
-The conclusion quantifies the observation stated in Section C.2.1.
-Intuitively, SoE shows how uneven the mean attention matrix is, thus showing the influence of position to the attention values.
-MoA’s generated mask density shows positive relation with SoE, suggesting that MoA successfully captures the semantic information of the dense language model as shown in Figure 12.
-Appendix D Automatic Pipeline Details
-D.1 Addtional Oracle on Elastic Pattern Design
+This conclusion quantifies the observation stated in Section D.3.1.
+Intuitively, SoE indicates how uneven the mean attention matrix is and thus reflects the influence of position on the attention values.
+MoA’s generated mask density shows a positive relationship with SoE, suggesting that MoA successfully captures the semantic information of the dense language model, as shown in Figure 13.
+Appendix E Automatic Pipeline Details
+E.1 Additional Oracle on Elastic Pattern Design
 Figure 14: Examples of attention matrices from different attention heads (columns) and tasks (rows) of the Vicuna-7B model.
@@
 The consistent attention span across tasks sheds light on the strong cross-dataset generalization ability of our MoA method.
-D.2 Derivation of Attention Influence
+E.2 Derivation of Attention Influence
 We use the first-order Taylor expansion to calculate the influence of each attention value.
-This approximation approach is supported by methodologies commonly employed in other LLM compression approaches (Li et al., 2024a; Shi et al., 2021; Das et al., 2023; Jiang et al., 2023).
+This approximation approach is supported by methodologies commonly employed in other LLM compression approaches (llm-mq; shi2021sparsebert; das2023beyond; jiang2023pushing).
 As discussed in Section 4.1, when masking out attention value $A_{h,i,j}$ at head $h$, row $i$, and column $j$, it also influences the attention values in the same row by $\Delta A_{h,i,n|j}$.
 $A_{h,i,n}$ $=\frac{e^{S_{h,i,n}}}{\sum_{j}e^{S_{h,i,j}}}$ (11)
-$\Delta A_{h,i,n|j}$ =$=\begin{cases}-A_{h,i,n},&n=j\\ A_{h,i,n}({\sum_{j}e^{S_{h,i,j}}}/{\sum_{j\neq n}e^{S_{h,i,j}}}-1),&n\neq j\\ \end{cases}$
+[…]
 Following the definition, the attention influence $\textbf{E}_{h}$ is calculated as follows:
 $E_{h,i,j}=\sum_{n}\frac{\partial L}{\partial A_{h,i,n}}\cdot\Delta A_{h,i,n|j}$ (12)
-Given Equation 11 and  10, we derive Equation 3 as follows. For notation simplicity, we omit the head index $h$ here.
+Given Equation 12 and  11, we derive Equation 3 as follows. For notation simplicity, we omit the head index $h$ here.
 $E_{i,j}$ $=\sum_{n}\frac{\partial L}{\partial A_{i,n}}\cdot\Delta A_{i,n|j}$ (13)
@@
 $\textbf{E}_{h}=\frac{\textbf{A}_{h}}{1-\textbf{A}_{h}}\cdot\left(\frac{\partial L}{\partial\textbf{A}_{h}}-\left(\frac{\partial L}{\partial\textbf{A}_{h}}\cdot\textbf{A}_{h}\right)\mathbbm{1}^{N\times N}\right).$ (14)
-D.3 Optimization Details
-D.3.1 Optimizing at Single Length
+E.3 Optimization Details
+E.3.1 Optimizing at Single Length
 The optimization problem is formulated as follows:
@@
 $0\leq X_{h,r_{h}}\leq 1,X_{h,r_{h}}\in\mathbb{Z},$ $\quad\forall h\in\left\{0,\cdots,H-1\right\},\forall r_{h}\in\mathbb{R}$ (16d)
-In this formulation, (15a) serves as the objective function to minimize the loss, subject to the constraints that each matrix selects exactly one compression plan (15b), and the average density does not exceed $d_{\mathrm{constr}}$ (15c).
+In this formulation, (16a) serves as the objective function to minimize the loss, subject to the constraints that each matrix selects exactly one MoA configuration (16b), and the average density does not exceed $d_{\mathrm{constr}}$ (16c).
 Finally, (16d) enforces that $X_{h,r_{h}}$ is a binary variable, indicating the selection of plans.
 Additionally, to enforce the restriction that each model layer only has a limited number of different plans, we bound the norm of element-wise multiplication of $\mathbf{X}_{h}=\begin{bmatrix}X_{h,0}&X_{h,1}&\cdots&X_{h,R_{h}-1}\end{bmatrix}^{\top}$ in a single layer.
-D.3.2 Optimizing at Multiple Lengths
+E.3.2 Optimizing at Multiple Lengths
+Figure 15: Illustration of our multi-objective Mixed-Integer Programming (MIP) approach, using a two-objective optimization example: (a) MoA first minimizes the loss for 4k inputs and records the corresponding loss for the current optimal plan at 8k.
+(b) Next, it minimizes the loss for 8k inputs and records the corresponding loss for the current optimal plan at 4k.
+These steps establish the loss ranges $R$ for both 4k and 8k input lengths.
+(c) MoA then re-optimizes the loss at 4k, this time using the loss intervals at 8k as different constraints.
+All plans generated under these constraints are recorded.
+(d) The last process (c) is repeated for 8k, using 4k intervals as constraints.
+Finally, plans meeting the Pareto front criteria for both 4k and 8k inputs are selected as the final outputs.
 With the ability to optimize at a single length, we utilize the same framework for multi-objective MIP across various lengths.
-The key is to transform the multi-objective MIP problem into several single-objective MIP problems (Paria et al., 2018).
-We utilize the idea of epsilon-constraint method (Yv et al., 1971).
+The key is to transform the multi-objective MIP problem into several single-objective MIP problems (Paria2018AFF).
+We utilize the idea of epsilon-constraint method (haimes1971bicriterion).
+Figure 15 illustrates the optimization process for two input lengths.
+We discuss the generalized approach to handle an arbitrary number of lengths.
 We first select each input length as our primary objective to perform the single-objective optimization on it while simultaneously recording the outcomes of other objectives.
@@
 $\mathop{\arg\min}_{r_{h}\in\mathbb{R}}\Delta L^{(N_{i})}\quad\operatorname{s.t.}\frac{1}{H}\sum_{h=1}^{H}d^{(N_{i})}_{r_{h}}\leq d^{(N_{i})}_{\text{constr}},\forall N_{i}\in\mathbb{N}_{\text{constr}};\quad\Delta L^{(N_{j})}\in S^{(N_{j})}_{k_{j}},\forall j\neq i.$ (17)
-where this optimization is performed for each $i$ ranging from 00 to $N$.
-For each $j$, $k_{j}$ can vary independently from 00 to $M$.
+where this optimization is performed for each $i$ ranging from 0 to $N$.
+For each $j$, $k_{j}$ can vary independently from 0 to $M$.
 For efficiency consideration, we set the number of intervals as five.
 Finally, the results that do not conform to the Pareto front requirements are removed, resulting in the final Pareto front set of our multi-objective optimization problem.
+E.4 Coding Interface
+MoA offers a straightforward and intuitive interface for configuration search and inference.
+For supported HuggingFace LLM architectures, the entire search pipeline for any given new model is fully automated with a single command:
+python scripts/pipeline/main.py –model_path <HF_MODEL_PATH> –model_name <NAME>
+Once search is completed, the MoA configuration is as straightforward to deploy as its dense counterpart.
+MoA seamlessly integrates with the HuggingFace framework and remains fully compatible with high-level methods like pipeline and generate.
+Switching from standard attention to our mixture-structured sliding-window attention requires minimal adjustments.
+We provide a specialized CUDA GPU kernel for heterogeneous sliding-window attention, integrated into HuggingFace’s AttentionModule interface.
+The adoption involves just two lines of code:
+model = update_model_function(model, model_name)
+model.model.set_mixture_of_attention(moa_config)
+By automating the MoA configuration search process and offering a simple deployment interface, our approach minimizes complexity and ensures practical usability.
+Appendix F Limitations and Future Work
+Under an extremely low-density budget, MoA fails to maintain good performance.
+Designing a dynamic MoA method has the potential to address this issue, which we leave for future work.
+Using non-linear elastic rules with bounded attention spans is also worth exploring.
+Additionally, MoA’s profiling method can be adapted to evaluate the influence of weights and other activations, facilitating other compression methods such as quantization.